### v1

# Global identification of dynamic panel models with interactive effects

Abstract
This paper examines the problem of global identification in dynamic panel models with interactive effects, a fundamental issue in econometric theory. We focus on the setting where the number of cross-sectional units ($N$) is large, but the time dimension ($T$) remains fixed. While local identification based on the Jacobian matrix is well understood and relatively straightforward to establish, achieving global identification remains a significant challenge. Under a set of mild and easily satisfied conditions, we demonstrate that the parameters of the model are globally identified, ensuring that no two distinct parameter values generate the same probability distribution of the observed data. Our findings contribute to the broader literature on identification in panel data models and have important implications for empirical research that relies on interactive effects.
Keywords: Dynamic panel models, global identification, factor models, panel data, interactive effects
1 Introduction
Consider the dynamic panel data model with interactive effects
$y_{it}=\alpha y_{it-1}+\delta_{t}+\lambda_{i}^{{}^{\prime}}f_{t}+\varepsilon_{it}$ (1)
where $i=1,2,\dots,N$ and $t=1,2,\dots,T$. $y_{it}$ is the outcome variable, $\alpha$ is the autoregressive parameter, whereas $\delta_{t}$ captures time-specific fixed effects. $f_{t}$ is an $\bar{r}\times 1$ vector containing the value of the $\bar{r}$ latent factors at time $t$, and $\lambda_{i}$ is an $\bar{r}\times 1$ vector of factor loadings for individual $i$. $\varepsilon_{it}$ is the idiosyncratic error term which we assume to be heteroskedastic but not autocorrelated i.e. $\mathbb{E}[\varepsilon_{i,t}^{2}]=\sigma_{t}^{2}\;\text{and}\;\mathbb{E}[\varepsilon_{i,t}\varepsilon_{i,s}]=0\;$ for $t\neq s$, and independent from the interactive effects. In this model, $y_{it}$ is the only observable variable.
This framework provides a flexible approach to estimate unobservable trends from the data. The model allows the common shocks (modeled by $f_{t}$) to have a heterogeneous effect in individuals (modeled by $\lambda_{i}$) on the outcome variable $y_{it}$. A notable special case is the standard additive fixed-effects model, where $y_{it}=\delta_{t}+\mu_{i}+\varepsilon_{it}$, which corresponds to $\alpha=0$ and a single constant factor $f_{t}=1$. Thus, this model generalizes the additive framework by allowing for richer dynamics and heterogeneity in the factor structure. Parameter estimation in this model can be conducted via quasi maximum likelihood based methods as in Hayakawa et al. (2023) and Bai (2024). This paper studies the problem of global identification of the model parameters.
Following Rothenberg (1971), we say that a parameter point $\theta^{0}$ in the parametric space $\Theta$ is globally identifiable if there is no other $\theta$ in $\Theta$ which is observationally equivalent. Establishing the identifiability of a model’s parameters is crucial, as it ensures that the estimation procedure yields meaningful and interpretable results.
Despite the apparent simplicity of the model, demonstrating parameter identifiability is a nontrivial task. The seminal work of Anderson and Rubin (1956) established that in the absence of the autoregressive term ($\alpha=0$), the parameters of the model are globally identified up to an indeterminacy given by an $\bar{r}\times\bar{r}$ rotation matrix, provided that $T\geq 2\bar{r}+1$.
To the best of our knowledge, no existing study has established the global identifiability of the parameters in the model described by (1). Bai (2024) demonstrates that the parameters of this model are locally identified when $\bar{r}$ does not exceed the modified Ledermann bound (Ledermann, 1937). In contrast, Hayakawa et al. (2023) argues that global identification is unattainable in dynamic panel data models with interactive effects.
We demonstrate that under standard assumptions in the factor analysis literature, the parameters of the model are almost surely globally identifiable when $T\geq 2(\bar{r}+1)$. The global identification is achieved using the first two moments, which is equivalent to global identification for the quasi Gaussian maximum likelihood method. This result is of particular importance, as it establishes a one-to-one correspondence between the population moments and the model parameters, despite high-degree of nonlinearity of the problem.
This paper contributes to the literature on global identification and panel data models. The question of global identification in parametric models has been the focus of extensive research (see, for example, Shapiro (1985), Komunjer and Ng (2011), Komunjer (2012), and Kocikecki and Kolasa (2023)), as well as trying to understand when and why identification fails (Forneron, 2024). A comprehensive review of this literature is provided in Lewbel (2019).
Panel data models have been a core focus in econometrics for many years, with extensive study and application over the past few decades. Notable monographs and textbooks on panel data include works by Arellano (2003), Baltagi (2008), Hsiao (2022), and Wooldridge (2010), among others. More recently, there has been significant research on dynamic panel models with interactive effects, including studies by Moon and Weidner (2017), Bai (2024), and Hayakawa et al. (2023) to name a few. However, these studies have not addressed the global identifiability of the model parameters, a gap this paper aims to fill.
The remainder of the paper is organized as follows: Section 2 formalizes the identification problem in our context. Section 3 outlines the assumptions required for our proof and presents some useful mathematical results that will be instrumental in our analysis. Section 4 derives intermediate results and provides the main proof of global identification under the stated assumptions. Finally, section 5 provides some concluding remarks with a discussion of potential extensions of the paper.
2 Framework
2.1 The Model in Matrix Notation
Before formally defining global identification, it is useful to express the model in matrix notation. Suppose we observe $y_{i}=(y_{i1},y_{i2},\dots y_{i,T})^{\prime}$. We begin by projecting the initial observation $y_{i1}$ onto $[1,\lambda_{i}]$, which gives
$y_{i1}=\delta_{1}^{*}+\lambda_{i}^{{}^{\prime}}f_{1}^{*}+\varepsilon_{i1}^{*}$
where $\delta_{1}^{*}$ and $f_{1}^{*}$ are the projection coefficients, and $\varepsilon_{i1}^{*}$ is the projection residual. Since $\delta_{1}^{*}$ and $f_{1}^{*}$ are free parameters, and $\varepsilon_{it}$ in the model are not required to have identical distributions, we can simply rename $\delta_{1}^{*},f_{1}^{*}$, and $\varepsilon_{i1}^{*}$ as $\delta_{1}$, $f_{1}$, and $\varepsilon_{i1}$, respectively, without any loss of generality. So we can write $y_{i1}=\delta_{1}+\lambda_{i}^{\prime}f_{1}+\varepsilon_{i1}$.
Next, stacking observations over $t$, we obtain
$\mathbf{B}y_{i}=\delta+\mathbf{F}\lambda_{i}+\varepsilon_{i}$
where $y_{i}=[y_{i1},y_{i2},\dots,y_{iT}]^{\prime}$ is a $T\times 1$ vector of individual observations for $t=1,\dots,T$, $\delta=[\delta_{1},\dots\delta_{T}]^{\prime}$ is a $T\times 1$ vector, and $\mathbf{F}=[f_{1},f_{2},\dots,f_{T}]^{\prime}$ is a $T\times\bar{r}$ matrix containing the $\bar{r}$ factors, $\lambda_{i}=[\lambda_{i1},\lambda_{i2},\dots,\lambda_{i\bar{r}}]^{\prime}$ is an $\bar{r}\times 1$ vector of factor loadings for individual $i$, and $\varepsilon_{i}=[\varepsilon_{i1},\varepsilon_{i2},\dots,\varepsilon_{i,T}]^{\prime}$ is a $T\times 1$ vector of idiosyncratic errors. Furthermore,
$\mathbf{B}=\begin{bmatrix}1&0&0&\dots&0\\ -\alpha&1&0&\dots&0\\ 0&-\alpha&1&\dots&0\\ \vdots&\vdots&\ddots&\ddots&\vdots\\ 0&\dots&0&-\alpha&1\end{bmatrix}$
a $T\times T$ matrix. Let $\mathbf{\Gamma}=\mathbf{B^{-1}}.$ Pre-multiplying by $\mathbf{B^{-1}}$ allows us to rewrite the model as
$y_{i}=\mathbf{\Gamma}\delta+\mathbf{\Gamma}\mathbf{F}\lambda_{i}+\mathbf{\Gamma}\varepsilon_{i}$ (2)
where $\mathbf{\Gamma}$ is $T\times T$ and is given by
$\mathbf{\Gamma}=\mathbf{B^{-1}}=\begin{bmatrix}1&0&\dots&0\\ \alpha&1&\dots&0\\ \vdots&\vdots&\ddots&\vdots\\ \alpha^{T-1}&\dots&\alpha&1\end{bmatrix}$
In the original model the errors had a diagonal covariance matrix given by $\mathbf{D}=\operatorname*{diag}(d_{1},d_{2},\dots,d_{T}):=\operatorname*{diag}(\sigma_{1}^{2},\dots,\sigma_{T}^{2})$ (for notational simplicty, we write $\sigma_{t}^{2}$ as $d_{t}$). After transformation, the idiosyncratic error $\mathbf{\Gamma}\varepsilon_{i}$ has covariance matrix $\mathbf{\Gamma}\mathbf{D}\mathbf{\Gamma}^{\prime}$, which is no longer diagonal. Let
$\mathbf{\Psi}_{N}=\frac{1}{N-1}\sum_{i=1}^{N}(\lambda_{i}-\bar{\lambda})(\lambda_{i}-\bar{\lambda})^{\prime}$
denote the sample covariance matrix of $\lambda_{i}$, where $\bar{\lambda}=\frac{1}{N}\sum_{n=1}^{N}\lambda_{i}$ and let $\mathbb{E}[\mathbf{\Psi}_{N}]=\mathbf{\Psi}$. Similarly, define
$\mathbf{S}_{N}=\frac{1}{N-1}\sum_{i=1}^{N}(y_{i}-\bar{y})(y_{i}-\bar{y})^{\prime}$
as the sample covariance matrix of $y_{i}$, where $\bar{y}=\frac{1}{N}\sum_{n=1}^{N}y_{i}$.
The expectation of the sample covariance matrix of $y_{i}$ is then given by
$\mathbb{E}[\mathbf{S}_{N}]=\mathbf{\Sigma}(\theta)=\mathbf{\Gamma}(\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}+\mathbf{D})\mathbf{\Gamma}^{\prime}=\mathbf{\Gamma}\mathbf{\Omega}\mathbf{\Gamma}^{\prime}$ (3)
where $\theta=\text{vec}(\alpha,\mathbf{F},\mathbf{\Psi},\mathbf{D})$ represents the model parameters. The matrix $\mathbf{\Sigma}$ is a $T\times T$ symmetric positive definite (PD) matrix, $\mathbf{\Psi}$ is an $\bar{r}\times\bar{r}$ symmetric PD matrix, $\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}$ is a $T\times T$ symmetric positive semi-definite (PSD) matrix of rank $\bar{r}$, and $\mathbf{D}$ is a $T\times T$ diagonal PD matrix. $\mathbf{\Omega}=\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}+\mathbf{D}$ is a $T\times T$ symmetric PD matrix corresponding to the expected value of sample variance of $y_{i}$ in a basic model without an autoregressive term $\alpha=0$ (Anderson and Rubin, 1956).
2.2 Global identification of the model
This paper aims to establish that the parameters of the model are globally identifiable from the quasi log-likelihood perspective, which is equivalent to demonstrating global identification from the first two moments. Specifically, this means that the expectation and variance of $y_{i}$ uniquely determine the structural parameters of the model, as there exists a one-to-one mapping between them.
To formalize this argument, define $\delta^{\dagger}=\bm{\Gamma}\delta$. The quasi Gaussian log-likelihood function is given by
$-\frac{N}{2}\ln|\bm{\Sigma}(\theta)|-\frac{1}{2}\sum_{i=1}^{N}(y_{i}-\delta^{\dagger})^{\prime}\bm{\Sigma}(\theta)^{-1}(y_{i}-\delta^{\dagger})$
Concentrating out $\delta^{\dagger}$, which is estimated by $\bar{y}$, and dividing by $N$, we obtain [^1] In the MLE, the definition of $\mathbf{S}_{N}$ needs a slight modification, namely, $1/(N-1)$ is replaced by $1/N$.
$\frac{1}{N}\ell_{N}(\theta)=-\frac{1}{2}\ln|\bm{\Sigma}(\theta)|-\frac{1}{2}\mathrm{tr}[\bm{\Sigma}(\theta)^{-1}\mathbf{S}_{N}]$
[…]
Furthermore, it is known (Arnold, 1981) that this function has a unique maximizer at $\mathbf{\Sigma}(\theta)=\mathbf{\Sigma}(\theta^{0})$. While this result ensures that $\theta^{0}$ is a global maximizer of the objective function, it does not immediately establish uniqueness—hence, global identification—because it does not preclude the possibility that other parameter values $\widetilde{\theta}\neq\theta^{0}$ satisfy $\mathbf{\Sigma(\widetilde{\theta})}=\mathbf{\Sigma}(\theta^{0})$. In the following sections, we investigate whether any alternative parameter set $\widetilde{\theta}\neq\theta^{0}$ satisfies $\mathbf{\Sigma}(\widetilde{\theta})=\mathbf{\Sigma}(\theta^{0})$. We demonstrate that, almost surely, no such alternative parametrization exists. Consequently, the quasi log-likelihood function is uniquely maximized at the true parameter value $\theta^{0}$, establishing global identification. This result, which has not been previously shown and was thought to be untrue (Hayakawa et al., 2023), constitutes a novel contribution.
To prove this claim, consider an alternative parameterization
  • $\mathbf{\widetilde{F}}=\left[\widetilde{f}_{1},\widetilde{f}_{2},\dots,\widetilde{f}_{T}\right]^{\prime}$;
  • $\mathbf{\widetilde{\Psi}}$ a symmetric PD $\bar{r}\times\bar{r}$ matrix;
  • $\mathbf{\widetilde{D}}=\operatorname*{diag}(\widetilde{d}_{1},\dots\widetilde{d}_{T})$ a diagonal PD $T\times T$ matrix;
  • $\widetilde{\mathbf{\Gamma}}$ the counterpart to $\mathbf{\Gamma}$ under an alternative parameter $\widetilde{\alpha}$, given by
$\mathbf{\widetilde{\Gamma}}=\begin{bmatrix}1&0&\dots&0\\ \widetilde{\alpha}&1&\dots&0\\ \vdots&\vdots&\ddots&\vdots\\ \widetilde{\alpha}^{T-1}&\dots&\widetilde{\alpha}&1\end{bmatrix}$
By definition, $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}$ is a symmetric PSD $T\times T$ matrix of rank $\bar{r}$.
We say that the parameters in the dynamic panel data model with interactive effects are globally identified at $\theta=\text{vec}(\alpha,\mathbf{F},\mathbf{\Psi},\mathbf{D})$ if there does not exist a diagonal and PD matrix $\mathbf{\widetilde{D}}$, a parameter $\widetilde{\alpha}$, and a symmetric PSD matrix $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}$ of rank $\bar{r}$, with at least one of them different from $\mathbf{D}$, $\alpha$, and $\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}$, respectively, such that
$\mathbf{\Sigma}=\mathbf{\Gamma}(\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}+\mathbf{D})\mathbf{\Gamma}^{\prime}=\mathbf{\widetilde{\Gamma}}(\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}+\mathbf{\widetilde{D}})\mathbf{\widetilde{\Gamma}}^{\prime}$
Equivalently, we want to show
$\mathbf{\Gamma}(\mathbf{F}\mathbf{\Psi}\mathbf{F^{\prime}}+\mathbf{D})\mathbf{\Gamma}^{\prime}=\mathbf{\widetilde{\Gamma}}(\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}+\mathbf{\widetilde{D}})\mathbf{\widetilde{\Gamma}}^{\prime}$ (5)
if and only if $\mathbf{\widetilde{D}}=\mathbf{D}$, $\mathbf{\widetilde{\Gamma}}=\mathbf{\Gamma}$ (or equivalently $\widetilde{\alpha}=\alpha$), and $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}=\mathbf{F}\mathbf{\Psi}\mathbf{F^{\prime}}$.
3 Assumptions and useful results
In order to proceed with our global identification proof, we need to make some assumptions. These assumptions are not restrictive and are rather standard in the factor analysis literature.
Assumption 1.
The number of distinct factors is known and equal to $\bar{r}$
Assumption 2.
Identification restrictions IC1 or IC2:
IC1: $\mathbf{\Psi}$ an unrestricted PD matrix and $\mathbf{F}=(\mathbf{I}_{\bar{r}},\mathbf{F}_{2}^{{}^{\prime}})^{\prime}$
IC2: $\mathbf{\Psi}$ is diagonal and $T^{-1}\mathbf{F}^{{}^{\prime}}\mathbf{D^{-1}}\mathbf{F}=\mathbf{I}_{\bar{r}}$
This assumption comes from the fact that factor models (and thus interactive effects) are identifiable up to a $\bar{r}\times\bar{r}$ matrix rotation, so $\bar{r}^{2}$ restrictions must be imposed for identification. A discussion of why this is the case can be found in Section 5 of Anderson and Rubin (1956). We will work with IC1 for the rest of the paper. But it can be shown that the model is globally identifiable under IC2 (with some modifications).
Assumption 3.
$T\geq 2(\bar{r}+1)$
Assumption 4.
$\lambda_{i}$ are either random or fixed constants with $\bm{\Psi}_{N}\to\bm{\Psi}$ as $N\to\infty$.
Assumption 5.
$\alpha,\mathbf{\Psi},d_{1},d_{2},\dots,d_{T}$ are fixed population parameters.
Assumption 6.
Let $\mathbf{F}_{2}=\begin{bmatrix}f_{\bar{r}+1,1}&f_{\bar{r}+1,2}&\dots&f_{\bar{r}+1,\bar{r}}\\ \vdots&\vdots&\vdots&\vdots\\ f_{T,1}&f_{T,2}&\dots&f_{T,\bar{r}}\end{bmatrix}$ be the random elements of $\mathbf{F}$. Then, $\mathbf{F}_{2}$ is a continuous random matrix with support in $\Xi\in\mathbb{R}^{(T-\bar{r})\times\bar{r}}$.
Additionally, throughout the paper, we will reference several established results from linear algebra and probability theory. To ensure they are easily accessible to the reader, we will present them as lemmas in this section.
Lemma 1.
Let $\mathbf{A}$ be a $T\times T$ matrix of rank $\bar{r}<T$. Then, the determinants of all the $(\bar{r}+1)\times(\bar{r}+1)$ minors of $\mathbf{A}$ are equal to 0.
Lemma 2.
Let $\mathbf{X}$ be a continuous random vector with $\mathrm{supp}(\mathbf{X})\subset\mathbb{R}^{q}$. Furthermore, let $\Upsilon$ be a subset of $\mathrm{supp}(\mathbf{X})$ with dimension lower than that of $\mathrm{supp}(\mathbf{X})$. Then
$P(\mathbf{X}\in\Upsilon)=0$
Lemma 3.
(Laplace Expansion) Let $\mathbf{A}$ be a $k\times k$ matrix. Then, we can write the determinant of $\mathbf{A}$ as a weighted sum of the determinants of some $(k-1)\times(k-1)$ minors of $\mathbf{A}$. In particular, for each row $i$, we can write
$\det\left(\mathbf{A}\right)=\sum_{j=1}^{k}(-1)^{i+j}A_{i,j}\det\left(\mathbf{A}_{-i,-j}\right)$
where $\mathbf{A}_{-i,-j}$ is a minor of $\mathbf{A}$ obtained from removing row $i$ and column $j$ of matrix $\mathbf{A}$, and $A_{i,j}$ represents the $(i,j)$-th element of the matrix $\mathbf{A}$.
In this paper, we will apply Lemma 3 for cases in which $\mathbf{A}$ itself is a minor of a larger matrix. Let $\mathbf{H}$ be a $T\times T$ matrix, and let $\mathbf{M}^{\mathbf{H}}_{R,C}$ be a minor of $\mathbf{H}$ of dimension $k$ including rows $R=\{r_{1},\dots,r_{k}\}$ and columns $C=\{c_{1},\dots,c_{k}\}$ of the original matrix. Therefore, for any $r\in R$ the Laplace decomposition of this minor is given by:
$\det(\mathbf{M}^{\mathbf{H}}_{R,C})=\sum_{j=1}^{k}(-1)^{r+j}H_{r,c_{j}}\det\left(\mathbf{M}^{\mathbf{H}}_{R-r,C-c_{j}}\right)$
where $H_{r,c_{j}}$ is the $(r,c_{j})$-th element of $\mathbf{H}$, and $\mathbf{M}^{\mathbf{H}}_{R-r,C-c_{j}}$ is the minor [^2] Where $R-r$ means removal $r$ from the set $R$, and $C-c_{j}$ means removal of $c_{j}$ from the set $C$. They correspond to the usual notation $R\setminus\{r\}$ and $C\setminus\{c_{j}\}$. of $\mathbf{H}$ that includes rows $R-r$, and columns $C-c_{j}$.
In our induction analysis to be used later, we will further need to indicate the dimension of the minor. The minor on the left hand side will be written as $\mathbf{M}^{\mathbf{H},k}_{R,C}$, and the minor on the right hand side will be written as $\mathbf{M}^{\mathbf{H},k-1}_{R-r,C-c_{j}}$
4 Global identification
In this section, we demonstrate that, under Assumptions 1-6, the parameters of the dynamic panel data model with interactive effects described in (1) are globally identified.
4.1 Proof insight
The proof we propose draws inspiration from Theorem 5.1 of Anderson and Rubin (1956). In that theorem, the authors establish a sufficient condition for global identification of a factor model that did not include an autoregressive component. In their simpler model, global identification requires
$\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}+\mathbf{D}=\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}+\mathbf{\widetilde{D}}$ (6)
If and only if $\mathbf{\widetilde{D}}=\mathbf{D}$ (a $T\times T$ diagonal PD matrix), and $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}=\mathbf{F}\mathbf{\Psi}\mathbf{F^{\prime}}$ (a $T\times T$ symmetric PSD matrix of rank $\bar{r}$).
Their proof relies on the fact that the off-diagonal elements of $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}$ must match those of $\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}$ for (6) to hold, because both $\mathbf{D}\ \text{and}\ \mathbf{\widetilde{D}}$ are diagonal matrices. Thus, the problem reduces to showing that the diagonal elements of these matrices are also equal. To establish this, they rewrite (6) as
$\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}=\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}+\mathbf{\widetilde{D}}-\mathbf{D}$
Then, they leverage the fact that all determinants of the $(\bar{r}+1)\times(\bar{r}+1)$ minors of $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}+\mathbf{\widetilde{D}}-\mathbf{D}$ must be zero for (6) to hold (Lemma 1). This follows as a necessary condition because, by assumption, $\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}$ is of rank $\bar{r}$. This procedure enables them to pin-down the diagonal terms of $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}$, thereby establishing that $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}=\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}$. Once this is established, it follows immediately that $\mathbf{\widetilde{D}}=\mathbf{D}$, completing the proof.
In their approach, the authors focus exclusively on the minors involving diagonal elements of the matrix and do not consider what we will refer to as diagonal exclusion minors (minors that do not include diagonal terms of the original matrix) as these do not contain relevant information for their analysis. In contrast, our proof takes a different approach: we rearrange the terms of our original problem—Equation (5)—in a way that allows us to exploit the information contained in the diagonal exclusion minors of this transformed problem. This novel approach enables us to demonstrate that (5) holds if and only if $\widetilde{\alpha}=\alpha$, thereby establishing the global identification of $\alpha$.
Once we establish that $\alpha$ is globally identified, it follows that our identification condition can be rewritten in the form proposed by Anderson and Rubin (1956), ensuring that all other parameters can also be globally identified.
Thus, the primary focus of our analysis is to show that $\alpha$ is globally identified (or identically, that it must be the case that $\widetilde{\alpha}=\alpha$ in order for (5) to hold). This task requires some intermediate results, which we will articulate and prove throughout the paper. The main contribution of this work is to formally establish the following Theorem.
Theorem 1.
Let $y_{it}$ be generated from (1). Then, under Assumptions 1-6, $\alpha$ is almost surely globally identified.
Once this shown, it is immediate to conclude that all the parameters of the model are almost surely globally identified as well, which is formally stated in the following Theorem
Theorem 2.
Let $y_{it}$ be generated from (1). Then, under Assumptions 1-6, all the parameters in the model are almost surely globally identified.
Remark: The almost sure identification refers to the distribution of $\mathbf{F}$, not to $\alpha$. Any value of $\alpha$ is globally identifiable. The practical implication is that as long as there is sufficient variation in $f_{t}$, both $\alpha$ and other parameters are globally identifiable. This is particularly true when $T$ is sufficiently large, ensuring a sufficient number of distinct realizations of $f_{t}$.
4.2 Intermediate results
In order to prove Theorem 1, we will rearrange the problem in a way that resembles the proof of Anderson and Rubin (1956) and then establish some properties of this modified problem. Start from (5). As $\mathbf{\widetilde{\Gamma}}$ is a unit-diagonal lower triangular matrix, it is invertible so that we can pre-multiply and post-multiply by $\mathbf{\widetilde{\Gamma}^{-1}}$ and $\mathbf{\widetilde{\Gamma}^{-1^{\prime}}}$ and re-write
$\mathbf{\widetilde{\Gamma}^{-1}}\mathbf{\Gamma}(\mathbf{F}\mathbf{\Psi}\mathbf{F^{\prime}}+\mathbf{D})\mathbf{\Gamma}^{\prime}\mathbf{\widetilde{\Gamma}^{-1^{\prime}}}=\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}+\mathbf{\widetilde{D}}$
Now, define
$\mathbf{Q}=\mathbf{\widetilde{\Gamma}^{-1}}\mathbf{\Gamma}=\begin{bmatrix}1&0&0&\dots&0\\ \alpha-\widetilde{\alpha}&1&0&\dots&0\\ \alpha(\alpha-\widetilde{\alpha})&\alpha-\widetilde{\alpha}&1&\dots&0\\ \vdots&\vdots&\ddots&\ddots&\vdots\\ \alpha^{T-2}(\alpha-\widetilde{\alpha})&\alpha^{T-3}(\alpha-\widetilde{\alpha})&\dots&\alpha-\widetilde{\alpha}&1\end{bmatrix}$
Note that $\mathbf{Q}$ can be decomposed into
$\mathbf{Q}=\mathbf{I_{T}}+(\alpha-\widetilde{\alpha})\mathbf{L}$
where
$\mathbf{L}=\begin{bmatrix}0&0&0&\dots&0\\ 1&0&0&\dots&0\\ \alpha&1&0&\dots&0\\ \vdots&\vdots&\ddots&\ddots&\vdots\\ \alpha^{T-2}&\alpha^{T-3}&\dots&1&0\end{bmatrix}$
Then, subtract $\mathbf{\widetilde{D}}$ on both sides of the equation
$\mathbf{Q}(\mathbf{F}\mathbf{\Psi}\mathbf{F^{\prime}}+\mathbf{D})\mathbf{Q}^{\prime}-\mathbf{\widetilde{D}}=\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}$ (7)
To simplify notation, denote the left-hand side of (7) by
$\mathbf{O}\equiv\mathbf{Q}(\mathbf{F}\mathbf{\Psi}\mathbf{F^{\prime}}+\mathbf{D})\mathbf{Q}^{\prime}-\mathbf{\widetilde{D}}$ (8)
We next establish some preliminary results.
Claim 1.
Let $\mathbf{M}$ be any minor of $\mathbf{O}$ of dimension $k>\bar{r}$. Then $\det(\mathbf{M})=0$ is a necessary condition for (5) to hold.
Proof: We have shown that (5) is equivalent to $\mathbf{O}=\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}$. Furthermore, $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}$ is a symmetric PSD matrix of rank $\bar{r}$ by assumption. Then, it must be the case that $\mathbf{O}$ is also a symmetric PSD matrix of rank $\bar{r}$ in order for (5) to hold. By Lemma 1 we know that if $\mathbf{O}$ is of rank $\bar{r}$ then the determinant of all the minors of $\mathbf{O}$ of dimension $k>\bar{r}$ must be equal to 0. $\square$
Furthermore, we can show that $\mathbf{O}$ can be decomposed in the following convenient way
Claim 2.
$\mathbf{O}$ can be decomposed into $\mathbf{O}=\mathbf{\Omega}-\mathbf{\widetilde{D}}+(\alpha-\widetilde{\alpha})\mathbf{J}(\widetilde{\alpha},\alpha,\mathbf{D},\mathbf{F},\mathbf{\Psi}),$
where $\mathbf{\Omega}$ is defined as in (3) and $\mathbf{J}(\widetilde{\alpha},\alpha,\mathbf{D},\mathbf{F},\mathbf{\Psi})$ is given by:
$\mathbf{J}=\mathbf{L}\mathbf{\Omega}+\mathbf{\Omega}\mathbf{L}^{\prime}+(\alpha-\widetilde{\alpha})\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}.$ (9)
Proof: By definition $\mathbf{\Omega}=\mathbf{F}\mathbf{\Psi}\mathbf{F^{\prime}}+\mathbf{D}$, so we can write $\mathbf{O}$ as
$\mathbf{O}=\mathbf{Q}\mathbf{\Omega}\mathbf{Q}^{\prime}-\mathbf{\widetilde{D}}.$
Recall that $\mathbf{Q}=\mathbf{I_{T}}+(\alpha-\widetilde{\alpha})\mathbf{L}$. Then
$\mathbf{O}=\left(\mathbf{I_{T}}+(\alpha-\widetilde{\alpha})\mathbf{L}\right)\mathbf{\Omega}\left(\mathbf{I_{T}}+(\alpha-\widetilde{\alpha})\mathbf{L}\right)^{\prime}-\mathbf{\widetilde{D}}.$
Distributing terms, we see that
$\mathbf{O}$ $=\mathbf{\Omega}-\mathbf{\widetilde{D}}+(\alpha-\widetilde{\alpha})\mathbf{L}\mathbf{\Omega}+\mathbf{\Omega}(\alpha-\widetilde{\alpha})\mathbf{L}^{\prime}+(\alpha-\widetilde{\alpha})^{2}\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}$
$=\mathbf{\Omega}-\mathbf{\widetilde{D}}+(\alpha-\widetilde{\alpha})\left[\mathbf{L}\mathbf{\Omega}+\mathbf{\Omega}\mathbf{L}^{\prime}+(\alpha-\widetilde{\alpha})\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}\right]$
$=\mathbf{\Omega}-\mathbf{\widetilde{D}}+(\alpha-\widetilde{\alpha})\mathbf{J}(\alpha,\widetilde{\alpha},\mathbf{D},\mathbf{F},\mathbf{\Psi})$
where $\mathbf{J}(\alpha,\widetilde{\alpha},\mathbf{D},\mathbf{F},\mathbf{\Psi})=\mathbf{L}\mathbf{\Omega}+\mathbf{\Omega}\mathbf{L}^{\prime}+(\alpha-\widetilde{\alpha})\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}.$ This proves the claim. $\Box$
From this point forward, we will use either $\mathbf{J}$ or $\mathbf{J}\left(\widetilde{\alpha},\theta\right)$ to refer to $\mathbf{J}(\widetilde{\alpha},\alpha,\mathbf{D},\mathbf{F},\mathbf{\Psi})$, noting that it depends on the true data generating process parameters $\theta=\text{vec}(\alpha,\mathbf{F},\mathbf{\Psi},\mathbf{D})$ and on $\widetilde{\alpha}$. With this result in hand, we can proceed to derive the following corollaries.
Corollary 1.
We can write every non-diagonal element of $\mathbf{O}$ as $O_{r,c}=\Omega_{r,c}+(\alpha-\widetilde{\alpha})J_{r,c},$ where $J_{r,c}$ is the $(r,c)$-th element of $\mathbf{J}$.
Proof: As $\mathbf{\widetilde{D}}$ is a diagonal matrix then $\widetilde{D}_{r,c}=0$ for all $r\neq c$. Using the expression for $\mathbf{O}$ derived in Claim 2:
$O_{r,c}=\Omega_{r,c}-\widetilde{D}_{r,c}+(\alpha-\widetilde{\alpha})J_{r,c}=\Omega_{r,c}+(\alpha-\widetilde{\alpha})J_{r,c}\quad r\neq c$
$\square$
Corollary 2.
For every row index $r\in\{\bar{r}+1,\dots T\}$ we have that $J_{r,c}\neq 0$ with probability 1 if $c\neq 1$. Similarly for every column index $c\in\{\bar{r}+1,\dots,T\}$ we have that $J_{r,c}\neq 0$ with probability 1 if $r\neq 1$.
Proof: We only provide the proof for the first statement, as the second one is completely analogous given the symmetry of the problem.
To show this result, we will use the fact that most of the positions of $\mathbf{\Omega}$ are non-zero with probability 1. First, note that under IC1 (Assumption 2) we know that the first $\bar{r}\times\bar{r}$ block of $\mathbf{\Omega}$ is given by:
$\mathbf{\Omega}_{1:\bar{r},1:\bar{r}}=\mathbf{\Psi}+\mathbf{D}_{1:\bar{r},1:\bar{r}}$
where $\mathbf{\Omega}_{1:\bar{r},1:\bar{r}}$ represents the first $\bar{r}\times\bar{r}$ block of $\mathbf{\Omega}$ (same for $\mathbf{D}_{1:\bar{r},1:\bar{r}}$). Therefore, we cannot guarantee that the positions of $\mathbf{\Omega}_{1:\bar{r},1:\bar{r}}$ are non-zero, as the covariance between the factor loadings (the off-diagonal elements of $\mathbf{\Psi}$) could be equal to zero. Then, we will use information contained in the remaining elements of $\mathbf{\Omega}$ to show this result.
From the definition of $\mathbf{\Omega}$ under IC1 and Assumption 6 we know that $\Omega_{r,c}\neq 0$ with probability 1 if $r>\bar{r}$, as every position depends on at least one factor realization, so the case in which the position is exactly equal to zero is an event of probability zero (Lemma 2). Now, recall that $\mathbf{J}=\mathbf{L}\mathbf{\Omega}+\mathbf{\Omega}\mathbf{L}^{\prime}+(\alpha-\widetilde{\alpha})\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}$ is a $T\times T$ matrix. By Assumptions 5 and 6 we know that the elements of $\mathbf{\Omega}_{(\bar{r}+1):T,1:T}$ follow a continuous distribution dependent on $\mathbf{F}$. We establish the corollary in three steps.
First, we show that $\mathbf{\Omega}\mathbf{L}^{\prime}$ is a random matrix with non-zero elements for rows $r>\bar{r}$ and $c\neq 1$. To see this, observe that
$\mathbf{\Omega}\mathbf{L}^{\prime}_{r,c}=\sum_{i=1}^{T}\Omega_{r,i}L^{{}^{\prime}}_{i,c}=\sum_{i=1}^{T}\Omega_{r,i}L_{c,i}$
If $c=1$ then we know that $L_{1,i}=0$ for all $i$, implying that $\mathbf{\Omega}\mathbf{L}^{\prime}_{r,1}=0$. Conversely, for $c>1$, the elements $L_{c,c-1},L_{c,c-2},\dots L_{c,1}$ are all nonzero if $\alpha\neq 0$. If $\alpha=0$ we still have that $L_{c,c-1}=1$, which suffices for our purpose (the element $\Omega_{r,c-1}$ is non-zero with probability 1 when $r>\bar{r}$). Then, by Lemma 2 and the definition of $\mathbf{\Omega}$ for row indexes $r>\bar{r}$, it follows that $\mathbf{\Omega}\mathbf{L}^{{}^{\prime}}_{r,c}\neq 0$ with probability 1 if $c\neq 1$.
By the same reasoning, the matrix $\mathbf{L}\mathbf{\Omega}$ has zeros in the first row and nonzero entries $\mathbf{L}\mathbf{\Omega}_{r,c}$ for every column $c>\bar{r}$ with probability 1 if $r\neq 1$.
Third, we show that $\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}$ has non-zero elements for rows $r>\bar{r}+1$ and $c\neq 1$ with probability 1. Using matrix associativity,
$\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}_{r,c}=\sum_{i=1}^{T}L_{r,i}(\Omega L^{{}^{\prime}})_{i,c}$
Since $L_{1,i}=0$ for all $i$ and $(\Omega L^{{}^{\prime}})_{i,1}=0$ for all $i$, it follows that $\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}$ has zeros in the first row and first column. Moreover, for row indexes $r>\bar{r}$ we know that the elements $L_{r,r-1},L_{r,r-2},\dots L_{r,1}$ are nonzero if $\alpha\neq 0$. Then, by Lemma 2 and the definition of $\mathbf{\Omega}$ for row indexes $r>\bar{r}$, it follows that $\mathbf{L}\mathbf{\Omega}\mathbf{L}^{{}^{\prime}}_{r,c}\neq 0$ with probability 1 if $\alpha\neq 0$. If $\alpha=0$ then we require $r>\bar{r}+1$ as $L_{r,r-2},\dots L_{r,1}$ are all equal to zero, so only $L_{r,r-1}=1$, but if $r=\bar{r}+1$ then $(\Omega L^{{}^{\prime}})_{r-1,c}=(\Omega L^{{}^{\prime}})_{\bar{r},c}$ may be equal to zero (the results derived before for $\mathbf{\Omega}\mathbf{L}^{\prime}$ hold under $r>\bar{r}$). Then, in the most general case, if $r>\bar{r}+1$ and $c\neq 1$ then $\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}_{r,c}\neq 0$ with probability 1.
Recall that $J_{r,c}=(\mathbf{\Omega}\mathbf{L^{\prime}})_{r,c}+(\mathbf{L}\mathbf{\Omega})_{r,c}+(\alpha-\widetilde{\alpha})(\mathbf{L}\mathbf{\Omega}\mathbf{L^{\prime}})_{r,c}$. We have shown that the first element is non-zero with probability 1 if $r>\bar{r}$, the second element is non-zero with probability 1 if $c>\bar{r}$ and $(\mathbf{L}\mathbf{\Omega}\mathbf{L^{\prime}})_{r,c}$ is non-zero with probability 1 if $r>\bar{r}+1$. It is clear that in any case the event in which the summands cancel out can only occur in a set of factor realizations that occur with probability zero (Lemma 2 and Assumption 6). This completes the proof. $\square$
Definition 1.
We refer to diagonal exclusion minors as minors that do not include any diagonal terms of the original matrix. Formally, let $\mathbf{A}$ be a $T\times T$ matrix, and define $\mathbf{M}^{\mathbf{A},k}_{R,C}$ as a diagonal exclusion minor of $\mathbf{A}$ of dimension $k$, consisting of rows $R=\{r_{1},r_{2},\dots,r_{k}\}$ and columns $C=\{c_{1},c_{2},\dots,c_{k}\}$.
By definition, $R$ and $C$ must satisfy
  1. 1.
    $|R|=|C|=k\leq\frac{T}{2}$.
  2. 2.
    $R\cap C=\emptyset$.
The minors in the definition are squared matrices, thus $|R|=|C|=k$. They do not contain diagonal element, thus $R\cap C=\emptyset$. It follows that $|R|+|C|=2k\leq T$.
Now, we are ready to state and show the following Propositions.
Proposition 1.
The determinant of every diagonal exclusion minor of dimension $k$ of $\mathbf{O}$ can be written as the sum of the determinant of the corresponding diagonal exclusion minor for $\mathbf{\Omega}$ and a term expressed as $(\alpha-\widetilde{\alpha})\widetilde{J}\left(\widetilde{\alpha},\theta\right)$
In mathematical terms, what we will show is that given any diagonal exclusion minor of $\mathbf{O}$ of dimension $k$, say $\mathbf{M}^{\mathbf{O},k}_{R,C}$, we can write its determinant as the sum of the determinant of the diagonal exclusion minor of $\mathbf{\Omega}$ associated with the same rows and columns, say $\mathbf{M}^{\mathbf{\Omega},k}_{R,C}$, and a term that can be written as $(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$
$\det\left(\mathbf{M}^{\mathbf{O},k}_{R,C}\right)=\det\left(\mathbf{M}^{\mathbf{\Omega},k}_{R,C}\right)+(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ (10)
where $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ is a scalar that is a function of the true structural parameters $\theta$ and $\widetilde{\alpha}$. Note that the value of $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ will depend on which columns and rows were taken from the original matrix so in general, $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)\neq\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta\right)$ for any other valid set of rows and columns $R^{\prime}$ and $C^{\prime}$. To keep notation as simple as possible, we sometimes use $\widetilde{J}_{R,C}$ instead of $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ to represent the same object.
Proof: We can prove Proposition 1 by induction.
1. Base case $k=1$: Without loss of generality $R=r$ and $C=c$. Then, a diagonal exclusion minor of $\mathbf{O}$ of dimension 1, say $\mathbf{M}^{\mathbf{O},1}_{r,c}$, is simply the term $O_{r,c}$ (for $r\neq c$) and its determinant is trivially given by
$\det\left(\mathbf{M}^{\mathbf{O},1}_{r,c}\right)=O_{r,c}$
From Corollary 1, we know that
$O_{r,c}=\Omega_{r,c}+(\alpha-\widetilde{\alpha})J_{r,c}\quad r\neq c$
Since $\Omega_{r,c}$ is a scalar we have that $\Omega_{r,c}=\det\left(\mathbf{M}^{\mathbf{\Omega},1}_{r,c}\right)$. We can rewrite the preceding equation as:
$\det\left(\mathbf{M}^{\mathbf{O},1}_{r,c}\right)=\det\left(\mathbf{M}^{\mathbf{\Omega},1}_{r,c}\right)+(\alpha-\widetilde{\alpha})\widetilde{J}_{r,c}$ (11)
where $\widetilde{J}_{r,c}=J_{r,c}$ when $k=1$. Therefore, the base step in the induction holds.
2. Inductive hypothesis: Assume our proposition holds for diagonal exclusion minors of dimension $(k-1)\times(k-1)$. This implies that for any set of valid $R$ and $C$ we have that
$\det\left(\mathbf{M}^{\mathbf{O},k-1}_{R,C}\right)=\det\left(\mathbf{M}^{\mathbf{\Omega},k-1}_{R,C}\right)+(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ (12)
3. Inductive step for $k$: We prove the desired result using both Corollary 1, and invoking the Laplace expansion for determinants (Lemma 3).
Let $\mathbf{M}^{\mathbf{O},k}_{R,C}$ be a diagonal exclusion minor of $\mathbf{O}$ of dimension $k$ with rows $R=\{r_{1},r_{2},...,r_{k}\}$ and columns $C=\{c_{1},c_{2},...,c_{k}\}$ satisfying the usual conditions. Before proceeding, we need to to establish a correspondence between the indexes of rows $R$ and columns $C$ in the original matrices (either $\mathbf{O}$ or $\mathbf{\Omega}$) and their new indexes (i.e. $\{1,2,...,k\}$) in the $k\times k$ minors $\mathbf{M}^{\mathbf{O},k}_{R,C}$ and $\mathbf{M}^{\mathbf{\Omega},k}_{R,C}$. Without loss of generality, let $(i,j)$ denote the position of the $i$-th row in $R$ and the $j$-th column in $C$ within the minors $\mathbf{M}^{\mathbf{O},k}_{R,C}$ and $\mathbf{M}^{\mathbf{\Omega},k}_{R,C}$. This establishes that the $(i,j)$-th element of the minors $\mathbf{M}^{\mathbf{O},k}_{R,C}$ and $\mathbf{M}^{\mathbf{\Omega},k}_{R,C}$ are respectively given by $O_{r_{i},c_{j}}$ and $\Omega_{r_{i},c_{j}}$.
With this established, we can use Lemma 3 (fixing the row $r_{k}\in R$ in the original matrix) to write the determinant of $\mathbf{M}^{\mathbf{O},k}_{R,C}$ as
$\det\left(\mathbf{M}^{\mathbf{O},k}_{R,C}\right)=\sum_{j=1}^{k}(-1)^{k+j}O_{r_{k},c_{j}}\det\left(\mathbf{M}^{\mathbf{O},k-1}_{R-r_{k},C-c_{j}}\right)$ (13)
where $k$ and $j$ represent the indexes in minor $\mathbf{M}^{\mathbf{O},k}_{R,C}$ of row $r_{k}$ and columns $c_{j}$ of the matrix $\mathbf{O}$. The scalar $O_{r_{k},c_{j}}$ is the $(k,j)$-th element of the minor $\mathbf{M}^{\mathbf{O},k}_{R,C}$; the matrix $\mathbf{M}^{\mathbf{O},k-1}_{R-r_{k},C-c_{j}}$ is the minor of matrix $\mathbf{O}$ with dimension $k-1$ that includes rows $R-r_{k}$ and columns $C-c_{j}$.
Because $r_{k}\neq c_{j}$ for all $j$ (recall $R\cap C=\emptyset$), we can use Corollary 1 to write
$O_{r_{k},c_{j}}=\Omega_{r_{k},c_{j}}+(\alpha-\widetilde{\alpha})J_{r_{k},c_{j}}$
Furthermore, by the inductive hypothesis
$\det\left(\mathbf{M}^{\mathbf{O},k-1}_{R-r_{k},C-c_{j}}\right)=\det\left(\mathbf{M}^{\mathbf{\Omega},k-1}_{R-r_{k},C-c_{j}}\right)+(\alpha-\widetilde{\alpha})\widetilde{J}_{R-r_{k},C-c_{j}}$
Plugging these results into (13) we have that
$\det\left(\mathbf{M}^{\mathbf{O},k}_{R,C}\right)=\sum_{j=1}^{k}(-1)^{k+j}\left(\Omega_{r_{k},c_{j}}+(\alpha-\widetilde{\alpha})J_{r_{k},c_{j}}\right)\left[\det\left(\mathbf{M}^{\mathbf{\Omega},k-1}_{R-r_{k},C-c_{j}}\right)+(\alpha-\widetilde{\alpha})\widetilde{J}_{R-r_{k},C-c_{j}}\right]$
Distributing terms we obtain
$\begin{split}&\det\left(\mathbf{M}^{\mathbf{O},k}_{R,C}\right)=\sum_{j=1}^{k}(-1)^{k+j}\Omega_{r_{k},c_{j}}\det\left(\mathbf{M}^{\mathbf{\Omega},k-1}_{R-r_{k},C-c_{j}}\right)+\\ &(\alpha-\widetilde{\alpha})\sum_{j=1}^{k}(-1)^{k+j}\left[J_{r_{k},c_{j}}\det\left(\mathbf{M}^{\mathbf{\Omega},k-1}_{R-r_{k},C-c_{j}}\right)+\Omega_{r_{k},c_{j}}\widetilde{J}_{R-r_{k},C-c_{j}}+(\alpha-\widetilde{\alpha})J_{r_{k},c_{j}}\widetilde{J}_{R-r_{k},C-c_{j}}\right]\end{split}$
Then, using Lemma 3 we can rewrite the first term of the expression above as
$\sum_{j=1}^{k}(-1)^{k+j}\Omega_{r_{k},c_{j}}\det\left(\mathbf{M}^{\mathbf{\Omega},k-1}_{R-r_{k},C-c_{j}}\right)=\det\left(\mathbf{M}^{\mathbf{\Omega},k}_{R,C}\right).$
Next, define $\widetilde{J}_{R,C}$ as
$\widetilde{J}_{R,C}=\sum_{j=1}^{k}(-1)^{k+j}\left[J_{r_{k},c_{j}}\det\left(\mathbf{M}^{\mathbf{\Omega},k-1}_{R-r,C-c_{j}}\right)+\Omega_{r_{k},c_{j}}\widetilde{J}_{R-r_{k},C-c_{j}}+(\alpha-\widetilde{\alpha})J_{r_{k},c_{j}}\widetilde{J}_{R-r_{k},C-c_{j}}\right]$ (14)
This permits us to conclude that
$\det\left(\mathbf{M}^{\mathbf{O},k}_{R,C}\right)=\det\left(\mathbf{M}^{\mathbf{\Omega},k}_{R,C}\right)+(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}$
$\Box$
Proposition 2.
Consider any diagonal exclusion minor of $\mathbf{O}$ of dimension $k$. Then, $\widetilde{J}_{R,C}(\widetilde{\alpha},\theta)$ is a random polynomial in $\widetilde{\alpha}$ of order in between $\max\{0,k-\bar{r}\}$ and $12k-$ 1.
Proof: First, note from (14) that $\widetilde{J}_{R,C}$ is recursively defined, so once we know $\widetilde{J}_{R,C}$ for $|R|=|C|=1$ we can compute $\widetilde{J}_{R,C}$ for any arbitrary set of rows and columns.
When $k=1$, we have shown that $\widetilde{J}_{R,C}=J_{r,c}$. Now, recall the definition of $\mathbf{J}$ in Claim 2. $\mathbf{\Omega}$ and $\mathbf{L}$ do not depend on $\widetilde{\alpha}$, so $J_{r,c}$ is a polynomial of order 1 in $\widetilde{\alpha}$ if $r>\bar{r}+1$ and $c\neq 1$ or $c>\bar{r}+1$ and $r\neq 1$, and it is a polynomial of order 0 in $\widetilde{\alpha}$ if $r$ or $c$ are equal to 1 (see the discussion about $\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}$ in Corollary 2). Intermediate cases will depend on the values of the other parameters, but it will be a polynomial of either order 0 or 1, depending on whether $\mathbf{L}\mathbf{\Omega}\mathbf{L}^{{}^{\prime}}_{r,c}$ is equal or different from zero. Given the recursive structure of $\widetilde{J}_{R,C}$ we can conclude that for any value of $k$, $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ is a polynomial in $\widetilde{\alpha}$.
Note that the coefficients of this polynomial depend on the true data-generating process parameters $\theta$, and that by Assumptions 5 and 6, $\mathbf{F}$ is drawn from a continuous probability distribution. Then, we can conclude that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ is a random polynomial in $\widetilde{\alpha}$ and that the randomness in the coefficients is given by the factor realizations $\mathbf{F}$.
Having this in mind, we first show that the proposed upper bound for the order of the polynomial holds. Then, we show that the lower bound holds as well.
1. Proof for the upper bound: We will show that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ is a random polynomial of at most order $12k-$ 1 when $|R|=|C|=k$. The proof will be done by induction.
Base case $k=1$: We have shown that when $k=1$ $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ is a random polynomial in $\widetilde{\alpha}$ of at most order $11=2k-$ 1, so the step base holds.
Inductive hypothesis: Assume our proposition holds for diagonal exclusion minors of dimension $(k-1)\times(k-1)$. This implies that for any set of valid $R$ and $C$ we have that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ is a random polynomial in $\widetilde{\alpha}$ of at most order $12(k-1)-$ 1.
Inductive step: Recall the recursive definition of $\widetilde{J}_{R,C}$ derived in (14). By the inductive hypothesis $\widetilde{J}_{R-r_{k},C-c_{j}}$ is a random polynomial in $\widetilde{\alpha}$ of at most order $12(k-1)-$ 1, $(\alpha-\widetilde{\alpha})$ is a polynomial of order 1 in $\widetilde{\alpha}$, and $J_{r_{k},c_{j}}$ is also a random polynomial in $\widetilde{\alpha}$ of at most order 1 (already discussed). Therefore, the term $(\alpha-\widetilde{\alpha})J_{r_{k},c_{j}}\widetilde{J}_{R-r_{k},C-c_{j}}$ is a random polynomial in $\widetilde{\alpha}$ of at most order $12k-$ 1. Since the first two terms are of lower order ($\mathbf{\Omega}$ does not depend on $\widetilde{\alpha}$), we conclude that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ is a random polynomial in $\widetilde{\alpha}$ of at most order $12k-$ 1. $\square$
2. Proof for the lower bound: First, note that when $k\leq\bar{r}$ our statement trivially holds because $\max\{0,k-\bar{r}\}=0$ and $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ is a polynomial in $\widetilde{\alpha}$. Then, we are left to show that our statement holds for $k>\bar{r}$. We prove this by induction:
Base case $k=\bar{r}+1$: If $k=\bar{r}+1$, then it must be the case that $\exists\ c_{j}\in C$ such that $c_{j}>\bar{r}$. Thus, by Corollary 2 we know there exists a $c_{j}$ such that $J_{r_{k},c_{j}}\neq 0$ with probability 1. As by the inductive hypothesis $\widetilde{J}_{R-r_{k},C-c_{j}}$ is a random polynomial in $\widetilde{\alpha}$ of at least order 00 and $(\alpha-\widetilde{\alpha})$ is of order 1, then $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ should be at least of order $1=k-\bar{r}=\max\{k-\bar{r},0\}$, as desired.
Inductive hypothesis: Assume our proposition holds for diagonal exclusion minors of dimension $(k-1)\times(k-1)$, with $k-1>\bar{r}$. This implies that for any set of valid $R$ and $C$ we have that $\widetilde{J}_{R,C}$ is a random polynomial in $\widetilde{\alpha}\left(\widetilde{\alpha},\theta\right)$ of at least order $\max\{0,k-1-\bar{r}\}=k-1-\bar{r}$.
Inductive step: For any $k>\bar{r}$ we know that there exists a $c_{j}$ such that $J_{r_{k},c_{j}}\neq 0$ with probability 1 (Corollary 2). By the inductive hypothesis $\widetilde{J}_{R-r_{k},C-c_{j}}$ is a random polynomial of at least order $k-1-\bar{r}$, $(\alpha-\widetilde{\alpha})$ is of order 1, which implies that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ should be a random polynomial of at least order $k-\bar{r}=\max\{0,k-\bar{r}\}$, as desired. $\square$
Propositions 1 and 2 provide the essential results for proving Theorem 1. However, before we proceed with the proof, we need to establish one additional result. Consider the following claim:
Claim 3.
Let $\mathbf{M}^{\mathbf{\Omega},k}_{R,C}$ be a diagonal exclusion minor of $\mathbf{\Omega}$ and let $\mathbf{M}^{\mathbf{F\Psi F^{\prime}},k}_{R,C}$ be the corresponding diagonal exclusion minor of $\mathbf{F\Psi F^{\prime}}$. Then $\mathbf{M}^{\mathbf{\Omega},k}_{R,C}=\mathbf{M}^{\mathbf{F\Psi F^{\prime}},k}_{R,C}$.
Proof: Recall
$\mathbf{\Omega}=\mathbf{F\Psi F^{\prime}+D}$
where $\mathbf{D}$ is a diagonal matrix. Then, one can see that all the non-diagonal terms of $\mathbf{\Omega}$ are equal to the non-diagonal terms of $\mathbf{F\Psi F^{\prime}}$. Thus, as all the terms are equal it is obvious that for every diagonal exclusion $k\times k$ minor we have that
$\mathbf{M}^{\mathbf{\Omega},k}_{R,C}=\mathbf{M}^{\mathbf{F\Psi F^{\prime}},k}_{R,C}$
$\square$
From this result, we can derive the following corollary
Corollary 3.
The determinant of every diagonal exclusion minor of $\mathbf{\Omega}$ of dimension $k>\bar{r}$ is equal to 0.
Proof: This is a direct implication from the previous Claim. First, note that from Claim 3, the determinant of every diagonal exclusion minor of $\mathbf{\Omega}$ is equal to the determinant of the corresponding non-diagonal minor of $\mathbf{F\Psi F^{\prime}}$
$\det\left(\mathbf{M}^{\mathbf{\Omega},k}_{R,C}\right)=\det\left(\mathbf{M}^{\mathbf{F\Psi F^{\prime}},k}_{R,C}\right)$
Now, let $k>\bar{r}$. Then, as $\mathbf{F\Psi F^{\prime}}$ is of rank $\bar{r}$ Lemma 1 implies that
$\det\left(\mathbf{M}^{\mathbf{\Omega},k}_{R,C}\right)=0$
proving the corollary. $\Box$
4.3 Proof of Theorem 1
In this subsection, we build on the intermediate results derived in subsection 4.2 to prove Theorem 1. Specifically, we demonstrate that (5) can only hold if $\widetilde{\alpha}=\alpha$, except for a set of factor realizations that occur with probability zero. We begin by outlining the structure of the proof, then proceed with a detailed argument for the case of a single factor to illustrate the reasoning. Finally, we generalize the proof to the case of $\bar{r}$ factors.
Proof of Theorem 1: Pick any diagonal exclusion minor of $\mathbf{O}$ of dimension $\bar{r}+1$, say $\mathbf{M}^{\mathbf{O},\bar{r}+1}_{R,C}$. By Claim 1 the determinant of this minor must satisfy
$\det\left(\mathbf{M}^{\mathbf{O},\bar{r}+1}_{R,C}\right)=0$
By Proposition 1 we know that
$\det\left(\mathbf{M}^{\mathbf{O},\bar{r}+1}_{R,C}\right)=\det\left(\mathbf{M}^{\mathbf{\Omega},\bar{r}+1}_{R,C}\right)+(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)=0$
Moreover, by Corollary 3 we have that $\det\left(\mathbf{M}^{\mathbf{\Omega},\bar{r}+1}_{R,C}\right)=0$. Then:
$\det\left(\mathbf{M}^{\mathbf{O},\bar{r}+1}_{R,C}\right)=(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)=0$ (15)
By Assumption 3 we know that $\mathbf{O}$ has at least $3$ different diagonal exclusion minors. Then, consider any other diagonal exclusion minor of $\mathbf{O}$, say $\mathbf{M}^{\mathbf{O},\bar{r}+1}_{R^{\prime},C^{\prime}}$. Using the same logic, we can see that:
$\det\left(\mathbf{M}^{\mathbf{O},\bar{r}+1}_{R^{\prime},C^{\prime}}\right)=(\alpha-\widetilde{\alpha})\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta\right)=0$
Then, it must be the case that either $\widetilde{\alpha}=\alpha$ or $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)=\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta\right)=0$ in order for (5) to hold (Claim 1). However, we will show that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)=\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta\right)=0$ is only possible in a set of probability zero.
Consider the case of a single factor $\bar{r}=1$
Consider the following $2\times 2$ diagonal exclusion minor of $\mathbf{O}$:
$\mathbf{M}_{(1,2),(3,4)}^{\mathbf{O},2}=\begin{bmatrix}O_{1,3}&O_{1,4}\\ O_{2,3}&O_{2,4}\end{bmatrix}$
See Appendix A for the explicit functional form of each $O_{i,j}$. Combining Proposition 1, Corollary 3, and Claim 1 we get that:
$\det\left(\mathbf{M}_{(1,2),(3,4)}^{\mathbf{O},2}\right)=\left(\alpha-\widetilde{\alpha}\right)\Big{(}\left[\Psi f_{3}(d_{2}-\alpha d_{1}f_{2})\right]\widetilde{\alpha}-\Psi f_{4}(d_{2}-\alpha d_{1}f_{2})\Big{)}=0$ (16)
where in this case $\widetilde{J}_{(1,2),(3,4)}\left(\widetilde{\alpha},\theta\right)$ is a random polynomial in $\widetilde{\alpha}$ of order 1. Note that one of the following four alternatives must be true in order for (16) to be satisfied:
Case 1.
$\Psi=0$
Case 2.
$f_{2}=\frac{d_{2}}{d_{1}\alpha}$
Case 3.
$\widetilde{\alpha}=\frac{f_{4}}{f_{3}}$
Case 4.
$\widetilde{\alpha}=\alpha$
Cases 1 and 2 correspond to the case in which $\widetilde{J}_{(1,2),(3,4)}\left(\widetilde{\alpha},\theta\right)$ is the 0-polynomial, so that all the coefficients are trivially equal to 0. Case 3 corresponds to the case in which we can found a root $\widetilde{\alpha}^{\star}$ that solves $\widetilde{J}_{(1,2),(3,4)}(\widetilde{\alpha}^{\star},\theta)=0$. However, once $\widetilde{\alpha}$ is pinned-down in this way, we have no other free parameter, so the same root must satisfy all the 0-determinant conditions (Claim 1) associated with the remaining diagonal exclusion minors of $\mathbf{O}$. Case 4 corresponds to the case where global identification is achieved.
We will show that Cases 1-3 are only possible in a set of factor realizations that occur with probability zero. Thus, with probability 1, Case 4 must hold so $\alpha$ is almost surely globally identified.
Case 1: $\Psi=0$
This case is obvious. We assumed $\Psi>0$, so we can instantly rule this out.
Case 2: $f_{2}=\frac{d_{2}}{d_{1}\alpha}$
From Assumptions 5 and 6, we know that $d_{1},d_{2},\alpha$ are fixed population parameters, whereas $f_{2}$ is drawn from a continuous probability distribution. Thus, we can invoke Lemma 2 and ensure that $P\left(f_{2}=\frac{d_{2}}{d_{1}\alpha}\right)=0$. Then, Case 2 can only occur in a set of probability 0.
Case 3: $\widetilde{\alpha}=\frac{f_{4}}{f_{3}}$
Setting $\widetilde{\alpha}^{\star}=\frac{f_{4}}{f_{3}}$ so that $\widetilde{J}_{(1,2),(3,4)}(\widetilde{\alpha}^{\star},\theta)=0$ leaves us with no other free parameter. Therefore, in order for Case 3 to be possible, it must be the case that this same $\widetilde{\alpha}^{\star}$ simultaneously make all the $\widetilde{J}_{R^{\prime},C^{\prime}}(\widetilde{\alpha}^{\star},\theta)$ equal to 0 for any other valid set of rows and columns (in particular, it requires $\widetilde{J}_{(2,3),(1,4)}(\widetilde{\alpha}^{\star},\theta)=0$). To see why this will not be the case, consider the following $2\times 2$ diagonal exclusion minor of $\mathbf{O}$:
$\mathbf{M}_{(2,3),(1,4)}^{\mathbf{O},2}=\begin{bmatrix}O_{2,1}&O_{2,4}\\ O_{3,1}&O_{3,4}\end{bmatrix}$
Following the same procedure as for (15),
$\det\left(\mathbf{M}_{(2,3),(1,4)}^{\mathbf{O},2}\right)=(\alpha-\widetilde{\alpha})\widetilde{J}_{(2,3),(1,4)}\left(\widetilde{\alpha},\theta\right)=0$
Algebra calculation shows that $\widetilde{J}_{(2,3),(1,4)}\left(\widetilde{\alpha},\theta\right)$ is a polynomial of order 2 in $\widetilde{\alpha}$:
$\widetilde{J}_{(2,3),(1,4)}\left(\widetilde{\alpha},\theta\right)=a\widetilde{\alpha}^{2}+b\widetilde{\alpha}+c$ (17)
with the coefficients given by:
$a$ $=\alpha d_{1}d_{2}+\alpha d_{2}\Psi+d_{1}f_{2}f_{3}\Psi$
$b$ $=-(d_{1}d_{3}+d_{3}\Psi+\alpha^{2}d_{1}d_{2}+\alpha^{2}d_{2}\Psi+d_{1}f_{3}^{2}\Psi+\alpha^{2}d_{1}f_{2}^{2}\Psi+d_{1}f_{2}f_{4}\Psi+2\alpha d_{1}f_{2}f_{3}\Psi)$
$c$ $=\alpha d_{1}d_{3}+\alpha d_{3}\Psi+d_{3}f_{2}\Psi+\alpha d_{1}f_{3}^{2}\Psi-\alpha d_{2}f_{3}\Psi+d_{1}f_{3}f_{4}\Psi+\alpha^{2}d_{1}f_{2}f_{3}\Psi$
Recall that $\widetilde{J}_{(2,3),(1,4)}\left(\widetilde{\alpha},\theta\right)=0$ must hold in order for Case 3 to be plausible. There are two ways in which $\widetilde{J}_{(2,3),(1,4)}\left(\widetilde{\alpha},\theta\right)$ can be 0: either all the coefficients are equal to 0, so that $\widetilde{J}_{(2,3),(1,4)}\left(\widetilde{\alpha},\theta\right)$ is the 0-polynomial; or the coefficients of the polynomial are such that real roots exist and at least one of them coincide with $\widetilde{\alpha}^{\star}=\frac{f_{4}}{f_{3}}$.
We can see that Assumption 6 implies that $a=b=c=0$ can only occur on a set with probability 0, as it involves the factor realizations lying on a particular low-dimensional manifold (see Lemma 2). That is, the first case has zero probability of occurring.
We next show that the second case also has zero probability of occurring. To see this, note that $\widetilde{J}_{(2,3),(1,4)}$ is a polynomial of order 2 in $\widetilde{\alpha}$. Its roots are:
$\widetilde{\alpha}=\frac{-b\pm\sqrt{b^{2}-4ac}}{2a}$
Conditional on the roots being real, it must be the case that either
$\frac{f_{4}}{f_{3}}=\frac{-b+\sqrt{b^{2}-4ac}}{2a}\quad\text{or}\quad\frac{f_{4}}{f_{3}}=\frac{-b-\sqrt{b^{2}-4ac}}{2a}$
Examining the expressions of $a$, $b$, and $c$ and making use of Lemma 2 and Assumption 6, we conclude that these events can never occur with positive probability (the factors $f_{2},f_{3},f_{4}$ must lie exactly in a lower dimensional manifold). Hence $\widetilde{\alpha}=\frac{f_{4}}{f_{3}}$ is only possible on a set of zero probability.
Having ruled out Cases 1-3, we are left with Case 4, which is the only option that can happen with positive probability when $\bar{r}=1$, so it must be the case that $\widetilde{\alpha}=\alpha$ almost surely in order for (5) to be satisfied. We have thus established global identification for $\bar{r}=1$. $\square$
Consider the general case with $\bar{r}$ factors
Start from (15)
$\det\left(\mathbf{M}^{\mathbf{O},\bar{r}+1}_{R,C}\right)=(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)=0$
Our goal is to show that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)=0$ can only occur with probability zero, implying that $\widetilde{\alpha}=\alpha$ must hold almost surely. Recall from Proposition 2 that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ is a random polynomial in $\widetilde{\alpha}$ of degree in between $1$ and $12\bar{r}+$ 1, given by:
$\widetilde{J}_{R,C}=\sum_{j=1}^{\bar{r}+1}(-1)^{\bar{r}+1+j}\left[J_{r_{\bar{r}+1},c_{j}}\det\left(\mathbf{M}^{\mathbf{\Omega},\bar{r}}_{R-r_{\bar{r}+1},C-c_{j}}\right)+\Omega_{r_{\bar{r}+1},c_{j}}\widetilde{J}_{R-r_{\bar{r}+1},C-c_{j}}+(\alpha-\widetilde{\alpha})J_{r_{\bar{r}+1},c_{j}}\widetilde{J}_{R-r_{\bar{r}+1},C-c_{j}}\right]$
By Assumptions 5 and 6, $\alpha,\mathbf{D}\ \text{and}\ \mathbf{\Psi}$ are deterministic population parameters, and $\mathbf{F}$ is drawn from a continuous distribution. The condition $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)=0$ can occur in two possible ways: either all of the coefficients of this random polynomial are equal to 0, or by setting $\widetilde{\alpha}$ equal to one of the roots of this polynomial $\widetilde{J}_{R,C}(\widetilde{\alpha}^{\star},\theta)=0$ (by the fundamental theorem of algebra we know that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ has at most $12\bar{r}+$ 1 real roots).
Consider the first case. Let $\kappa\in\{1,2,\dots,2\bar{r}+1\}$ be the degree of the random polynomial $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$. In order for all the coefficients to be equal to 0, the factor realizations must be such that the coefficients $a_{0}(\theta),a_{1}(\theta),\dots,a_{\kappa}(\theta)$ are all simultaneously equal to 0. By Lemma 2, this is only possible in a set of probability 0.
Now consider the second case. If $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ has no real roots, then it is clear that it must be that $\widetilde{\alpha}=\alpha$ in order for (15) to hold, so the proof would be over. Now, assume $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ has at least one real root. Pick any of these roots, say $\widetilde{\alpha}^{\star}_{R,C}(\theta)$. By definition, $\widetilde{\alpha}^{\star}_{R,C}(\theta)$ solves:
$\widetilde{J}_{R,C}\left(\widetilde{\alpha}^{\star}_{R,C},(\theta),\theta\right)=0$
Note that $\widetilde{\alpha}^{\star}_{R,C}(\theta)$ is a function of the other true data-generating process parameters and can be obtained from the above condition. Moreover, it also depends on the rows $R$ and columns $C$ chosen as in general the random polynomials $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ and $\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta\right)$ are different (except for a set of factor realizations of probability zero).
As we used the diagonal exclusion minor associated with rows and columns $R,C$ to pin down $\widetilde{\alpha}$ we have no other free parameter, so this same $\widetilde{\alpha}^{\star}_{R,C}(\theta)$ must satisfy:
$\widetilde{J}_{R^{\prime},C^{\prime}}(\widetilde{\alpha}^{\star}_{R,C}(\theta),\theta)=0$
for any other valid set of $R^{\prime},C^{\prime}$. This can only occur in two different ways: either $\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta\right)$ is the 0-polynomial, or the random polynomial $\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta\right)$ shares at least one root with $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$.
As we have previously discussed, the case in which $\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta\right)$ is the 0-polynomial is an event that occurs with probability zero. Therefore, consider the second case. We know that the coefficients of $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ and $\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta\right)$ can be exactly equal only in a set of probability zero, so the polynomials are almost surely different. Then, the case in which they share the root $\widetilde{\alpha}^{\star}_{R,C}(\theta)$ can be expressed as:
$\begin{split}\widetilde{J}_{R^{\prime},C^{\prime}}&=\sum_{j=1}^{\bar{r}+1}(-1)^{\bar{r}+1+j^{{}^{\prime}}}\Bigg{[}J_{r_{\bar{r}+1}^{{}^{\prime}},c_{j}^{{}^{\prime}}}\det\left(\mathbf{M}^{\mathbf{\Omega},\bar{r}}_{R^{\prime}-r_{\bar{r}+1}^{{}^{\prime}},C^{\prime}-c_{j}^{{}^{\prime}}}\right)+\Omega_{r_{\bar{r}+1}^{{}^{\prime}},c_{j}^{{}^{\prime}}}\widetilde{J}_{R^{\prime}-r_{\bar{r}+1}^{{}^{\prime}},C^{\prime}-c_{j}^{{}^{\prime}}}\\ &\quad+\left(\alpha-\widetilde{\alpha}^{\star}_{R,C}(\theta)\right)J_{r_{\bar{r}+1}^{{}^{\prime}},c_{j}^{{}^{\prime}}}\widetilde{J}_{R^{\prime}-r_{\bar{r}+1}^{{}^{\prime}},C^{\prime}-c_{j}^{{}^{\prime}}}\Bigg{]}=0\end{split}$ (18)
Note that the above condition depends only on the structural parameters of the model. Moreover, it defines a particular low-dimensional manifold in which the random factors $\mathbf{F}$ must lie in order for both polynomials to share the root $\widetilde{\alpha}^{\star}_{R,C}(\theta)$. However, by Assumption 6 and Lemma 2, the event in which the random factors lie in this particular low-dimensional manifold occurs with probability zero.
Then, the case $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)=\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta\right)=0$ can only occur in a set of probability 0. Thus, it must be the case that $\widetilde{\alpha}=\alpha$ in order for (5) to be satisfied. This implies that $\alpha$ is almost surely globally identified. $\square$
4.4 Global identification of $\mathbf{F}$, $\mathbf{\Psi}$ and $\mathbf{D}$
We have shown that $\alpha$ is almost surely globally identified, meaning that (5) is satisfied if and only if $\widetilde{\alpha}=\alpha$ with probability 1. We can then conclude that all the parameters in our model are globally identified. This is stated in Theorem 2. We next provide a proof.
Proof of Theorem 2: As $\alpha$ is almost surely globally identified, then $\widetilde{\alpha}=\alpha$ with probability 1 and $\mathbf{Q}=\mathbf{I_{T}}$. Then, we are left to show:
$\mathbf{F}\mathbf{\Psi}\mathbf{F^{\prime}}+\mathbf{D}=\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}+\mathbf{\widetilde{D}}$
if and only if $\mathbf{\widetilde{D}}=\mathbf{D}$, and $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}=\mathbf{F}\Psi\mathbf{F^{\prime}}$, where $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}$ must be a $T\times T$ symmetric PSD matrix of rank $\bar{r}$ and $\mathbf{\widetilde{D}}$ a $T\times T$ PD diagonal matrix.
Under the Assumptions of our model, this is the case shown by Anderson and Rubin (1956), so all the rest of the parameters are almost surely globally identified as well. $\square$
5 Individual fixed-effects
So far, we have established that the model parameters are globally identified under the assumption that all factors in $\mathbf{F}$ follow a continuous probability distribution (Assumption 6). While this assumption is quite general and encompasses many relevant applications, it notably excludes one important case: individual fixed-effects.
$y_{it}=\alpha y_{it-1}+\delta_{t}+\gamma_{i}+\lambda_{i}^{{}^{\prime}}f_{t}+\varepsilon_{it}$ (19)
where $\gamma_{i}$ denotes the individual fixed effect for observation $i$, and the rest of the parameters are defined as in (1). It is important to note that we will not estimate the fixed effects $\gamma_{i}$, doing so will cause the incidental parameters problem. We simply treat $\gamma_{i}$ as the factor loadings of a special factor, which taking on the value of 1 for all $t$. The number of factors in this model is $\bar{r}=dim(f_{t})+1$.
Using the same procedure as in Section 2.1, we project the first observation $y_{i1}$ onto the vector $[1,\gamma_{i},\lambda_{i}]$, yielding the following expression: $y_{i1}=\delta_{1}^{*}+\gamma_{i}f_{\gamma}^{*}+\lambda_{i}^{\prime}f_{1}^{*}+\varepsilon_{i1}^{*}$, where $(\delta_{1}^{*},f_{\gamma}^{*},f_{1}^{*})$ are the projection coefficients and $\varepsilon_{i1}^{*}$ is the projection residual. For notational simplicity, we drop the asterisks from the coefficients and from the residual, as these are free parameters, and we do not require the error terms $\varepsilon_{it}$ to have identical distributions across $t$. Therefore, for $t=1$, we can rewrite the equation as:
$y_{i1}=\delta_{1}+\gamma_{i}f_{\gamma}+\lambda_{i}^{\prime}f_{1}+\varepsilon_{i1}.$
Combining this with equation (19) for $t\geq 2$ and stacking the observations over $t$, we obtain:
$\mathbf{B}y_{i}=\delta+\mathbf{F}\eta_{i}+\varepsilon_{i}$
where $\mathbf{B}$, $y_{i}$, $\delta$, and $\varepsilon_{i}$ are all defined earlier, but
$\mathbf{F}=\begin{bmatrix}f_{\gamma}&f_{1}^{{}^{\prime}}\\ 1&f_{2}^{{}^{\prime}}\\ \vdots&\vdots\\ 1&f_{T}^{{}^{\prime}}\end{bmatrix},\quad\text{ and }\eta_{i}=\begin{bmatrix}\gamma_{i}\\ \lambda_{i}\end{bmatrix}$
Note that the first element in the first column of $\mathbf{F}$ is $f_{\gamma}$, not 1. Importantly, the first column of $\mathbf{F}$ consists of many 1’s, which are fixed constants and follow degenerate distributions. As a result, the factor matrix does not satisfy the conditions of Assumption 6. However, since this assumption is a sufficient but not necessary condition, we will show that the model remains globally identified. The key insight is that, as long as there is sufficient variation in some of the factors, identification can still be achieved.
For simplicity, we provide a proof for the global identification of $\alpha$ in the presence of individual fixed-effects and a single unrestricted factor ($dim(f_{t})=1$), and $f_{t}$ is drawn from a continuous distribution. In this case, $\bar{r}=2$. However, the proof can be extended to more general cases encompassing multiple factors.
Note that the presence of individual fixed-effects imposes structure on $\mathbf{F}$, so we require less restrictions to eliminate rotational indeterminacy. We only need to impose two restrictions on $\mathbf{F}$ to prevent indeterminacy (see Appendix B for detailed explanation). We shall impose the normalization in the following way, referred to as tail normalization:
$\mathbf{F}=\begin{bmatrix}f_{\gamma}&f_{1}\\ 1&f_{2}\\ \vdots&\vdots\\ 1&f_{T-2}\\ 1&0\\ 1&1\end{bmatrix}$
This normalization restricts the last two realizations of the factor $f$ to be 0 and 1 respectively. This is purely for convenience in term of exposition. Normalizing other entries (except the first row) has the same effect. Under this normalization, $\mathbf{F}A=\mathbf{F}$, if and only if $A=I_{2}$, thus eliminating rotational indeterminacy.
Before showing that $\alpha$ can be almost surely identified in this context we shall explicitly state the identification assumptions:
Assumption 1’.
There is one factor associated with the individual fixed-effects and an unrestricted factor ($\bar{r}=2$)
Assumption 2’.
$T\geq 6$.
Assumption 3’.
$\eta_{i}$ are either random or fixed constants with sample covariance matrix $\bm{\Psi}_{N}$ satisfying $\bm{\Psi}_{N}\to\bm{\Psi}$ with $\bm{\Psi}$ being positive definite, as $N\to\infty$.
Assumption 4’.
$\alpha,\mathbf{\Psi},d_{1},d_{2},\dots,d_{T}$ are fixed population parameters.
Assumption 5’.
$f_{\gamma},f_{1},f_{2},\dots f_{T-2}$ are drawn from a continuous probability distribution.
Assumption 6’.
$d_{t}\neq\alpha d_{t-1}$ when $\alpha\neq 1$.
Assumptions 1’-5’ are an adaptation of the ones outlined in Section 3 for the particular case we are studying. While we conjecture that $f_{\gamma}$ drawn from a continuous distribution is not strictly needed to attain identification, it substantially simplifies the proof exposition.
Assumption 6’ is imposed because our strategy is not able to show identification in that particular case (it will become clearer why during the proof). We argue that this does not undermine the generality of our results for two main reasons. First, the case in which $d_{t}=\alpha d_{t-1}$ with $\alpha\neq 1$ is a restriction on the values of the parameters that have 0-lebesgue measure in the parameter space. Second, and more importantly, it is hard to think any true data generating process such that the variance of the idiosyncratic errors should evolve as $d_{t}=\alpha d_{t-1}$ when $\alpha\neq 1$, so this is a case we believe is never observed in practice. Note that the case in which $\alpha=1$ and $d_{t}=\alpha d_{t-1}$ is identified. This implies that if our process exhibits an unit root and homoskedastic variance we can almost surely identify $\alpha$, which is a case that should have more empirical relevance.
Under this assumptions, we can show that $\alpha$ is almost surely globally identified:
Theorem 3.
Let $y_{it}$ be generated from (19). Then, under Assumptions 1’-6’, $\alpha$ is almost surely globally identified.
Proof: The proof follows in a similar way as in the previous section, as most of the results derived did not depend on the distributional assumption imposed on $\mathbf{F}$. For instance, the definition of $\mathbf{O}$ is the same as before (for the case of $\bar{r}=2$) and Claim 1 still holds, so for any $3\times 3$ diagonal exclusion minor of $\mathbf{O}$ we should have:
$\det\left(\mathbf{M}^{\mathbf{O},3}_{R,C}\right)=0$
for any set of valid rows and columns. Moreover, Proposition 1, Claims 2,3, and Corollary 3 did not depend on the distributional assumptions imposed on $\mathbf{F}$, so we know that:
$\det\left(\mathbf{M}^{\mathbf{O},3}_{R,C}\right)=(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)=0$
We will show that $\alpha$ is almost surely identified when $T=6$. This is a sufficient condition for identification for any bigger $T$, as it is clear that having a higher number of observations can only help identification, not make it worse.
Let $T=6$ and $R=(1,2,3)$, $C=(4,5,6)$. The determinant of the diagonal exclusion minor associated with these rows and columns is given by:
$\det\left(\mathbf{M}^{\mathbf{O},3}_{(1,2,3),(4,5,6)}\right)$ $=(\alpha-\widetilde{\alpha})\left(\Psi_{1,1}\Psi_{2,2}-\Psi_{1,2}^{2}\right)$ (20)
×$\quad\times\left(d_{3}f_{1}-\alpha d_{2}f_{1}-d_{3}f_{2}f_{\gamma}+\alpha^{2}d_{1}f_{2}-\alpha^{2}d_{1}f_{3}+\alpha d_{2}f_{3}f_{\gamma}\right)$
×$\quad\times(\widetilde{\alpha}f_{4}+1)(\widetilde{\alpha}-1)=0$
where, in this case, $\widetilde{J}_{(1,2,3),(4,5,6)}\left(\widetilde{\alpha},\theta\right)$ is a random polynomial in $\widetilde{\alpha}$ of order 2. The proof is similar to the case of $\bar{r}=1$ in Section 4.3. Note that for equation (20) to hold, one of the following five alternatives must be true:
Case 1’.
$\Psi_{1,2}^{2}=\Psi_{1,1}\Psi_{2,2}$
Case 2’.
$(d_{3}f_{1}-\alpha d_{2}f_{1}-d_{3}f_{2}f_{\gamma}+\alpha^{2}d_{1}f_{2}-\alpha^{2}d_{1}f_{3}+\alpha d_{2}f_{3}f_{\gamma})=0$
Case 3’.
$\widetilde{\alpha}=-\frac{1}{f_{4}}$
Case 4’.
$\widetilde{\alpha}=1$
Case 5’.
$\widetilde{\alpha}=\alpha$
Cases 1’ and 2’ correspond to the case in which $\widetilde{J}_{(1,2,3),(4,5,6)}$ is the 0-polynomial, so that all the coefficients are trivially equal to 0. Cases 3’ and 4’ correspond to the case in which we can found a root $\widetilde{\alpha}^{\star}$ that solves $\widetilde{J}_{(1,2,3),(4,5,6)}(\widetilde{\alpha}^{\star},\theta)=0$. However, once $\widetilde{\alpha}$ is pinned-down in this way, we have no other free parameter, so the same root must satisfy all the 0-determinant conditions (Claim 1) associated with the remaining diagonal exclusion minors of $\mathbf{O}$. Case 5’ corresponds to the case where global identification is achieved.
We show that Cases 1’-4’ are only possible in a set of factor realizations that occur with probability zero. Thus, with probability 1, Case 5’ must hold so $\alpha$ is almost surely globally identified.
Case 1’: $\Psi_{1,2}^{2}=\Psi_{1,1}\Psi_{2,2}$
This case implies that $\det(\mathbf{\Psi})=0$, which is analogous to the case where $\Psi=0$ in subsection 4.3. However, this case is excluded because we assume that $\mathbf{\Psi}$ is positive definite.
Case 2’: $d_{3}f_{1}-\alpha d_{2}f_{1}-d_{3}f_{2}f_{\gamma}+\alpha^{2}d_{1}f_{2}-\alpha^{2}d_{1}f_{3}+\alpha d_{2}f_{3}f_{\gamma}=0$
From Assumptions 4’ and 5’, we know that $d_{1},d_{2},d_{3},\alpha$ are fixed population parameters, while $f_{1},f_{2},f_{3},f_{\gamma}$ are drawn from a continuous probability distribution. Therefore, we can apply Lemma 2 to conclude that Case 2’ occurs only with probability zero.
Cases 3’ and 4’: $\widetilde{\alpha}=-\frac{1}{f_{4}}$ and $\widetilde{\alpha}=1$
We demonstrate that the case $\widetilde{\alpha}=1$ can never satisfy 5 with positive probability. The proof for $\widetilde{\alpha}=-\frac{1}{f_{4}}$ is analogous and thus omitted.
If we set $\widetilde{\alpha}=1$ in order to satisfy Equation 20, then we have no other free parameters. Thus, this same $\widetilde{\alpha}$ must satisfy all the 0-determinant conditions (Claim 1). However, this cannot occur with positive probability.
To see this, consider the determinant of the diagonal exclusion minor associated with rows $R=(1,5,6)$ and columns $C=(2,3,4)$. By Proposition 1 and Corollary 3, we know it can be written as:
$\det\left(\mathbf{M}^{\mathbf{O},3}_{(1,5,6),(2,3,4)}\right)=(\alpha-\widetilde{\alpha})\widetilde{J}_{(1,5,6),(2,3,4)}\left(\widetilde{\alpha},\theta\right)$
where $\widetilde{J}_{(1,5,6),(2,3,4)}\left(\widetilde{\alpha},\theta\right)$ is a random polynomial in $\widetilde{\alpha}$ of order 4. Case 4’ is plausible if both $\widetilde{J}_{(1,2,3),(4,5,6)}\left(\widetilde{\alpha},\theta\right)$ and $\widetilde{J}_{(1,5,6),(2,3,4)}\left(\widetilde{\alpha},\theta\right)$ share the root $\widetilde{\alpha}=1$. Note that this is equivalent to stating that $\widetilde{J}_{(1,5,6),(2,3,4)}\left(\widetilde{\alpha},\theta\right)|_{\widetilde{\alpha}=1}=0$. Evaluating $\widetilde{J}_{(1,5,6),(2,3,4)}\left(\widetilde{\alpha},\theta\right)$ at $\widetilde{\alpha}=1$ yields:
$\widetilde{J}_{(1,5,6),(2,3,4)}\left(\widetilde{\alpha},\theta\right)|_{\widetilde{\alpha}=1}$ $=\left(d_{4}f_{2}-d_{4}f_{3}-\alpha d_{3}f_{2}+\alpha d_{3}f_{4}+\alpha^{2}d_{2}f_{3}-\alpha^{2}d_{2}f_{4}\right)$
×$\quad\times\left(f_{\gamma}^{2}\Psi_{1,2}^{2}-f_{\gamma}\Psi_{1,2}^{2}-d_{1}\Psi_{2,2}-f_{\gamma}^{2}\Psi_{1,1}\Psi_{2,2}\right.$
$\quad\left.-\alpha f_{\gamma}^{2}\Psi_{1,2}^{2}+\alpha d_{1}\Psi_{2,2}+f_{\gamma}\Psi_{1,1}\Psi_{2,2}+\alpha f_{\gamma}^{2}\Psi_{1,1}\Psi_{2,2}\right)$
Thus, Case 4’ is plausible if either term is equal to zero. However, this can never occur with positive probability. Let’s begin with the first term. Using Assumption 6’, the case in which this term equals zero can be rewritten as:
$f_{2}=\frac{(d_{4}-\alpha^{2}d_{2})f_{3}+(\alpha^{2}d_{2}-\alpha d_{3})f_{4}}{d_{4}-\alpha d_{3}}$
This equation defines a plane on which the three-dimensional random variable $(f_{2},f_{3},f_{4})$ must lie. Since this is a probability zero event (Lemma 2), it cannot occur with positive probability. Now, consider the second term. It can be expressed as a second-order polynomial in $f_{\gamma}$:
$\mathcal{P}(f_{\gamma})=af_{\gamma}^{2}+bf_{\gamma}+c$
where
$a$ $=(\alpha-1)\left(\Psi_{1,1}\Psi_{2,2}-\Psi_{1,2}^{2}\right)$
$b$ $=\Psi_{1,1}\Psi_{2,2}-\Psi_{1,2}^{2}$
$c$ $=(\alpha-1)d_{1}\Psi_{2,2}$
Setting the second term equal to zero is equivalent to finding the roots of this polynomial. First, note that $\mathcal{P}$ can never be the zero polynomial since $b>0$, which holds because $\mathbf{\Psi}$ is positive definite by assumption. Then, by the fundamental theorem of algebra, there are at most two values of $f_{\gamma}$ that satisfy $\mathcal{P}(f_{\gamma})=0$. Since $f_{\gamma}$ is assumed to be drawn from a continuous distribution, the event $\mathcal{P}(f_{\gamma})=0$ has zero probability, because these two roots have zero Lebesgue measure in the support of $f_{\gamma}$.
Thus, we can conclude that almost surely, Case 4’ is not plausible. A similar argument leads to the conclusion that Case 3’ can only occur with probability zero.
We have shown that Case 5’ is the only alternative that makes $\det\left(\mathbf{M}^{\mathbf{O},3}_{(1,2,3),(4,5,6)}\right)=0$ with positive probability. Thus, $\alpha$ is almost surely globally identified in the presence of individual fixed effects.
Remark. We have proven identification for the case when $T=6$. Larger $T$ only strengthens the identification. One way to see this is by applying the same normalization used for the case of $T=6$ even when $T>6$. This corresponds to a non-tail normalization. That is, the transpose of $\mathbf{F}$ is written as:
$\mathbf{F}^{\prime}=\begin{bmatrix}f_{\gamma}&1&1&1&1&1&1&\cdots&1\\ f_{1}&f_{2}&f_{3}&f_{4}&0&1&f_{7}&\cdots&f_{T}\end{bmatrix}$
Our entire proof carries over without any modification. $\Box$
Given global identification of $\alpha$, it becomes clear that the rest of the parameters are almost surely globally identified as well. This can be formalized in the following Theorem.
Theorem 4.
Let $y_{it}$ be generated from (19). Then, under Assumptions 1’-6’, all the parameters in the model are almost surely globally identified.
Proof: Identical to the Proof of Theorem 2 in Section 4.4.
6 Concluding remarks
In this paper, we demonstrate that the parameters of a dynamic panel data model with interactive effects can be globally identified. We extend the seminal result of Anderson and Rubin (1956) to dynamic panels with interactive effects, a problem that was previously thought to be non-globally identifiable.
The significance of our result lies in the fact that, despite the high degree of nonlinearity in the problem, when the data-generating process corresponds to the specified model, there exists a one-to-one correspondence between the population moments and the model parameters.
Our proof shows that global identification is achievable when $T\geq 2(\bar{r}+1)$. An interesting open question is whether global identification can be attained under weaker conditions, such as when $T$ satisfies the Lebdermann bound plus one. The requirement for an additional period, beyond the Lebdermann bound, arises from the dynamic nature of the model, which includes an extra parameter compared to the static case.
Appendix A Explicit expression for $\mathbf{O}_{i,j}$ in Section 4 and $\bar{r}=1$
As $\mathbf{O}$ is symmetric, then $\mathbf{O}_{i,j}=\mathbf{O}_{j,i}$.The positions of $\mathbf{O}$ used in section 4.3 are:
$\mathbf{O}_{1,2}$ $=\left(d_{1}+\Psi\right)\,\left(\alpha-\widetilde{\alpha}\right)+f_{2}\Psi$
$\mathbf{O}_{1,3}$ $=f_{3}\,\Psi-\left(d_{1}+\Psi\right)\,\left(\alpha\,\widetilde{\alpha}-\alpha^{2}\right)+f_{2}\,\Psi\,\left(\alpha-\widetilde{\alpha}\right)$
$\mathbf{O}_{1,4}$ $=f_{4}\,\Psi-\left(\alpha^{2}\,\widetilde{\alpha}-\alpha^{3}\right)\,\left(d_{1}+\Psi\right)+f_{3}\,\Psi_{1}\,\left(\alpha-\widetilde{\alpha}\right)$
$\quad-f_{2}\,\Psi\,\left(\alpha\,\widetilde{\alpha}-\alpha^{2}\right)$
$\mathbf{O}_{2,3}$ $=\left(\alpha-\widetilde{\alpha}\right)\,\left(\Psi\,{f_{2}}^{2}+\Psi\,\left(\alpha-\widetilde{\alpha}\right)\,f_{2}+d_{2}\right)$
$\quad-\left(\alpha\,\widetilde{\alpha}-\alpha^{2}\right)\,\left(\left(d_{1}+\Psi\right)\,\left(\alpha-\widetilde{\alpha}\right)+f_{2}\,\Psi\right)$
$\quad+f_{2}\,f_{3}\,\Psi+f_{3}\,\Psi\,\left(\alpha-\widetilde{\alpha}\right)$
$\mathbf{O}_{2,4}$ $=\left(\alpha-\widetilde{\alpha}\right)\,\left(f_{2}\,f_{3}\,\Psi+f_{3}\,\Psi\,\left(\alpha-\widetilde{\alpha}\right)\right)$
$\quad-\left(\alpha\,\widetilde{\alpha}-\alpha^{2}\right)\,\left(\Psi\,{f_{2}}^{2}+\Psi\,\left(\alpha-\widetilde{\alpha}\right)\,f_{2}+d_{2}\right)$
$\quad-\left(\alpha^{2}\,\widetilde{\alpha}-\alpha^{3}\right)\,\left(f_{2}\,\Psi+\left(d_{1}+\Psi\right)\,\left(\alpha-\widetilde{\alpha}\right)\right)$
$\quad+f_{2}\,f_{4}\,\Psi+f_{4}\,\Psi\,\left(\alpha-\widetilde{\alpha}\right)$
$\mathbf{O}_{3,4}$ $=\left(\alpha-\widetilde{\alpha}\right)\,\left(d_{3}+{f_{3}}^{2}\,\Psi-f_{3}\,\Psi\,\left(\alpha\,\widetilde{\alpha}-\alpha^{2}\right)+f_{2}\,f_{3}\,\Psi\,\left(\alpha-\widetilde{\alpha}\right)\right)$
$\quad-\left(\alpha\,\widetilde{\alpha}-\alpha^{2}\right)\,\left(\left(\Psi\,{f_{2}}^{2}+d_{2}\right)\,\left(\alpha-\widetilde{\alpha}\right)+f_{2}\,f_{3}\,\Psi-f_{2}\,\Psi\,\left(\alpha\,\widetilde{\alpha}-\alpha^{2}\right)\right)$
$\quad-\left(\alpha^{2}\,\widetilde{\alpha}-\alpha^{3}\right)\,\left(f_{3}\,\Psi-\left(d_{1}+\Psi\right)\,\left(\alpha\,\widetilde{\alpha}-\alpha^{2}\right)+f_{2}\,\Psi\,\left(\alpha-\widetilde{\alpha}\right)\right)$
$\quad+f_{3}\,f_{4}\,\Psi-f_{4}\,\Psi\,\left(\alpha\,\widetilde{\alpha}-\alpha^{2}\right)+f_{2}\,f_{4}\,\Psi\,\left(\alpha-\widetilde{\alpha}\right)$
Appendix B Rotational indeterminacy in the presence of individual fixed-effects and $\bar{r}=2$
In general, factor models exhibit rotational indeterminacy, meaning that identification is possible up to a $\bar{r}\times\bar{r}$ rotation matrix $\mathbf{A}$. The presence of individual fixed effects imposes a restriction on the possible values of the factor matrix $\mathbf{F}$. We explain why, in the case of individual fixed effects and $\bar{r}=2$, only two restrictions are needed to eliminate rotational indeterminacy, as opposed to $\bar{r}^{2}=4$ restrictions in the unrestricted case. Let $\mathbf{A}$ be an invertible $2\times 2$ matrix, defined as:
$\mathbf{A}=\begin{bmatrix}a&b\\ c&d\\ \end{bmatrix}$
Next, recall the factor structure with individual fixed effects:
$\mathbf{F}=\begin{bmatrix}f_{\gamma}&f_{1}\\ 1&f_{2}\\ \vdots&\vdots\\ 1&f_{T}\\ \end{bmatrix},\quad\eta_{i}=\begin{bmatrix}\gamma_{i}\\ \lambda_{i}\end{bmatrix}$
Rotational indeterminacy implies that:
$\mathbf{F}\eta_{i}=\mathbf{F}\mathbf{A}\mathbf{A}^{-1}\eta_{i}$
However, for $\mathbf{F}\mathbf{A}$ to maintain the structure of $\mathbf{F}$ (which includes a column of nearly all elements equal to one), we immediately conclude that $a=1$ and $c=0$, leaving only two free parameters, $b$ and $d$. Hence, only two restrictions are required on $\mathbf{F}$ to resolve rotational indeterminacy.
Under the tail normalization discussed in the main text, we implicitly assume that the last $2\times 2$ block of $\mathbf{F}$ is an invertible matrix, which is equivalent to the condition $f_{T-1}\neq f_{T}$. The case where $f_{T-1}=f_{T}$ is considered a zero-probability event by assumption.