### v2

# Global identification of dynamic panel models with interactive effects

Abstract
We investigate the problem of global identification in dynamic panel models with interactive effects, under the large-$N$, fixed-$T$ setting. While local identification, typically established via the Jacobian matrix, is well understood, global identification has remained a more elusive and challenging issue. It is commonly believed to be unachievable in this context. However, we demonstrate that the model is, in fact, globally identified. Our analysis also covers models with additive fixed effects, including cases with unit roots, where previous studies have reported non-identification based on certain estimators. We show that, even in these settings, global identification holds.
Keywords: Dynamic panel models, global identification, factor models, panel data, interactive effects
1 Introduction
Consider the dynamic panel data model with interactive effects
$y_{it}=\alpha y_{it-1}+\delta_{t}+\lambda_{i}^{\prime}f_{t}+\varepsilon_{it}$ (1)
where $i=1,2,\dots,N$ and $t=1,2,\dots,T$. $y_{it}$ is the outcome variable, $\alpha$ is the autoregressive parameter, whereas $\delta_{t}$ captures time-specific fixed effects. $f_{t}$ is an $\bar{r}\times 1$ vector containing the value of the $\bar{r}$ latent factors at time $t$, and $\lambda_{i}$ is an $\bar{r}\times 1$ vector of factor loadings for individual $i$. $\varepsilon_{it}$ is the idiosyncratic error term, which we assume to be i.i.d. across $i$, heteroskedastic across $t$ but serially uncorrelated, i.e. $\mathbb{E}[\varepsilon_{it}^{2}]=\sigma_{t}^{2}\;\text{and}\;\mathbb{E}[\varepsilon_{it}\varepsilon_{is}]=0\;$ for $t\neq s$, and independent from the interactive effects. In this model, $y_{it}$ is the only observable variable.
This framework provides a flexible way to model unobservable common trends and heterogeneous responses. The model allows the common shocks (modeled by $f_{t}$) to have a heterogeneous effect in individuals (modeled by $\lambda_{i}$) on the outcome variable $y_{it}$. A notable special case is the standard additive fixed-effects model, where $y_{it}=\delta_{t}+\gamma_{i}+\varepsilon_{it}$, which corresponds to $\alpha=0$ and a single constant factor $f_{t}=1$. Thus, this model generalizes the additive framework by allowing for richer dynamics and heterogeneity in the factor structure. Parameter estimation in this model can be conducted via quasi maximum likelihood based methods as in Hayakawa et al., (2023) and Bai, (2024).
This paper studies the problem of global identification of the model parameters when the number of cross-sectional units $(N)$ is large, but the time dimension $(T)$ is fixed. This case is particularly relevant in applied economics because many panel datasets, such as those involving countries, firms, or households, span a large number of cross-sectional units but only cover a limited time horizon. For instance, macroeconomic panels often include many countries but only a few decades of annual data, making the short-$T$, large-NN framework empirically realistic.
Following Rothenberg, (1971), we say that a parameter point $\theta^{0}$ in the parametric space $\Theta$ is globally identifiable if there is no other $\widetilde{\theta}\in\Theta$ which is observationally equivalent. Establishing the identifiability of a model’s parameters is crucial, as it ensures that the estimation procedure yields meaningful and interpretable results.
Despite the apparent simplicity of the model, demonstrating parameter identifiability is a nontrivial task. The seminal work of Anderson and Rubin, (1956) established that in the absence of the autoregressive term ($\alpha=0$), the parameters of the model are globally identified up to an indeterminacy given by an $\bar{r}\times\bar{r}$ rotation matrix, provided that $T\geq 2\bar{r}+1$.
To the best of our knowledge, no existing study has established the global identifiability of the parameters in the model described by (1). Bai, (2024) demonstrates that the parameters of this model are locally identified when $\bar{r}$ does not exceed the modified Ledermann bound (Ledermann,, 1937). In contrast, Hayakawa et al., (2023) argues that global identification is unattainable in dynamic panel data models with interactive effects.
We demonstrate that under standard assumptions in the factor analysis literature, the parameters of the model are almost surely globally identifiable when $T\geq 2(\bar{r}+1)$. The global identification is achieved using the first two moments. Since the quasi-Gaussian likelihood depends only on the first two moments, this equivalence implies global identification for quasi-Gaussian maximum likelihood estimation. This result is of particular importance, as it establishes a one-to-one correspondence between the population moments and the model parameters, despite high degree of nonlinearity of the problem.
This paper contributes to the literature on global identification and panel data models. The question of global identification in parametric models has been the focus of extensive research (see, for example, Shapiro, (1985), Komunjer and Ng, (2011), Komunjer, (2012), and Kociecki and Kołasa, (2023)), while other authors have focused on trying to understand when and why identification fails (Forneron,, 2024). A comprehensive review of this literature is provided in Lewbel, (2019).
Panel data models have been a core focus in econometrics for many years, with extensive study and application over the past few decades. Notable monographs and textbooks on panel data include works by Arellano, (2003), Baltagi, (2008), Hsiao, (2022), and Wooldridge, (2010), among others. Moreover, the study of panel data models with interactive fixed effects [^1] Sometimes referred to in the literature as multiple time-varying individual effects (Ahn et al.,, 2013) or multifactor error structure (Pesaran,, 2006). has gained notable attention during the recent years. While there is an extensive literature in this area, some of the most noteworthy papers include Holtz-Eakin et al., (1988), Ahn et al., (2001), Pesaran, (2006), Bai, (2009), and Ahn et al., (2013). More recently, there has been significant research on dynamic panel models with interactive effects, including studies by Moon and Weidner, (2017), Hayakawa et al., (2023), and Bai, (2024), to name a few. However, these studies have not addressed the global identifiability of the model parameters for the short-$T$ case, a gap this paper aims to fill.
We also argue that the techniques developed in this paper can be used to prove identification in a broad class of dynamic models featuring time, individual and interactive effects. For example, we show that in a dynamic panel with individual effects (Arellano and Bond,, 1991), the autoregressive coefficient can be globally identified even when $\alpha=1$, a case previously thought to be non-identifiable (see Blundell and Bond, 1998; Sentana, 2024).
The remainder of the paper is organized as follows: Section 2 formalizes the identification problem in our context. Section 3 outlines the assumptions required for our proof and presents some useful mathematical results that will be instrumental in our analysis. Section 4 derives intermediate results and provides the main proof of global identification under the stated assumptions. Section 5 discusses identification in the presence of individual fixed effects. Finally, Section 6 provides some concluding remarks with a discussion of potential extensions of the paper.
2 Framework
2.1 The Model in Matrix Notation
Before formally defining global identification, it is useful to express the model in matrix notation. Suppose we observe $y_{i}=(y_{i1},y_{i2},\dots,y_{iT})^{\prime}$. We begin by projecting the initial observation $y_{i1}$ onto $1,\lambda_{i}]$, which gives
$y_{i1}=\delta_{1}^{*}+\lambda_{i}^{\prime}f_{1}^{*}+\varepsilon_{i1}^{*}$
where $\delta_{1}^{*}$ and $f_{1}^{*}$ are the projection coefficients, and $\varepsilon_{i1}^{*}$ is the projection residual. Since $\delta_{1}^{*}$ and $f_{1}^{*}$ are free parameters, and $\varepsilon_{it}$ in the model are not required to have identical distributions, we can simply rename $\delta_{1}^{*},f_{1}^{*}$, and $\varepsilon_{i1}^{*}$ as $\delta_{1}$, $f_{1}$, and $\varepsilon_{i1}$, respectively, without any loss of generality. So we can write $y_{i1}=\delta_{1}+\lambda_{i}^{\prime}f_{1}+\varepsilon_{i1}$.
Next, stacking observations over $t$, we obtain
$\mathbf{B}y_{i}=\delta+\mathbf{F}\lambda_{i}+\varepsilon_{i}$
where $y_{i}=[y_{i1},y_{i2},\dots,y_{iT}]^{\prime}$ is a $T\times 1$ vector of individual observations for $t=1,\dots,T$, $\delta=[\delta_{1},\dots,\delta_{T}]^{\prime}$ is a $T\times 1$ vector, and $\mathbf{F}=[f_{1},f_{2},\dots,f_{T}]^{\prime}$ is a $T\times\bar{r}$ matrix containing the $\bar{r}$ factors, $\lambda_{i}=[\lambda_{i1},\lambda_{i2},\dots,\lambda_{i\bar{r}}]^{\prime}$ is an $\bar{r}\times 1$ vector of factor loadings for individual $i$, and $\varepsilon_{i}=[\varepsilon_{i1},\varepsilon_{i2},\dots,\varepsilon_{iT}]^{\prime}$ is a $T\times 1$ vector of idiosyncratic errors. Furthermore,
$\mathbf{B}=\begin{bmatrix}1&0&0&\dots&0\\ -\alpha&1&0&\dots&0\\ 0&-\alpha&1&\dots&0\\ \vdots&\vdots&\ddots&\ddots&\vdots\\ 0&\dots&0&-\alpha&1\end{bmatrix}$
is a nonsingular lower-triangular $T\times T$ matrix, hence invertible. Let $\mathbf{\Gamma}=\mathbf{B}^{-1}.$ Pre-multiplying by $\mathbf{B}^{-1}$ allows us to rewrite the model as
$y_{i}=\mathbf{\Gamma}\delta+\mathbf{\Gamma}\mathbf{F}\lambda_{i}+\mathbf{\Gamma}\varepsilon_{i}$ (2)
where $\mathbf{\Gamma}$ is $T\times T$ and is given by
$\mathbf{\Gamma}=\mathbf{B}^{-1}=\begin{bmatrix}1&0&\dots&0\\ \alpha&1&\dots&0\\ \vdots&\vdots&\ddots&\vdots\\ \alpha^{T-1}&\dots&\alpha&1\end{bmatrix}$
In the original model, the errors had a diagonal covariance matrix given by $\mathbf{D}=\operatorname*{diag}(d_{1},d_{2},\dots,d_{T}):=\operatorname*{diag}(\sigma_{1}^{2},\dots,\sigma_{T}^{2})$ (for notational simplicity, we write $\sigma_{t}^{2}$ as $d_{t}$). After transformation, the idiosyncratic error $\mathbf{\Gamma}\varepsilon_{i}$ has covariance matrix $\mathbf{\Gamma}\mathbf{D}\mathbf{\Gamma}^{\prime}$, which is no longer diagonal. Let
$\mathbf{\Psi}_{N}=\frac{1}{N-1}\sum_{i=1}^{N}(\lambda_{i}-\bar{\lambda})(\lambda_{i}-\bar{\lambda})^{\prime}$
denote the sample covariance matrix of $\lambda_{i}$, where $\bar{\lambda}=\frac{1}{N}\sum_{i=1}^{N}\lambda_{i}$ and let $\mathbb{E}[\mathbf{\Psi}_{N}]=\mathbf{\Psi}$. Similarly, define
$\mathbf{S}_{N}=\frac{1}{N-1}\sum_{i=1}^{N}(y_{i}-\bar{y})(y_{i}-\bar{y})^{\prime}$
as the sample covariance matrix of $y_{i}$, where $\bar{y}=\frac{1}{N}\sum_{i=1}^{N}y_{i}$.
The expectation of the sample covariance matrix of $y_{i}$ is then given by
$\mathbb{E}[\mathbf{S}_{N}]=\mathbf{\Sigma}(\theta^{0})=\mathbf{\Gamma}(\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}+\mathbf{D})\mathbf{\Gamma}^{\prime}=\mathbf{\Gamma}\mathbf{\Omega}\mathbf{\Gamma}^{\prime}$ (3)
where $\theta^{0}=(\alpha,\mathbf{F},\mathbf{\Psi},\mathbf{D})$ is the true vector of structural parameters. The matrix $\mathbf{\Sigma}$ is a $T\times T$ symmetric positive definite (PD) matrix, $\mathbf{\Psi}$ is an $\bar{r}\times\bar{r}$ symmetric PD matrix, $\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}$ is a $T\times T$ symmetric positive semi-definite (PSD) matrix of rank $\bar{r}$, and $\mathbf{D}$ is a $T\times T$ diagonal PD matrix. $\mathbf{\Omega}=\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}+\mathbf{D}$ is a $T\times T$ symmetric PD matrix corresponding to the expected value of the sample covariance matrix of $y_{i}$ in a basic model without an autoregressive term $\alpha=0$ (Anderson and Rubin,, 1956).
2.2 Global identification of the model
This paper aims to establish that the parameters of the model are globally identifiable from the Gaussian quasi log-likelihood perspective, which is equivalent to demonstrating global identification from the first two moments. Specifically, this means that the expectation and variance of $y_{i}$ uniquely determine the structural parameters of the model, as there exists a one-to-one mapping between them.
To formalize this argument, define $\delta^{\dagger}=\mathbf{\Gamma}\delta$. The Gaussian quasi log-likelihood function is given by
$\frac{N}{2}\ln|\mathbf{\Sigma}(\theta)|-\frac{1}{2}\sum_{i=1}^{N}(y_{i}-\delta^{\dagger})^{\prime}\mathbf{\Sigma}(\theta)^{-1}(y_{i}-\delta^{\dagger})$
Concentrating out $\delta^{\dagger}$, which is estimated by $\bar{y}$, and dividing by $N$, we obtain [^2] In the MLE, the definition of $\mathbf{S}_{N}$ needs a slight modification, namely, 1/(N1)1/(N-1) is replaced by $1/N$.
$\frac{1}{N}\ell_{N}(\theta)=-\frac{1}{2}\ln|\mathbf{\Sigma}(\theta)|-\frac{1}{2}\mathrm{tr}[\mathbf{\Sigma}(\theta)^{-1}\mathbf{S}_{N}]$
For a fixed $T$ and as $N\to\infty$, by the law of large numbers, $\mathbf{S}_{N}\overset{p}{\to}\mathbf{\Sigma}(\theta^{0})$, we have
$\frac{2}{N}\ell_{N}(\theta)\overset{p}{\to}-\ln|\mathbf{\Sigma}(\theta)|-\mathrm{tr}\left[\mathbf{\Sigma}(\theta^{0})\mathbf{\Sigma}(\theta)^{-1}\right]$ (4)
Furthermore, it is known (Arnold,, 1981) that this function has a unique maximizer at $\mathbf{\Sigma}(\theta)=\mathbf{\Sigma}(\theta^{0})$. While this result ensures that $\theta^{0}$ is a global maximizer of the objective function, it does not immediately establish uniqueness—hence, global identification—because it does not preclude the possibility that other parameter values $\widetilde{\theta}\neq\theta^{0}$ satisfy $\mathbf{\Sigma}(\widetilde{\theta})=\mathbf{\Sigma}(\theta^{0})$. In the following sections, we investigate whether any alternative parameter set $\widetilde{\theta}\neq\theta^{0}$ satisfies $\mathbf{\Sigma}(\widetilde{\theta})=\mathbf{\Sigma}(\theta^{0})$. We demonstrate that, almost surely, no such alternative parametrization exists. Consequently, the quasi log-likelihood function is uniquely maximized at the true parameter value $\theta^{0}$, establishing global identification. This result, which has not been previously shown and was thought to be untrue (Hayakawa et al.,, 2023), constitutes a novel contribution.
To prove this claim, consider an alternative parameterization
  • $\mathbf{\widetilde{F}}=\left[\widetilde{f}_{1},\widetilde{f}_{2},\dots,\widetilde{f}_{T}\right]^{\prime}$;
  • $\mathbf{\widetilde{\Psi}}$ a symmetric PD $\bar{r}\times\bar{r}$ matrix;
  • $\mathbf{\widetilde{D}}=\operatorname*{diag}(\widetilde{d}_{1},\dots\widetilde{d}_{T})$ a diagonal PD $T\times T$ matrix;
  • $\widetilde{\mathbf{\Gamma}}$ the analogue of $\mathbf{\Gamma}$ under an alternative parameter $\widetilde{\alpha}$, given by
$\mathbf{\widetilde{\Gamma}}=\begin{bmatrix}1&0&\dots&0\\ \widetilde{\alpha}&1&\dots&0\\ \vdots&\vdots&\ddots&\vdots\\ \widetilde{\alpha}^{T-1}&\dots&\widetilde{\alpha}&1\end{bmatrix}$
By definition, $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}$ is a symmetric PSD $T\times T$ matrix of rank $\bar{r}$.
The model is globally identified at $\theta^{0}$ if whenever $\mathbf{\Sigma}(\widetilde{\theta})=\mathbf{\Sigma}(\theta^{0})$, it must hold that $\widetilde{\alpha}=\alpha$, $\mathbf{\widetilde{D}}=\mathbf{D}$, and $\mathbf{\widetilde{F}\widetilde{\Psi}\widetilde{F}}’=\mathbf{F\Psi F’}$. Formally:
$\mathbf{\Gamma}(\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}+\mathbf{D})\mathbf{\Gamma}^{\prime}=\mathbf{\widetilde{\Gamma}}(\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}+\mathbf{\widetilde{D}})\mathbf{\widetilde{\Gamma}}^{\prime}$ (5)
If and only if $\widetilde{\alpha}=\alpha$, $\mathbf{\widetilde{D}}=\mathbf{D}$, and $\mathbf{\widetilde{F}\widetilde{\Psi}\widetilde{F}}’=\mathbf{F\Psi F’}$ (or simply $\widetilde{\theta}=\theta^{0}$).
3 Assumptions and useful results
In order to proceed with our global identification proof, we need to make some assumptions. These assumptions are not restrictive and are rather standard in the factor analysis literature.
Assumption A1.
The number of distinct factors is known and equal to $\bar{r}$.
Assumption A2.
$\mathbf{\Psi}$ is an unrestricted symmetric and PD matrix, and $\mathbf{F}=(\mathbf{I}_{\bar{r}},\mathbf{F}_{2}^{\prime})^{\prime}$.
This assumption comes from the fact that factor models (and thus interactive effects) are identifiable up to a $\bar{r}\times\bar{r}$ matrix rotation, so $\bar{r}^{2}$ restrictions must be imposed for identification. A discussion of why this is the case can be found in Section 5 of Anderson and Rubin, (1956). We work with this restriction for convenience, but any other normalization that imposes $\bar{r}^{2}$ restriction would also work.
Assumption A3.
$T\geq 2(\bar{r}+1)$.
Assumption A4.
$\varepsilon_{i}$ are i.i.d. over $i$. $\mathbb{E}[\varepsilon_{it}]=0$, $\mathbb{V}(\varepsilon_{it})=d_{t}>0$, and $\varepsilon_{it}$ are independent over $t$. $\varepsilon_{it}$ is independent from the factor loadings $\lambda_{i}$ and the factors $\mathbf{F}$.
Assumption A5.
$\lambda_{i}$ are either random or fixed constants with $\mathbf{\Psi}_{N}\to\mathbf{\Psi}$ as $N\to\infty$.
Assumption A6.
$\alpha,d_{1},d_{2},\dots,d_{T},\mathbf{\Psi}$ are fixed population parameters.
Assumption A7.
Let $\mathbf{F}_{2}=\begin{bmatrix}f_{\bar{r}+1,1}&f_{\bar{r}+1,2}&\dots&f_{\bar{r}+1,\bar{r}}\\ \vdots&\vdots&\vdots&\vdots\\ f_{T,1}&f_{T,2}&\dots&f_{T,\bar{r}}\end{bmatrix}$ be the random elements of $\mathbf{F}$. Then, $\mathbf{F}_{2}$ is a continuous random matrix with support in $\Pi\in\mathbb{R}^{(T-\bar{r})\times\bar{r}}$.
Additionally, throughout the paper, we will reference several established results from linear algebra and probability theory. To ensure they are easily accessible to the reader, we present them as lemmas in this section.
Lemma 1.
Let $\mathbf{A}$ be a $T\times T$ matrix of rank $\bar{r}<T$. Then, the determinants of all the $(\bar{r}+1)\times(\bar{r}+1)$ minors of $\mathbf{A}$ are equal to 0.
Lemma 2.
Let $\mathbf{X}$ be a continuous random vector with $\mathrm{supp}(\mathbf{X})\subset\mathbb{R}^{q}$. Furthermore, let $\Upsilon$ be a subset of $\mathrm{supp}(\mathbf{X})$ with dimension lower than that of $\mathrm{supp}(\mathbf{X})$. Then
$P(\mathbf{X}\in\Upsilon)=0$
Lemma 3.
(Laplace Expansion) Let $\mathbf{A}$ be a $k\times k$ matrix. Then, we can write the determinant of $\mathbf{A}$ as a weighted sum of the determinants of some $(k-1)\times(k-1)$ minors of $\mathbf{A}$. In particular, for each row $i$, we can write
$\det\left(\mathbf{A}\right)=\sum_{j=1}^{k}(-1)^{i+j}A_{i,j}\det\left(\mathbf{A}_{-i,-j}\right)$
where $\mathbf{A}_{-i,-j}$ is the minor obtained by removing row $i$ and column $j$, and $A_{i,j}$ represents the $(i,j)$-th element of the matrix $\mathbf{A}$.
In this paper, we will apply Lemma 3 for cases in which $\mathbf{A}$ itself is a minor of a larger matrix. Let $\mathbf{H}$ be a $T\times T$ matrix, and let $\mathbf{M}^{\mathbf{H}}_{R,C}$ be a minor of $\mathbf{H}$ of dimension $k$ including rows $R=\{r_{1},\dots,r_{k}\}$ and columns $C=\{c_{1},\dots,c_{k}\}$ of the original matrix. Therefore, for any $r\in R$ we can denote $i\in\{1,\dots,k\}$ as the position of row $r$ in the minor $\mathbf{M}^{\mathbf{H}}_{R,C}$, and the Laplace decomposition of this minor is given by:
$\det(\mathbf{M}^{\mathbf{H}}_{R,C})=\sum_{j=1}^{k}(-1)^{i+j}H_{r,c_{j}}\det\left(\mathbf{M}^{\mathbf{H}}_{R-r,C-c_{j}}\right)$
where $H_{r,c_{j}}$ is the $(r,c_{j})$-th element of $\mathbf{H}$, and $\mathbf{M}^{\mathbf{H}}_{R-r,C-c_{j}}$ is the minor [^3] Where RrR-r means removal $r$ from the set $R$, and $C-c_{j}$ means removal of $c_{j}$ from the set $C$. They correspond to the usual notation $R\setminus\{r\}$ and $C\setminus\{c_{j}\}$. of $\mathbf{H}$ that includes rows RrR-r, and columns $C-c_{j}$.
In our induction analysis to be used later, we will further need to indicate the dimension of the minor. The minor on the left hand side will be written as $\mathbf{M}^{\mathbf{H},k}_{R,C}$, and the minor on the right hand side will be written as $\mathbf{M}^{\mathbf{H},k-1}_{R-r,C-c_{j}}$
4 Global identification
In this section, we demonstrate that, under Assumptions A1-A7, the parameters of the dynamic panel data model with interactive effects described in (1) are globally identified.
4.1 Proof insight
The proof we propose draws inspiration from Theorem 5.1 of Anderson and Rubin, (1956). In that theorem, the authors establish a sufficient condition for global identification of a factor model that did not include an autoregressive component. In their simpler model, global identification requires
$\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}+\mathbf{D}=\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}+\mathbf{\widetilde{D}}$ (6)
if and only if $\mathbf{\widetilde{D}}=\mathbf{D}$ (a $T\times T$ diagonal PD matrix), and $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}=\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}$ (a $T\times T$ symmetric PSD matrix of rank $\bar{r}$).
Their proof relies on the fact that the off-diagonal elements of $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}$ must match those of $\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}$ for (6) to hold, because both $\mathbf{D}\ \text{and}\ \mathbf{\widetilde{D}}$ are diagonal matrices. Thus, the problem reduces to showing that the diagonal elements of these matrices are also equal. To establish this, they rewrite (6) as
$\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}=\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}+\mathbf{\widetilde{D}}-\mathbf{D}$
Then, they leverage the fact that all determinants of the $(\bar{r}+1)\times(\bar{r}+1)$ minors of $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}+\mathbf{\widetilde{D}}-\mathbf{D}$ must be zero for (6) to hold (Lemma 1). This follows as a necessary condition because, by assumption, $\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}$ is of rank $\bar{r}$. This procedure enables them to pin down the diagonal terms of $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}$, thereby establishing that $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}=\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}$. Once this is established, it follows immediately that $\mathbf{\widetilde{D}}=\mathbf{D}$, completing the proof.
In their approach, the authors focus exclusively on the minors involving diagonal elements of the matrix and do not consider what we will refer to as diagonal exclusion minors (minors that do not include diagonal terms of the original matrix) as these do not contain relevant information for their analysis. In contrast, our proof takes a different approach: we rearrange the terms of our original problem—Equation (5)—in a way that allows us to exploit the information contained in the diagonal exclusion minors of this transformed problem. This novel approach enables us to demonstrate that (5) holds if and only if $\widetilde{\alpha}=\alpha$, thereby establishing the global identification of $\alpha$.
Once we establish that $\alpha$ is globally identified, it follows that our identification condition can be rewritten in the form proposed by Anderson and Rubin, (1956), ensuring that all other parameters can also be globally identified.
Thus, the primary focus of our analysis is to show that $\alpha$ is globally identified (or equivalently, that it must be the case that $\widetilde{\alpha}=\alpha$ in order for (5) to hold). This task requires some intermediate results, which we will articulate and prove throughout the paper. The main contribution of this work is to formally establish the following Theorem.
Theorem 1.
Let $y_{it}$ be generated from (1). Then, under Assumptions A1-A7, $\alpha$ is almost surely globally identified.
Once this is shown, it is immediate to conclude that all the parameters of the model are almost surely globally identified as well, which is formally stated in the following Theorem
Theorem 2.
Let $y_{it}$ be generated from (1). Then, under Assumptions A1-A7, all the parameters in the model are almost surely globally identified.
Remark: The qualifier ‘almost surely’ is with respect to the distribution of $\mathbf{F}$. For any fixed $\alpha$, identification holds for almost all realizations of $\mathbf{F}$. The practical implication is that as long as there is sufficient variation in $f_{t}$, both $\alpha$ and other parameters are globally identifiable. This is particularly true when $T$ is sufficiently large, ensuring a sufficient number of distinct realizations of $f_{t}$.
4.2 Intermediate results
In order to prove Theorem 1, we will rearrange the problem in a way that resembles the proof of Anderson and Rubin, (1956) and then establish some properties of this modified problem. Start from (5). As $\mathbf{\widetilde{\Gamma}}$ is a unit-diagonal lower triangular matrix, it is invertible so that we can pre-multiply and post-multiply by $\mathbf{\widetilde{\Gamma}}^{-1}$ and $\mathbf{\widetilde{\Gamma}}^{-1^{\prime}}$ and re-write
$\mathbf{\widetilde{\Gamma}}^{-1}\mathbf{\Gamma}(\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}+\mathbf{D})\mathbf{\Gamma}^{\prime}\mathbf{\widetilde{\Gamma}}^{-1^{\prime}}=\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}+\mathbf{\widetilde{D}}$
Now, define
$\mathbf{Q}=\mathbf{\widetilde{\Gamma}^{-1}}\mathbf{\Gamma}=\begin{bmatrix}1&0&0&\dots&0\\ \alpha-\widetilde{\alpha}&1&0&\dots&0\\ \alpha(\alpha-\widetilde{\alpha})&\alpha-\widetilde{\alpha}&1&\dots&0\\ \vdots&\vdots&\ddots&\ddots&\vdots\\ \alpha^{T-2}(\alpha-\widetilde{\alpha})&\alpha^{T-3}(\alpha-\widetilde{\alpha})&\dots&\alpha-\widetilde{\alpha}&1\end{bmatrix}$
Note that $\mathbf{Q}$ can be decomposed into
$\mathbf{Q}=\mathbf{I}_{T}+(\alpha-\widetilde{\alpha})\mathbf{L}$
where $\mathbf{L}$ is a strictly lower-triangular Toeplitz matrix given by
$\mathbf{L}=\begin{bmatrix}0&0&0&\dots&0\\ 1&0&0&\dots&0\\ \alpha&1&0&\dots&0\\ \vdots&\vdots&\ddots&\ddots&\vdots\\ \alpha^{T-2}&\alpha^{T-3}&\dots&1&0\end{bmatrix}$
Then, subtract $\mathbf{\widetilde{D}}$ on both sides of the equation
$\mathbf{Q}(\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}+\mathbf{D})\mathbf{Q}^{\prime}-\mathbf{\widetilde{D}}=\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}$ (7)
To simplify notation, denote the left-hand side of (7) by
$\mathbf{O}\equiv\mathbf{Q}(\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}+\mathbf{D})\mathbf{Q}^{\prime}-\mathbf{\widetilde{D}}$ (8)
We next establish some preliminary results.
Claim 1.
Let $\mathbf{M}$ be any minor of $\mathbf{O}$ of dimension $k>\bar{r}$. Then $\det(\mathbf{M})=0$ is a necessary condition for (5) to hold.
Proof.
We have shown that (5) is equivalent to $\mathbf{O}=\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}$. Furthermore, $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}$ is a symmetric PSD matrix of rank $\bar{r}$ by assumption. Then, it must be the case that $\mathbf{O}$ is also a symmetric PSD matrix of rank $\bar{r}$ in order for (5) to hold. By Lemma 1 we know that if $\mathbf{O}$ is of rank $\bar{r}$ then the determinant of all the minors of $\mathbf{O}$ of dimension $k>\bar{r}$ must be equal to 0. $\square$
Furthermore, we can show that $\mathbf{O}$ can be decomposed in the following convenient way
Claim 2.
$\mathbf{O}$ can be decomposed into $\mathbf{O}=\mathbf{\Omega}-\mathbf{\widetilde{D}}+(\alpha-\widetilde{\alpha})\mathbf{J}(\widetilde{\alpha},\theta^{0}),$
where $\mathbf{\Omega}$ is defined as in (3) and $\mathbf{J}(\widetilde{\alpha},\theta^{0})$ is given by:
$\mathbf{J}=\mathbf{L}\mathbf{\Omega}+\mathbf{\Omega}\mathbf{L}^{\prime}+(\alpha-\widetilde{\alpha})\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}.$ (9)
Proof.
By definition $\mathbf{\Omega}=\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}+\mathbf{D}$, so we can write $\mathbf{O}$ as
$\mathbf{O}=\mathbf{Q}\mathbf{\Omega}\mathbf{Q}^{\prime}-\mathbf{\widetilde{D}}$.
Recall that $\mathbf{Q}=\mathbf{I}_{T}+(\alpha-\widetilde{\alpha})\mathbf{L}$. Then
$\mathbf{O}=\left(\mathbf{I}_{T}+(\alpha-\widetilde{\alpha})\mathbf{L}\right)\mathbf{\Omega}\left(\mathbf{I}_{T}+(\alpha-\widetilde{\alpha})\mathbf{L}\right)^{\prime}-\mathbf{\widetilde{D}}$.
Distributing terms, we see that
$\mathbf{O}$ $=\mathbf{\Omega}-\mathbf{\widetilde{D}}+(\alpha-\widetilde{\alpha})\mathbf{L}\mathbf{\Omega}+\mathbf{\Omega}(\alpha-\widetilde{\alpha})\mathbf{L}^{\prime}+(\alpha-\widetilde{\alpha})^{2}\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}$
$=\mathbf{\Omega}-\mathbf{\widetilde{D}}+(\alpha-\widetilde{\alpha})\left[\mathbf{L}\mathbf{\Omega}+\mathbf{\Omega}\mathbf{L}^{\prime}+(\alpha-\widetilde{\alpha})\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}\right]$
$=\mathbf{\Omega}-\mathbf{\widetilde{D}}+(\alpha-\widetilde{\alpha})\mathbf{J}(\widetilde{\alpha},\theta^{0})$
where $\mathbf{J}(\widetilde{\alpha},\theta^{0})=\mathbf{L}\mathbf{\Omega}+\mathbf{\Omega}\mathbf{L}^{\prime}+(\alpha-\widetilde{\alpha})\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}.$ This proves the claim. $\square$
With this result in hand, we can proceed to derive the following corollaries.
Corollary 1.
We can write every non-diagonal element of $\mathbf{O}$ as $O_{r,c}=\Omega_{r,c}+(\alpha-\widetilde{\alpha})J_{r,c},$ where $J_{r,c}$ is the $(r,c)$-th element of $\mathbf{J}$.
Proof.
As $\mathbf{\widetilde{D}}$ is a diagonal matrix then $\widetilde{D}_{r,c}=0$ for all $r\neq c$. Using the expression for $\mathbf{O}$ derived in Claim 2:
$O_{r,c}=\Omega_{r,c}-\widetilde{D}_{r,c}+(\alpha-\widetilde{\alpha})J_{r,c}=\Omega_{r,c}+(\alpha-\widetilde{\alpha})J_{r,c}\quad r\neq c$
$\square$
Corollary 2.
For every row index $r\in\{\bar{r}+1,\dots T\}$ we have that $J_{r,c}\neq 0$ with probability 1 if $c\geq 2$. Similarly for every column index $c\in\{\bar{r}+1,\dots,T\}$ we have that $J_{r,c}\neq 0$ with probability 1 if $r\geq 2$.
Proof.
We only provide the proof for the first statement, as the second one is completely analogous given the symmetry of the problem.
To show this result, we use the fact that most of the positions of $\mathbf{\Omega}$ are non-zero with probability 1. First, note that by Assumption A2 we know that the first $\bar{r}\times\bar{r}$ block of $\mathbf{\Omega}$ is given by:
$\mathbf{\Omega}_{1:\bar{r},1:\bar{r}}=\mathbf{\Psi}+\mathbf{D}_{1:\bar{r},1:\bar{r}}$
where $\mathbf{\Omega}_{1:\bar{r},1:\bar{r}}$ represents the first $\bar{r}\times\bar{r}$ block of $\mathbf{\Omega}$ (same for $\mathbf{D}_{1:\bar{r},1:\bar{r}}$). Therefore, we cannot guarantee that the positions of $\mathbf{\Omega}_{1:\bar{r},1:\bar{r}}$ are non-zero, as the covariance between the factor loadings (the off-diagonal elements of $\mathbf{\Psi}$) could be equal to zero. Then, we use information contained in the remaining elements of $\mathbf{\Omega}$ to show this result.
From the definition of $\mathbf{\Omega}$ under Assumption A2 and Assumption A7 we know that $\Omega_{r,c}\neq 0$ with probability 1 if $r>\bar{r}$, as every position depends on at least one factor realization, so the case in which the position is exactly equal to zero is an event of probability zero (Lemma 2). Now, recall that $\mathbf{J}=\mathbf{L}\mathbf{\Omega}+\mathbf{\Omega}\mathbf{L}^{\prime}+(\alpha-\widetilde{\alpha})\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}$ is a $T\times T$ matrix. By Assumptions A6 and A7 we know that the elements of $\mathbf{\Omega}_{(\bar{r}+1):T,1:T}$ follow a continuous distribution dependent on $\mathbf{F}$. We establish the corollary in three steps.
First, we show that $\mathbf{\Omega}\mathbf{L}^{\prime}$ is a random matrix with non-zero elements for rows $r>\bar{r}$ and $c\neq 1$. From the definition of $\mathbf{L}$ in 2 we know that:
$L_{r,c}=\begin{cases}0,\quad r\leq c\\ 1,\quad r=c+1\\ \alpha^{r-c-1},\quad r>c+1\end{cases}$
Then, observe that
$(\mathbf{\Omega}\mathbf{L}^{\prime})_{r,c}=\sum_{i=1}^{T}\Omega_{r,i}L^{\prime}_{i,c}=\sum_{i=1}^{T}\Omega_{r,i}L_{c,i}=\sum_{i=1}^{c-1}\Omega_{r,i}\alpha^{c-i-1}$
where from the definition of $\mathbf{L}$ we know that all indexes $i\geq c$ are such that $L_{c,i}=0$. For $c\geq 2$, the elements $L_{c,c-1},L_{c,c-2},\dots L_{c,1}$ are all nonzero if $\alpha\neq 0$. If $\alpha=0$ we still have that $L_{c,c-1}=1$, which suffices for our purpose (the element $\Omega_{r,c-1}$ is non-zero with probability 1 when $r>\bar{r}$). Then, by Lemma 2 and the definition of $\mathbf{\Omega}$ for row indices $r>\bar{r}$, it follows that $\mathbf{\Omega}\mathbf{L}^{\prime}_{r,c}\neq 0$ with probability 1 if $c\geq 2$.
By the same reasoning, the matrix $\mathbf{L}\mathbf{\Omega}$ has zeros in the first row and nonzero entries $\mathbf{L}\mathbf{\Omega}_{r,c}$ for every column $c>\bar{r}$ with probability 1 if $r\geq 2$.
Third, we show that $\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}$ has non-zero elements for rows $r>\bar{r}+1$ and $c\geq 2$ with probability 1. Using matrix associativity,
$\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}_{r,c}=\sum_{i=1}^{T}L_{r,i}(\Omega L^{\prime})_{i,c}$
Since $L_{1,i}=0$ for all $i$ and $(\Omega L^{\prime})_{i,1}=0$ for all $i$, it follows that $\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}$ has zeros in the first row and first column. Moreover, for row indices $r>\bar{r}$ we know that the elements $L_{r,r-1},L_{r,r-2},\dots L_{r,1}$ are nonzero if $\alpha\neq 0$. Then, by Lemma 2 and the definition of $\mathbf{\Omega}$ for row indices $r>\bar{r}$, it follows that $\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}_{r,c}\neq 0$ with probability 1 if $\alpha\neq 0$. If $\alpha=0$ then we require $r>\bar{r}+1$ as $L_{r,r-2},\dots L_{r,1}$ are all equal to zero, so only $L_{r,r-1}=1$, but if $r=\bar{r}+1$ then $(\Omega L^{\prime})_{r-1,c}=(\Omega L^{\prime})_{\bar{r},c}$ may be equal to zero (the results derived before for $\mathbf{\Omega}\mathbf{L}^{\prime}$ hold under $r>\bar{r}$). Then, in the most general case, if $r>\bar{r}+1$ and $c\geq 2$ then $\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}_{r,c}\neq 0$ with probability 1.
Recall that $J_{r,c}=(\mathbf{\Omega}\mathbf{L^{\prime}})_{r,c}+(\mathbf{L}\mathbf{\Omega})_{r,c}+(\alpha-\widetilde{\alpha})(\mathbf{L}\mathbf{\Omega}\mathbf{L^{\prime}})_{r,c}$. We have shown that the first element is non-zero with probability 1 if $r>\bar{r}$, the second element is non-zero with probability 1 if $c>\bar{r}$ and $(\mathbf{L}\mathbf{\Omega}\mathbf{L^{\prime}})_{r,c}$ is non-zero with probability 1 if $r>\bar{r}+1$. The event in which the summands cancel out can only occur in a set of factor realizations that occur with probability zero (Lemma 2 and Assumption A7). This completes the proof. $\square$
Definition 1.
We refer to diagonal exclusion minors as minors that do not include any diagonal terms of the original matrix. Formally, let $\mathbf{A}$ be a $T\times T$ matrix, and define $\mathbf{M}^{\mathbf{A},k}_{R,C}$ as a diagonal exclusion minor of $\mathbf{A}$ of dimension $k$, consisting of rows $R=\{r_{1},r_{2},\dots,r_{k}\}$ and columns $C=\{c_{1},c_{2},\dots,c_{k}\}$.
By definition, $R$ and $C$ must satisfy
  1. 1.
    $|R|=|C|=k\leq\frac{T}{2}$.
  2. 2.
    $R\cap C=\emptyset$.
The minors in the definition are squared matrices, thus $|R|=|C|=k$. They do not contain diagonal elements, thus $R\cap C=\emptyset$. It follows that $|R|=|C|=k\leq\frac{T}{2}$.
Now, we are ready to state and show the following Propositions.
Proposition 1.
The determinant of every diagonal exclusion minor of dimension $k$ of $\mathbf{O}$ can be written as the sum of the determinant of the corresponding diagonal exclusion minor for $\mathbf{\Omega}$ and a term expressed as $(\alpha-\widetilde{\alpha})\widetilde{J}\left(\widetilde{\alpha},\theta^{0}\right)$
In mathematical terms, what we show is that given any diagonal exclusion minor of $\mathbf{O}$ of dimension $k$, say $\mathbf{M}^{\mathbf{O},k}_{R,C}$, we can write its determinant as the sum of the determinant of the diagonal exclusion minor of $\mathbf{\Omega}$ associated with the same rows and columns, say $\mathbf{M}^{\mathbf{\Omega},k}_{R,C}$, and a term that can be written as $(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$
$\det\left(\mathbf{M}^{\mathbf{O},k}_{R,C}\right)=\det\left(\mathbf{M}^{\mathbf{\Omega},k}_{R,C}\right)+(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ (10)
where $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ is a scalar that is a function of the true structural parameters $\theta^{0}$ and $\widetilde{\alpha}$. Note that the value of $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ will depend on which columns and rows were taken from the original matrix so in general, $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)\neq\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta^{0}\right)$ for any other valid set of rows and columns $R^{\prime}$ and $C^{\prime}$. To keep notation as simple as possible, we sometimes use $\widetilde{J}_{R,C}$ instead of $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ to represent the same object.
Proof.
We can prove Proposition 1 by induction.
1. Base case $k=1$: Without loss of generality $R=r$ and $C=c$. Then, a diagonal exclusion minor of $\mathbf{O}$ of dimension 1, say $\mathbf{M}^{\mathbf{O},1}_{r,c}$, is simply the term $O_{r,c}$ (for $r\neq c$) and its determinant is trivially given by
$\det\left(\mathbf{M}^{\mathbf{O},1}_{r,c}\right)=O_{r,c}$
From Corollary 1, we know that
$O_{r,c}=\Omega_{r,c}+(\alpha-\widetilde{\alpha})J_{r,c}\quad r\neq c$
Since $\Omega_{r,c}$ is a scalar we have that $\Omega_{r,c}=\det\left(\mathbf{M}^{\mathbf{\Omega},1}_{r,c}\right)$. We can rewrite the preceding equation as:
$\det\left(\mathbf{M}^{\mathbf{O},1}_{r,c}\right)=\det\left(\mathbf{M}^{\mathbf{\Omega},1}_{r,c}\right)+(\alpha-\widetilde{\alpha})\widetilde{J}_{r,c}$ (11)
where $\widetilde{J}_{r,c}=J_{r,c}$ when $k=1$. Therefore, the base step in the induction holds.
2. Inductive hypothesis: Assume our proposition holds for diagonal exclusion minors of dimension $(k-1)\times(k-1)$. This implies that for any set of valid $R$ and $C$ we have that
$\det\left(\mathbf{M}^{\mathbf{O},k-1}_{R,C}\right)=\det\left(\mathbf{M}^{\mathbf{\Omega},k-1}_{R,C}\right)+(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ (12)
3. Inductive step for $k$: We prove the desired result using both Corollary 1, and invoking the Laplace expansion for determinants (Lemma 3).
Let $\mathbf{M}^{\mathbf{O},k}_{R,C}$ be a diagonal exclusion minor of $\mathbf{O}$ of dimension $k$ with rows $R=\{r_{1},r_{2},...,r_{k}\}$ and columns $C=\{c_{1},c_{2},...,c_{k}\}$ satisfying the usual conditions. Before proceeding, we need to establish a correspondence between the indices of rows $R$ and columns $C$ in the original matrices (either $\mathbf{O}$ or $\mathbf{\Omega}$) and their new indices (i.e. $\{1,2,...,k\}$) in the $k\times k$ minors $\mathbf{M}^{\mathbf{O},k}_{R,C}$ and $\mathbf{M}^{\mathbf{\Omega},k}_{R,C}$. Without loss of generality, let $(i,j)$ denote the position of the $i$-th row in $R$ and the $j$-th column in $C$ within the minors $\mathbf{M}^{\mathbf{O},k}_{R,C}$ and $\mathbf{M}^{\mathbf{\Omega},k}_{R,C}$. This establishes that the $(i,j)$-th element of the minors $\mathbf{M}^{\mathbf{O},k}_{R,C}$ and $\mathbf{M}^{\mathbf{\Omega},k}_{R,C}$ are respectively given by $O_{r_{i},c_{j}}$ and $\Omega_{r_{i},c_{j}}$.
With this established, we can use Lemma 3 (fixing the row $r_{k}\in R$ in the original matrix) to write the determinant of $\mathbf{M}^{\mathbf{O},k}_{R,C}$ as
$\det\left(\mathbf{M}^{\mathbf{O},k}_{R,C}\right)=\sum_{j=1}^{k}(-1)^{k+j}O_{r_{k},c_{j}}\det\left(\mathbf{M}^{\mathbf{O},k-1}_{R-r_{k},C-c_{j}}\right)$ (13)
where $k$ and $j$ represent the indices in minor $\mathbf{M}^{\mathbf{O},k}_{R,C}$ of row $r_{k}$ and columns $c_{j}$ of the matrix $\mathbf{O}$. The scalar $O_{r_{k},c_{j}}$ is the $(k,j)$-th element of the minor $\mathbf{M}^{\mathbf{O},k}_{R,C}$; the matrix $\mathbf{M}^{\mathbf{O},k-1}_{R-r_{k},C-c_{j}}$ is the minor of matrix $\mathbf{O}$ with dimension k1k-1 that includes rows $R-r_{k}$ and columns $C-c_{j}$.
Because $r_{k}\neq c_{j}$ for all $j$ (recall $R\cap C=\emptyset$), we can use Corollary 1 to write
$O_{r_{k},c_{j}}=\Omega_{r_{k},c_{j}}+(\alpha-\widetilde{\alpha})J_{r_{k},c_{j}}$
Furthermore, by the inductive hypothesis
$\det\left(\mathbf{M}^{\mathbf{O},k-1}_{R-r_{k},C-c_{j}}\right)=\det\left(\mathbf{M}^{\mathbf{\Omega},k-1}_{R-r_{k},C-c_{j}}\right)+(\alpha-\widetilde{\alpha})\widetilde{J}_{R-r_{k},C-c_{j}}$
Plugging these results into (13) we have that
$\det\left(\mathbf{M}^{\mathbf{O},k}_{R,C}\right)=\sum_{j=1}^{k}(-1)^{k+j}\left(\Omega_{r_{k},c_{j}}+(\alpha-\widetilde{\alpha})J_{r_{k},c_{j}}\right)\left[\det\left(\mathbf{M}^{\mathbf{\Omega},k-1}_{R-r_{k},C-c_{j}}\right)+(\alpha-\widetilde{\alpha})\widetilde{J}_{R-r_{k},C-c_{j}}\right]$
Distributing terms we obtain
$\begin{split}&\det\left(\mathbf{M}^{\mathbf{O},k}_{R,C}\right)=\sum_{j=1}^{k}(-1)^{k+j}\Omega_{r_{k},c_{j}}\det\left(\mathbf{M}^{\mathbf{\Omega},k-1}_{R-r_{k},C-c_{j}}\right)+\\ &(\alpha-\widetilde{\alpha})\sum_{j=1}^{k}(-1)^{k+j}\left[J_{r_{k},c_{j}}\det\left(\mathbf{M}^{\mathbf{\Omega},k-1}_{R-r_{k},C-c_{j}}\right)+\Omega_{r_{k},c_{j}}\widetilde{J}_{R-r_{k},C-c_{j}}+(\alpha-\widetilde{\alpha})J_{r_{k},c_{j}}\widetilde{J}_{R-r_{k},C-c_{j}}\right]\end{split}$
Then, using Lemma 3 we can rewrite the first term of the expression above as
$\sum_{j=1}^{k}(-1)^{k+j}\Omega_{r_{k},c_{j}}\det\left(\mathbf{M}^{\mathbf{\Omega},k-1}_{R-r_{k},C-c_{j}}\right)=\det\left(\mathbf{M}^{\mathbf{\Omega},k}_{R,C}\right).$
Next, define $\widetilde{J}_{R,C}(\widetilde{\alpha},\theta^{0})$ as
$\widetilde{J}_{R,C}(\widetilde{\alpha},\theta^{0})=\sum_{j=1}^{k}(-1)^{k+j}\left[J_{r_{k},c_{j}}\det\left(\mathbf{M}^{\mathbf{\Omega},k-1}_{R-r_{k},C-c_{j}}\right)+\Omega_{r_{k},c_{j}}\widetilde{J}_{R-r_{k},C-c_{j}}+(\alpha-\widetilde{\alpha})J_{r_{k},c_{j}}\widetilde{J}_{R-r_{k},C-c_{j}}\right]$ (14)
This permits us to conclude that
$\det\left(\mathbf{M}^{\mathbf{O},k}_{R,C}\right)=\det\left(\mathbf{M}^{\mathbf{\Omega},k}_{R,C}\right)+(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}(\widetilde{\alpha},\theta^{0})$
$\square$
Proposition 2.
Consider any diagonal exclusion minor of $\mathbf{O}$ of dimension $k$. Then, $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ is a random polynomial in $\widetilde{\alpha}$ of degree almost surely in between $\max\{0,k-\bar{r}\}$ and 2k12k-1.
Proof.
First, note from (14) that $\widetilde{J}_{R,C}(\widetilde{\alpha},\theta^{0})$ is recursively defined, so once we know $\widetilde{J}_{R,C}(\widetilde{\alpha},\theta^{0})$ for $|R|=|C|=1$ we can compute $\widetilde{J}_{R,C}(\widetilde{\alpha},\theta^{0})$ for any arbitrary set of rows and columns.
When $k=1$, we have shown that $\widetilde{J}_{R,C}=J_{r,c}$. Now, recall the definition of $\mathbf{J}$ in Claim 2. $\mathbf{\Omega}$ and $\mathbf{L}$ do not depend on $\widetilde{\alpha}$, so $J_{r,c}$ is a polynomial of degree 1 in $\widetilde{\alpha}$ if $r>\bar{r}+1$ and $c\neq 1$ or $c>\bar{r}+1$ and $r\neq 1$, and it is a polynomial of degree 0 in $\widetilde{\alpha}$ if $r$ or $c$ are equal to 1 (see the discussion about $\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}$ in Corollary 2). Intermediate cases will depend on the values of the other parameters, but it will be a polynomial of either degree 0 or 1, depending on whether $\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}_{r,c}$ is equal or different from zero. Given the recursive structure of $\widetilde{J}_{R,C}(\widetilde{\alpha},\theta^{0})$ we can conclude that for any value of $k$, $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ is a polynomial in $\widetilde{\alpha}$.
Note that the coefficients of this polynomial depend on the true data-generating process parameters $\theta^{0}$, and that by Assumptions A6 and A7, $\mathbf{F}$ is drawn from a continuous probability distribution. Then, we can conclude that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ is a random polynomial in $\widetilde{\alpha}$ and that the randomness in the coefficients is given by the factor realizations $\mathbf{F}$.
Having this in mind, we first show that the proposed upper bound for the degree of the polynomial holds. Then, we show that the lower bound holds as well.
1. Proof for the upper bound: We show that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ is a random polynomial of at most degree 2k12k-1 when $|R|=|C|=k$. The proof will be done by induction.
Base case $k=1$: We have shown that when $k=1$ $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ is a random polynomial in $\widetilde{\alpha}$ of at most degree 1=2k11=2k-1, so the step base holds.
Inductive hypothesis: Assume our proposition holds for diagonal exclusion minors of dimension $(k-1)\times(k-1)$. This implies that for any set of valid $R$ and $C$ we have that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ is a random polynomial in $\widetilde{\alpha}$ of at most degree 2(k1)12(k-1)-1.
[…]
2. Proof for the lower bound: First, note that when $k\leq\bar{r}$ our statement trivially holds because $\max\{0,k-\bar{r}\}=0$ and $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ is a polynomial in $\widetilde{\alpha}$. Then, we are left to show that our statement holds for $k>\bar{r}$. We prove this by induction:
[…]
Inductive hypothesis: Assume our proposition holds for diagonal exclusion minors of dimension $(k-1)\times(k-1)$, with $k-1>\bar{r}$. This implies that for any set of valid $R$ and $C$ we have that $\widetilde{J}_{R,C}(\widetilde{\alpha},\theta^{0})$ is a random polynomial in $\widetilde{\alpha}$ of at least degree $\max\{0,k-1-\bar{r}\}=k-1-\bar{r}$.
[…]
Propositions 1 and 2 provide the essential results for proving Theorem 1. However, before we proceed with the proof, we need to establish one additional result. Consider the following claim:
Claim 3.
Let $\mathbf{M}^{\mathbf{\Omega},k}_{R,C}$ be a diagonal exclusion minor of $\mathbf{\Omega}$ and let $\mathbf{M}^{\mathbf{F\Psi F^{\prime}},k}_{R,C}$ be the corresponding diagonal exclusion minor of $\mathbf{F\Psi F^{\prime}}$. Then $\mathbf{M}^{\mathbf{\Omega},k}_{R,C}=\mathbf{M}^{\mathbf{F\Psi F^{\prime}},k}_{R,C}$.
Proof.
Recall
$\mathbf{\Omega}=\mathbf{F\Psi F^{\prime}+D}$
where $\mathbf{D}$ is a diagonal matrix. Then, one can see that all the non-diagonal terms of $\mathbf{\Omega}$ are equal to the non-diagonal terms of $\mathbf{F\Psi F^{\prime}}$. Thus, as all the terms are equal it is evident that for every diagonal exclusion $k\times k$ minor we have that
$\mathbf{M}^{\mathbf{\Omega},k}_{R,C}=\mathbf{M}^{\mathbf{F\Psi F^{\prime}},k}_{R,C}$
$\square$
From this result, we can derive the following corollary
Corollary 3.
The determinant of every diagonal exclusion minor of $\mathbf{\Omega}$ of dimension $k>\bar{r}$ is equal to 0.
Proof.
This is a direct implication from the previous Claim. First, note that from Claim 3, the determinant of every diagonal exclusion minor of $\mathbf{\Omega}$ is equal to the determinant of the corresponding non-diagonal minor of $\mathbf{F\Psi F^{\prime}}$
$\det\left(\mathbf{M}^{\mathbf{\Omega},k}_{R,C}\right)=\det\left(\mathbf{M}^{\mathbf{F\Psi F^{\prime}},k}_{R,C}\right)$
Now, let $k>\bar{r}$. Then, as $\mathbf{F\Psi F^{\prime}}$ is of rank $\bar{r}$ Lemma 1 implies that
$\det\left(\mathbf{M}^{\mathbf{\Omega},k}_{R,C}\right)=0$
proving the corollary. $\square$
4.3 Proof of Theorem 1
In this subsection, we build on the intermediate results derived in subsection 4.2 to prove Theorem 1. Specifically, we demonstrate that (5) can only hold if $\widetilde{\alpha}=\alpha$, except for a set of factor realizations that occur with probability zero. We begin by outlining the structure of the proof, then proceed with a detailed argument for the case of a single factor to illustrate the reasoning. Finally, we generalize the proof to the case of $\bar{r}$ factors.
Proof of Theorem 1: Pick any diagonal exclusion minor of $\mathbf{O}$ of dimension $\bar{r}+1$, say $\mathbf{M}^{\mathbf{O},\bar{r}+1}_{R,C}$. By Claim 1 the determinant of this minor must satisfy
$\det\left(\mathbf{M}^{\mathbf{O},\bar{r}+1}_{R,C}\right)=0$
By Proposition 1 we know that
$\det\left(\mathbf{M}^{\mathbf{O},\bar{r}+1}_{R,C}\right)=\det\left(\mathbf{M}^{\mathbf{\Omega},\bar{r}+1}_{R,C}\right)+(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)=0$
Moreover, by Corollary 3 we have that $\det\left(\mathbf{M}^{\mathbf{\Omega},\bar{r}+1}_{R,C}\right)=0$. Then:
$\det\left(\mathbf{M}^{\mathbf{O},\bar{r}+1}_{R,C}\right)=(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)=0$ (15)
By Assumption A3 we know that $\mathbf{O}$ has at least $2$ different diagonal exclusion minors. [^4] In particular, when $T=2(\bar{r}+1)$ and $\bar{r}>0$ there are at least $\frac{\binom{2(\bar{r}+1)}{\bar{r}+1}}{2}\geq 2$ diagonal exclusion minors Then, consider any other diagonal exclusion minor of $\mathbf{O}$, say $\mathbf{M}^{\mathbf{O},\bar{r}+1}_{R^{\prime},C^{\prime}}$. Using the same logic, we can see that:
$\det\left(\mathbf{M}^{\mathbf{O},\bar{r}+1}_{R^{\prime},C^{\prime}}\right)=(\alpha-\widetilde{\alpha})\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta^{0}\right)=0$
Then, it must be the case that either $\widetilde{\alpha}=\alpha$ or $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)=\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta^{0}\right)=0$ in order for (5) to hold (Claim 1). However, we show that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)=\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta^{0}\right)=0$ is only possible in a set of probability zero.
Consider the case of a single factor $\bar{r}=1$
Consider the following $2\times 2$ diagonal exclusion minor of $\mathbf{O}$:
$\mathbf{M}_{(1,2),(3,4)}^{\mathbf{O},2}=\begin{bmatrix}O_{1,3}&O_{1,4}\\ O_{2,3}&O_{2,4}\end{bmatrix}$
See Appendix A for the explicit functional form of each $O_{i,j}$. Combining Proposition 1, Corollary 3, and Claim 1 we get that:
$\det\left(\mathbf{M}_{(1,2),(3,4)}^{\mathbf{O},2}\right)=\left(\alpha-\widetilde{\alpha}\right)\Psi(d_{2}-\alpha d_{1}f_{2})\left(f_{3}\widetilde{\alpha}-f_{4}\right)=0$ (16)
where in this case $\widetilde{J}_{(1,2),(3,4)}\left(\widetilde{\alpha},\theta^{0}\right)$ is a random polynomial in $\widetilde{\alpha}$ of degree 1. Note that one of the following four alternatives must be true in order for (16) to be satisfied:
Case 1.
$\Psi=0$
Case 2.
$f_{2}=\frac{d_{2}}{d_{1}\alpha}$
Case 3.
$\widetilde{\alpha}=\frac{f_{4}}{f_{3}}$
Case 4.
$\widetilde{\alpha}=\alpha$
Cases 1 and 2 correspond to the case in which $\widetilde{J}_{(1,2),(3,4)}\left(\widetilde{\alpha},\theta^{0}\right)$ is the zero polynomial, so that all the coefficients are trivially equal to 0. Case 3 corresponds to the case in which we can find a root $\widetilde{\alpha}^{\star}$ that solves $\widetilde{J}_{(1,2),(3,4)}(\widetilde{\alpha}^{\star},\theta^{0})=0$. However, once $\widetilde{\alpha}$ is pinned-down in this way, we have no other free parameter, so the same root must satisfy all the zero-determinant conditions (Claim 1) associated with the remaining diagonal exclusion minors of $\mathbf{O}$. Case 4 corresponds to the case where global identification is achieved.
We show that Cases 1-3 are only possible in a set of factor realizations that occur with probability zero. Thus, with probability 1, Case 4 must hold so $\alpha$ is almost surely globally identified.
Case 1: $\Psi=0$
This case is obvious. We assumed $\Psi>0$, so we can instantly rule this out.
Case 2: $f_{2}=\frac{d_{2}}{d_{1}\alpha}$
From Assumptions A6 and A7, we know that $d_{1},d_{2},\alpha$ are fixed population parameters, whereas $f_{2}$ is drawn from a continuous probability distribution. Thus, we can invoke Lemma 2 and ensure that $P\left(f_{2}=\frac{d_{2}}{d_{1}\alpha}\right)=0$. Then, Case 2 can only occur in a set of probability 0.
Case 3: $\widetilde{\alpha}=\frac{f_{4}}{f_{3}}$
Setting $\widetilde{\alpha}^{\star}=\frac{f_{4}}{f_{3}}$ so that $\widetilde{J}_{(1,2),(3,4)}(\widetilde{\alpha}^{\star},\theta^{0})=0$ leaves us with no other free parameter. Therefore, in order for Case 3 to be possible, it must be the case that this same $\widetilde{\alpha}^{\star}$ simultaneously make all the $\widetilde{J}_{R^{\prime},C^{\prime}}(\widetilde{\alpha}^{\star},\theta^{0})$ equal to 0 for any other valid set of rows and columns (in particular, it requires $\widetilde{J}_{(2,3),(1,4)}(\widetilde{\alpha}^{\star},\theta^{0})=0$). To see why this will not be the case, consider the following $2\times 2$ diagonal exclusion minor of $\mathbf{O}$:
$\mathbf{M}_{(2,3),(1,4)}^{\mathbf{O},2}=\begin{bmatrix}O_{2,1}&O_{2,4}\\ O_{3,1}&O_{3,4}\end{bmatrix}$
Following the same procedure as for (15),
$\det\left(\mathbf{M}_{(2,3),(1,4)}^{\mathbf{O},2}\right)=(\alpha-\widetilde{\alpha})\widetilde{J}_{(2,3),(1,4)}\left(\widetilde{\alpha},\theta^{0}\right)=0$
Algebra calculation shows that $\widetilde{J}_{(2,3),(1,4)}\left(\widetilde{\alpha},\theta^{0}\right)$ is a polynomial of degree 2 in $\widetilde{\alpha}$:
$\widetilde{J}_{(2,3),(1,4)}\left(\widetilde{\alpha},\theta^{0}\right)=a\widetilde{\alpha}^{2}+b\widetilde{\alpha}+c$ (17)
with the coefficients given by:
$a$ $=\alpha d_{1}d_{2}+\alpha d_{2}\Psi+d_{1}f_{2}f_{3}\Psi$
$b$ $=-(d_{1}d_{3}+d_{3}\Psi+\alpha^{2}d_{1}d_{2}+\alpha^{2}d_{2}\Psi+d_{1}f_{3}^{2}\Psi+\alpha^{2}d_{1}f_{2}^{2}\Psi+d_{1}f_{2}f_{4}\Psi+2\alpha d_{1}f_{2}f_{3}\Psi)$
$c$ $=\alpha d_{1}d_{3}+\alpha d_{3}\Psi+d_{3}f_{2}\Psi+\alpha d_{1}f_{3}^{2}\Psi-\alpha d_{2}f_{3}\Psi+d_{1}f_{3}f_{4}\Psi+\alpha^{2}d_{1}f_{2}f_{3}\Psi$
Recall that $\widetilde{J}_{(2,3),(1,4)}\left(\widetilde{\alpha},\theta^{0}\right)=0$ must hold in order for Case 3 to be plausible. There are two ways in which $\widetilde{J}_{(2,3),(1,4)}\left(\widetilde{\alpha},\theta^{0}\right)$ can be 0: either all the coefficients are equal to 0, so that $\widetilde{J}_{(2,3),(1,4)}\left(\widetilde{\alpha},\theta^{0}\right)$ is the zero polynomial; or the coefficients of the polynomial are such that real roots exist and at least one of them coincide with $\widetilde{\alpha}^{\star}=\frac{f_{4}}{f_{3}}$.
We can see that Assumption A7 implies that $a=b=c=0$ can only occur on a set with probability 0, as it involves the factor realizations lying on a particular low-dimensional manifold (see Lemma 2). That is, the first case has zero probability of occurring.
We next show that the second case also has zero probability of occurring. To see this, note that $\widetilde{J}_{(2,3),(1,4)}(\widetilde{\alpha},\theta^{0})$ is a polynomial of degree 2 in $\widetilde{\alpha}$. Its roots are:
$\widetilde{\alpha}=\frac{-b\pm\sqrt{b^{2}-4ac}}{2a}$
Conditional on the roots being real, it must be the case that either
$\frac{f_{4}}{f_{3}}=\frac{-b+\sqrt{b^{2}-4ac}}{2a}\quad\text{or}\quad\frac{f_{4}}{f_{3}}=\frac{-b-\sqrt{b^{2}-4ac}}{2a}$
Examining the expressions of $a$, $b$, and $c$ and making use of Lemma 2 and Assumption A7, we conclude that these events can never occur with positive probability (the factors $f_{2},f_{3},f_{4}$ must lie exactly in a lower dimensional manifold). Hence $\widetilde{\alpha}=\frac{f_{4}}{f_{3}}$ is only possible on a set of zero probability.
Having ruled out Cases 1-3, we are left with Case 4, which is the only option that can happen with positive probability when $\bar{r}=1$, so it must be the case that $\widetilde{\alpha}=\alpha$ almost surely in order for (5) to be satisfied. We have thus established global identification for $\bar{r}=1$. $\square$
Consider the general case with $\bar{r}$ factors
Start from (15)
$\det\left(\mathbf{M}^{\mathbf{O},\bar{r}+1}_{R,C}\right)=(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)=0$
Our goal is to show that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)=0$ for all possible valid combinations of $R$ and $C$ can only occur with probability zero, implying that $\widetilde{\alpha}=\alpha$ must hold almost surely. Recall from Proposition 2 that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ is a random polynomial in $\widetilde{\alpha}$ of degree in between $1$ and $2\bar{r}+1$, given by:
$\widetilde{J}_{R,C}=\sum_{j=1}^{\bar{r}+1}(-1)^{\bar{r}+1+j}\left[J_{r_{\bar{r}+1},c_{j}}\det\left(\mathbf{M}^{\mathbf{\Omega},\bar{r}}_{R-r_{\bar{r}+1},C-c_{j}}\right)+\Omega_{r_{\bar{r}+1},c_{j}}\widetilde{J}_{R-r_{\bar{r}+1},C-c_{j}}+(\alpha-\widetilde{\alpha})J_{r_{\bar{r}+1},c_{j}}\widetilde{J}_{R-r_{\bar{r}+1},C-c_{j}}\right]$
By Assumptions A6 and A7, $\alpha,\mathbf{D}\ \text{and}\ \mathbf{\Psi}$ are deterministic population parameters, and $\mathbf{F}$ is drawn from a continuous distribution. The condition $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)=0$ can occur in two possible ways: either all of the coefficients of this random polynomial are equal to 0, or by setting $\widetilde{\alpha}$ equal to one of the roots of this polynomial $\widetilde{J}_{R,C}\left(\widetilde{\alpha}^{\star},\theta^{0}\right)=0$ (by the fundamental theorem of algebra we know that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ has at most $2\bar{r}+1$ real roots).
The first case is automatically ruled-out by Proposition 2, as it states that $\widetilde{J}_{R,C}(\widetilde{\alpha},\theta^{0})$ is a random polynomial in $\widetilde{\alpha}$ of degree in between $1$ and $2\bar{r}+1$ almost surely. This result guarantees that with probability one $\widetilde{J}_{R,C}(\widetilde{\alpha},\theta^{0})$ is not the zero polynomial.
Now consider the second case. If $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ has no real roots, then it is clear that it must be that $\widetilde{\alpha}=\alpha$ in order for (15) to hold, so the proof would be over. Now, assume $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ has at least one real root. Pick any of these roots, say $\widetilde{\alpha}^{\star}_{R,C}(\theta^{0})$. By definition, $\widetilde{\alpha}^{\star}_{R,C}(\theta^{0})$ solves:
$\widetilde{J}_{R,C}\left(\widetilde{\alpha}^{\star}_{R,C}(\theta^{0}),\theta^{0}\right)=0$
Note that $\widetilde{\alpha}^{\star}_{R,C}(\theta^{0})$ is a function of the other true data-generating process parameters and can be obtained from the above condition. Moreover, it also depends on the rows $R$ and columns $C$ chosen as in general the random polynomials $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ and $\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta^{0}\right)$ are different (except for a set of factor realizations of probability zero).
As we used the diagonal exclusion minor associated with rows and columns $R,C$ to pin down $\widetilde{\alpha}$ we have no other free parameter, so this same $\widetilde{\alpha}^{\star}_{R,C}\left(\theta^{0}\right)$ must satisfy:
$\widetilde{J}_{R^{\prime},C^{\prime}}(\widetilde{\alpha}^{\star}_{R,C}\left(\theta^{0}\right),\theta^{0})=0$
for any other valid set of $R^{\prime},C^{\prime}$. This can only occur in two different ways: either $\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta^{0}\right)$ is the zero polynomial, or the random polynomial $\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta^{0}\right)$ shares at least one root with $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$.
As we have previously discussed, the case in which $\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta^{0}\right)$ is the zero polynomial is an event that occurs with probability zero. Therefore, consider the second case. We know that the coefficients of $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ and $\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta^{0}\right)$ can be exactly equal only in a set of probability zero, so the polynomials are almost surely different. Then, the case in which they share the root $\widetilde{\alpha}^{\star}_{R,C}\left(\theta^{0}\right)$ can be expressed as:
$\begin{split}\widetilde{J}_{R^{\prime},C^{\prime}}(\widetilde{\alpha}^{\star}_{R,C}\left(\theta^{0}\right),\theta^{0})&=\sum_{j=1}^{\bar{r}+1}(-1)^{\bar{r}+1+j}\Bigg[J_{r_{\bar{r}+1}^{\prime},c_{j}^{\prime}}\det\left(\mathbf{M}^{\mathbf{\Omega},\bar{r}}_{R^{\prime}-r_{\bar{r}+1}^{\prime},C^{\prime}-c_{j}^{\prime}}\right)\\ &+\quad\Omega_{r_{\bar{r}+1}^{\prime},c_{j}^{\prime}}\widetilde{J}_{R^{\prime}-r_{\bar{r}+1}^{\prime},C^{\prime}-c_{j}^{\prime}}(\widetilde{\alpha}^{\star}_{R,C}\left(\theta^{0}\right),\theta^{0})\\ &+\quad\left(\alpha-\widetilde{\alpha}^{\star}_{R,C}\left(\theta^{0}\right)\right)J_{r_{\bar{r}+1}^{\prime},c_{j}^{\prime}}\widetilde{J}_{R^{\prime}-r_{\bar{r}+1}^{\prime},C^{\prime}-c_{j}^{\prime}}(\widetilde{\alpha}^{\star}_{R,C}\left(\theta^{0}\right),\theta^{0})\Bigg]=0\end{split}$ (18)
Note that the above condition depends only on the structural parameters of the model. Moreover, it defines a particular low-dimensional manifold in which the random factors $\mathbf{F}$ must lie in order for both polynomials to share the root $\widetilde{\alpha}^{\star}_{R,C}\left(\theta^{0}\right)$. However, by Assumption A7 and Lemma 2, the event in which the random factors lie in this particular low-dimensional manifold occurs with probability zero.
Then, the case $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)=\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta^{0}\right)=0$ can only occur in a set of probability 0. Thus, it must be the case that $\widetilde{\alpha}=\alpha$ in order for (5) to be satisfied. This implies that $\alpha$ is almost surely globally identified. $\square$
4.4 Global identification of $\mathbf{F}$, $\mathbf{\Psi}$, and $\mathbf{D}$
We have shown that $\alpha$ is almost surely globally identified, meaning that (5) is satisfied if and only if $\widetilde{\alpha}=\alpha$ with probability 1. We can then conclude that all the parameters in our model are globally identified. This is stated in Theorem 2. We next provide a proof.
Proof of Theorem 2: As $\alpha$ is almost surely globally identified, then $\widetilde{\alpha}=\alpha$ with probability 1 and $\mathbf{Q}=\mathbf{I}_{T}$. Then, we are left to show:
$\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}+\mathbf{D}=\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}+\mathbf{\widetilde{D}}$
if and only if $\mathbf{\widetilde{D}}=\mathbf{D}$, and $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}=\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}$, where $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}$ must be a $T\times T$ symmetric PSD matrix of rank $\bar{r}$ and $\mathbf{\widetilde{D}}$ a $T\times T$ PD diagonal matrix.
Under the Assumptions of our model, this is the case shown by Anderson and Rubin, (1956), so all the rest of the parameters are almost surely globally identified as well. $\square$
5 Individual fixed-effects
Consider a panel data model with time, individual and interactive effects:
$y_{it}=\alpha y_{it-1}+\delta_{t}+\gamma_{i}+\lambda_{i}^{\prime}f_{t}+\varepsilon_{it}$ (19)
where $\gamma_{i}$ denotes the individual fixed effect for observation $i$, and the rest of the parameters are defined as in (1). It is a well-known result that individual fixed-effects can be absorbed into the factor structure by treating $\gamma_{i}$ as the factor loadings of a special factor, which takes the value of 1 for all $t$. As a result, the factor matrix does not satisfy the conditions of Assumption A7 because the factor associated with the individual fixed-effects follows a degenerate distribution.
Therefore, the results derived in Theorem 1 do not apply in this case. We can overcome this issue in two ways. The first approach consists of showing that our proof can be adapted to allow for one factor with a degenerate distribution. In practice, global identification requires that there is enough variation in some of the factors, but there could be some other factors following degenerate distributions. It is important to note that in this case we do not estimate $\gamma_{i}$ (doing so would cause incidental parameter problems), but we absorb it in the factor structure and identify the variance of $\gamma_{i}$ and the covariance with the rest of the factor loadings through $\mathbf{\Psi}$.
The second approach consists of differencing the data to get rid of the individual fixed effects as in Hayakawa et al., (2023). While Hayakawa et al., (2023) claims that global identification of $\alpha$ cannot be attained in the differenced data model we show that their conjecture is incorrect and that we can use a slightly modified version of our proof to derive global identification when we work in differences.
In practice, we advocate for working with the data in levels. Differencing the data eliminates important information about the process contained when the model is studied in levels. Moreover, it introduces time dependence between the errors as a mechanical result from differencing the idiosyncratic error. It also requires additional periods for identification as a result of both losing one observation by differencing and the temporal dependence between the errors.
5.1 Analysis in levels
We start from the model (19) and absorb the individual fixed effects into the factor structure:
$y_{it}=\alpha y_{it-1}+\delta_{t}+\eta_{i}^{\prime}F_{t}+\varepsilon_{it}$ (20)
where everything is defined as in (1) but $\eta_{i}=[\gamma_{i},\lambda_{i}]$ and $F_{t}=[1,f_{t}]$. Using the same procedure as in Section 2.1, we project the first observation $y_{i1}$ onto the vector $1,\gamma_{i},\lambda_{i}]$, yielding the following expression: $y_{i1}=\delta_{1}^{*}+\gamma_{i}f_{\gamma}^{*}+\lambda_{i}^{\prime}f_{1}^{*}+\varepsilon_{i1}^{*}$, where $(\delta_{1}^{*},f_{\gamma}^{*},f_{1}^{*})$ are the projection coefficients and $\varepsilon_{i1}^{*}$ is the projection residual. For notational simplicity, we drop the asterisks from the coefficients and from the residual, as these are free parameters, and we do not require the error terms $\varepsilon_{it}$ to have identical distributions across $t$. Therefore, for $t=1$, we can rewrite the equation as:
$y_{i1}=\delta_{1}+\gamma_{i}f_{\gamma}+\lambda_{i}^{\prime}f_{1}+\varepsilon_{i1}.$
Combining this with equation (20) and stacking the observations over $t$, we obtain:
$\mathbf{B}y_{i}=\delta+\mathbf{F}\eta_{i}+\varepsilon_{i}$
where $\mathbf{B}$, $y_{i}$, $\delta$, and $\varepsilon_{i}$ are all defined earlier, but
$\mathbf{F}=\begin{bmatrix}f_{\gamma}&f_{1}^{\prime}\\ 1&f_{2}^{\prime}\\ \vdots&\vdots\\ 1&f_{T}^{\prime}\end{bmatrix},\quad\text{ and }\eta_{i}=\begin{bmatrix}\gamma_{i}\\ \lambda_{i}\end{bmatrix}$
Note that the first element in the first column of $\mathbf{F}$ is $f_{\gamma}$, not 1. Importantly, the first column of $\mathbf{F}$ consists of many 1’s, which are fixed constants and follow degenerate distributions. As a result, the factor matrix does not satisfy the conditions of Assumption A7. However, since this assumption is a sufficient but not necessary condition, we show that the model remains globally identified. The key insight is that, as long as there is sufficient variation in some of the factors, identification can still be achieved.
For simplicity, we provide a proof for the global identification of $\alpha$ in the presence of individual fixed-effects and a single unrestricted factor ($dim(f_{t})=1$), and $f_{t}$ is drawn from a continuous distribution. In this case, $\bar{r}=2$. However, the proof can be extended to more general cases encompassing multiple factors.
Note that the presence of individual fixed-effects imposes structure on $\mathbf{F}$, so we require fewer restrictions to eliminate rotational indeterminacy. We only need to impose two restrictions on $\mathbf{F}$ to prevent indeterminacy (see Appendix B for detailed explanation). We shall impose the normalization in the following way, referred to as tail normalization:
$\mathbf{F}=\begin{bmatrix}f_{\gamma}&f_{1}\\ 1&f_{2}\\ \vdots&\vdots\\ 1&f_{T-2}\\ 1&0\\ 1&1\end{bmatrix}$ (21)
This normalization restricts the last two realizations of the factor $f$ to be 0 and 1 respectively. This is purely for convenience in terms of exposition. Normalizing other entries (except the first row) has the same effect. Under this normalization, $\mathbf{F}\mathbf{A}=\mathbf{F}$, if and only if $\mathbf{A}=\mathbf{I}_{2}$, thus eliminating rotational indeterminacy.
Before showing that $\alpha$ can be almost surely identified in this context, we shall explicitly state the identification assumptions:
Assumption B1.
There is one factor associated with the individual fixed-effects and an unrestricted factor ($\bar{r}=2$)
Assumption B2.
$\mathbf{\Psi}$ is an unrestricted symmetric and PD matrix, and $\mathbf{F}$ is normalized as (21)
Assumption B3.
$T\geq 6$.
Assumption B4.
$\varepsilon_{i}$ are i.i.d. over $i$. $\mathbb{E}[\varepsilon_{it}]=0$, $\mathbb{V}(\varepsilon_{it})=d_{t}>0$, and $\varepsilon_{it}$ are independent over $t$. $\varepsilon_{it}$ is independent from the factor loadings $\eta_{i}$ and the factors $\mathbf{F}$.
Assumption B5.
$\eta_{i}$ are either random or fixed constants with sample covariance matrix $\mathbf{\Psi}_{N}$ satisfying $\mathbf{\Psi}_{N}\to\mathbf{\Psi}$ with $\mathbf{\Psi}$ being PD, as $N\to\infty$.
Assumption B6.
$\alpha,d_{1},d_{2},\dots,d_{T},\mathbf{\Psi}$ are fixed population parameters.
Assumption B7.
$f_{\gamma},f_{1},f_{2},\dots f_{T-2}$ are drawn from a continuous probability distribution.
Assumption B8.
There exists a time period $t$ such that $d_{t}\neq\alpha d_{t-1}$ when $\alpha\neq 1$.
Assumptions B1-B7 are an adaptation of the ones outlined in Section 3 for the particular case we are studying. While we conjecture that $f_{\gamma}$ drawn from a continuous distribution is not strictly needed to attain identification, it substantially simplifies the proof exposition.
Assumption B8 is imposed because our strategy is not able to show identification in that particular case (it will become clearer why during the proof). We argue that this does not undermine the generality of our results for two main reasons. First, the case in which $d_{t}=\alpha d_{t-1}$ with $\alpha\neq 1$ is a restriction on the values of the parameters that have Lebesgue measure zero in the parameter space. Second, and more importantly, it is hard to think any true data generating process such that the variance of the idiosyncratic errors should evolve as $d_{t}=\alpha d_{t-1}$ when $\alpha\neq 1$, so this is a case unlikely to be of practical relevance. Note that the case in which $\alpha=1$ and $d_{t}=d_{t-1}$ is identifiable. Thus, when for $\alpha=1$, the model is identifiable under both homoskedasticity and heterokedasticity.
Under these assumptions, we can show that $\alpha$ is almost surely globally identified:
Theorem 3.
Let $y_{it}$ be generated from (19). Then, under Assumptions B1-B8, $\alpha$ is almost surely globally identified.
Proof.
The proof follows in a similar way as in the previous section, as most of the results derived did not depend on the distributional assumption imposed on $\mathbf{F}$. For instance, the definition of $\mathbf{O}$ is the same as before (for the case of $\bar{r}=2$) and Claim 1 still holds, so for any $3\times 3$ diagonal exclusion minor of $\mathbf{O}$ we should have:
$\det\left(\mathbf{M}^{\mathbf{O},3}_{R,C}\right)=0$
for any set of valid rows and columns. Moreover, Proposition 1, Claims 2, 3, and Corollary 3 did not depend on the distributional assumptions imposed on $\mathbf{F}$, so we know that:
$\det\left(\mathbf{M}^{\mathbf{O},3}_{R,C}\right)=(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)=0$
We show that $\alpha$ is almost surely identified when $T=6$. This is a sufficient condition for identification for any bigger $T$, since having more observations can only help identification, not make it worse. Let $T=6$ and $R=(1,2,3)$, $C=(4,5,6)$. The determinant of the diagonal exclusion minor associated with these rows and columns is given by:
[…]
$\quad\times\left(d_{3}f_{1}-\alpha d_{2}f_{1}-d_{3}f_{2}f_{\gamma}+\alpha^{2}d_{1}f_{2}-\alpha^{2}d_{1}f_{3}+\alpha d_{2}f_{3}f_{\gamma}\right)$
$\quad\times(\widetilde{\alpha}f_{4}+1)(\widetilde{\alpha}-1)=0$
where, in this case, $\widetilde{J}_{(1,2,3),(4,5,6)}\left(\widetilde{\alpha},\theta^{0}\right)$ is a random polynomial in $\widetilde{\alpha}$ of degree 2. The proof is similar to the case of $\bar{r}=1$ in Section 4.3. Note that for (22) to hold, one of the following five alternatives must be true:
Case 1’.
$\Psi_{1,2}^{2}=\Psi_{1,1}\Psi_{2,2}$
Case 2’.
$(d_{3}f_{1}-\alpha d_{2}f_{1}-d_{3}f_{2}f_{\gamma}+\alpha^{2}d_{1}f_{2}-\alpha^{2}d_{1}f_{3}+\alpha d_{2}f_{3}f_{\gamma})=0$
Case 3’.
$\widetilde{\alpha}=-\frac{1}{f_{4}}$
Case 4’.
$\widetilde{\alpha}=1$
Case 5’.
$\widetilde{\alpha}=\alpha$
Cases 1’ and 2’ correspond to the case in which $\widetilde{J}_{(1,2,3),(4,5,6)}$ is the zero polynomial, so that all the coefficients are trivially equal to 0. Cases 3’ and 4’ correspond to the case in which we can find a root $\widetilde{\alpha}^{\star}$ that solves $\widetilde{J}_{(1,2,3),(4,5,6)}\left(\widetilde{\alpha}^{\star},\theta^{0}\right)=0$. However, once $\widetilde{\alpha}$ is pinned-down in this way, we have no other free parameter, so the same root must satisfy all the zero-determinant conditions (Claim 1) associated with the remaining diagonal exclusion minors of $\mathbf{O}$. Case 5’ corresponds to the case where global identification is achieved.
We show that Cases 1’-4’ are only possible in a set of factor realizations that occur with probability zero. Thus, with probability 1, Case 5’ must hold so $\alpha$ is almost surely globally identified.
Case 1’: $\Psi_{1,2}^{2}=\Psi_{1,1}\Psi_{2,2}$
This implies that $\det(\mathbf{\Psi})=0$, which is analogous to the case where $\Psi=0$ in subsection 4.3. However, this possibility is excluded because we assume that $\mathbf{\Psi}$ is PD.
Case 2’: $d_{3}f_{1}-\alpha d_{2}f_{1}-d_{3}f_{2}f_{\gamma}+\alpha^{2}d_{1}f_{2}-\alpha^{2}d_{1}f_{3}+\alpha d_{2}f_{3}f_{\gamma}=0$
From Assumptions B6 and B7, we know that $d_{1},d_{2},d_{3},\alpha$ are fixed population parameters, while $f_{1},f_{2},f_{3},f_{\gamma}$ are drawn from a continuous probability distribution. Therefore, we can apply Lemma 2 to conclude that Case 2’ occurs only with probability zero. Note that given that $d_{3}>0$ then the expression cannot be trivially equal to 0.
Cases 3’ and 4’: $\widetilde{\alpha}=-\frac{1}{f_{4}}$ and $\widetilde{\alpha}=1$
We demonstrate that the case $\widetilde{\alpha}=1$ can never satisfy (5) with positive probability. The proof for $\widetilde{\alpha}=-\frac{1}{f_{4}}$ is analogous and thus omitted. If we set $\widetilde{\alpha}=1$ in order to satisfy (22), then we have no other free parameters. Thus, this same $\widetilde{\alpha}$ must satisfy all the zero-determinant conditions (Claim 1). However, this cannot occur with positive probability. To see this, consider the determinant of the diagonal exclusion minor associated with rows $R=(1,5,6)$ and columns $C=(2,3,4)$. By Proposition 1 and Corollary 3, we know it can be written as:
$\det\left(\mathbf{M}^{\mathbf{O},3}_{(1,5,6),(2,3,4)}\right)=(\alpha-\widetilde{\alpha})\widetilde{J}_{(1,5,6),(2,3,4)}\left(\widetilde{\alpha},\theta^{0}\right)$
where $\widetilde{J}_{(1,5,6),(2,3,4)}\left(\widetilde{\alpha},\theta^{0}\right)$ is a random polynomial in $\widetilde{\alpha}$ of degree 4. Case 4’ is plausible if both $\widetilde{J}_{(1,2,3),(4,5,6)}\left(\widetilde{\alpha},\theta^{0}\right)$ and $\widetilde{J}_{(1,5,6),(2,3,4)}\left(\widetilde{\alpha},\theta^{0}\right)$ share the root $\widetilde{\alpha}=1$. Note that this is equivalent to stating that $\widetilde{J}_{(1,5,6),(2,3,4)}\left(\widetilde{\alpha},\theta^{0}\right)|_{\widetilde{\alpha}=1}=0$. Evaluating $\widetilde{J}_{(1,5,6),(2,3,4)}\left(\widetilde{\alpha},\theta^{0}\right)$ at $\widetilde{\alpha}=1$ yields:
$\widetilde{J}_{(1,5,6),(2,3,4)}\left(\widetilde{\alpha},\theta^{0}\right)|_{\widetilde{\alpha}=1}$ $=\left(d_{4}f_{2}-d_{4}f_{3}-\alpha d_{3}f_{2}+\alpha d_{3}f_{4}+\alpha^{2}d_{2}f_{3}-\alpha^{2}d_{2}f_{4}\right)$
$\quad\times\left(f_{\gamma}^{2}\Psi_{1,2}^{2}-f_{\gamma}\Psi_{1,2}^{2}-d_{1}\Psi_{2,2}-f_{\gamma}^{2}\Psi_{1,1}\Psi_{2,2}\right$.
$\quad\left.-\alpha f_{\gamma}^{2}\Psi_{1,2}^{2}+\alpha d_{1}\Psi_{2,2}+f_{\gamma}\Psi_{1,1}\Psi_{2,2}+\alpha f_{\gamma}^{2}\Psi_{1,1}\Psi_{2,2}\right)$
Thus, Case 4’ is plausible if either term is equal to zero. However, this can never occur with positive probability. Let’s begin with the first term. The case in which this term equals zero can be rewritten as:
$(d_{4}-\alpha d_{3})f_{2}-(d_{4}-\alpha^{2}d_{2})f_{3}-(\alpha^{2}d_{2}-\alpha d_{3})f_{4}=0$
This equation defines a plane on which the three-dimensional random variable $(f_{2},f_{3},f_{4})$ must lie. By Lemma 2 and Assumption B8, this is a zero probability event. Now, consider the second term. It can be expressed as a second-degree polynomial in $f_{\gamma}$:
$\mathcal{P}(f_{\gamma})=af_{\gamma}^{2}+bf_{\gamma}+c$
where
$a$ $=(\alpha-1)\left(\Psi_{1,1}\Psi_{2,2}-\Psi_{1,2}^{2}\right)$
$b$ $=\Psi_{1,1}\Psi_{2,2}-\Psi_{1,2}^{2}$
$c$ $=(\alpha-1)d_{1}\Psi_{2,2}$
Setting $\mathcal{P}(f_{\gamma})$ equal to zero is equivalent to finding the roots of this polynomial. First, note that $\mathcal{P}$ can never be the zero polynomial since $b>0$, which holds because $\mathbf{\Psi}$ is PD by assumption. Then, by the fundamental theorem of algebra, there are at most two real values of $f_{\gamma}$ that satisfy $\mathcal{P}(f_{\gamma})=0$. Since $f_{\gamma}$ is assumed to be drawn from a continuous distribution, the event $\mathcal{P}(f_{\gamma})=0$ has zero probability, because these two roots have zero Lebesgue measure in the support of $f_{\gamma}$.
Thus, we can conclude that almost surely, Case 4’ is not plausible. A similar argument leads to the conclusion that Case 3’ can only occur with probability zero. We have shown that Case 5’ is the only alternative that makes $\det\left(\mathbf{M}^{\mathbf{O},3}_{(1,2,3),(4,5,6)}\right)=0$ with positive probability. This implies that $\widetilde{\alpha}=\alpha$ is a necessary condition for (5) to hold in the presence of individual fixed effects. Thus, $\alpha$ is almost surely globally identified.
Remark. We have proven identification for the case when $T=6$. Larger $T$ only strengthens the identification. One way to see this is by applying the same normalization used for the case of $T=6$ even when $T>6$. This corresponds to a non-tail normalization. That is, the transpose of $\mathbf{F}$ is written as:
$\mathbf{F}^{\prime}=\begin{bmatrix}f_{\gamma}&1&1&1&1&1&1&\cdots&1\\ f_{1}&f_{2}&f_{3}&f_{4}&0&1&f_{7}&\cdots&f_{T}\end{bmatrix}$
Our entire proof carries over without any modification. $\square$
Given global identification of $\alpha$, it becomes clear that the rest of the parameters are almost surely globally identified as well. This can be formalized in the following Theorem.
Theorem 4.
Let $y_{it}$ be generated from (19). Then, under Assumptions B1-B8, all the parameters in the model are almost surely globally identified.
Proof.
Identical to the Proof of Theorem 2 in Section $4.$ $\square$
5.2 Analysis in differences
Another alternative to dealing with the presence of individual fixed effects is to difference the data as in Hayakawa et al., (2023). While the authors claim that global identification is not necessarily attainable when we consider the data in differences, we show that our proof and results can be adapted to achieve global identification of $\alpha$ in this context. To see this, consider a re-labeled version of the DGP studied in the previous subsection: [^5] The DGP is exactly equivalent to the one considered in the previous subsection and in Hayakawa et al., (2023). The choice of notation is arbitrary and done for convenience.
$z_{it}=\alpha z_{it-1}+\nu_{t}+\gamma_{i}+\lambda_{i}^{\prime}g_{t}+u_{it}$ (23)
where everything is defined exactly as in (19), except that we assume that $u_{it}$ is homoskedastic with variance $\sigma^{2}<\infty$. [^6] Hayakawa et al., (2023) assumes that $u_{it}$ is homoskedastic in their proof, so we follow this assumption in order to simplify the comparison and exposition. However, the proof would work in the same way if $u_{it}$ had heteroskedastic variance $V(u_{it})=\sigma_{t}^{2}$ as in the previous sections of the paper. Now, consider the model for the differenced data:
$\Delta z_{it}=\alpha\Delta z_{it-1}+\Delta\nu_{t}+\lambda_{i}^{\prime}\Delta g_{t}+\Delta u_{it}$
Let:
$y_{it}=\Delta z_{it},\quad\delta_{t}=\Delta\nu_{t},\quad f_{t}=\Delta g_{t},\quad\varepsilon_{it}=\Delta u_{it}$
And write the model for the differenced data as:
$y_{it}=\alpha y_{it-1}+\delta_{t}+\lambda_{i}^{\prime}f_{t}+\varepsilon_{it}$
We can see that this model looks exactly like the one we have considered throughout the paper derived from (1) with the only difference that $\varepsilon_{it}$ are no longer independent over $t$. The reason is that, as $u_{it}$ was assumed to be a white noise process, then its first difference follows a non-invertible MA process. Aside from that, the models are equal. Thus, we can project the initial observation of the differenced process $y_{i1}$ onto $1,\lambda_{i}]$ and stack observations over $t$ the same way we did in Section 2 and write:
$y_{i}=\mathbf{\Gamma}\delta+\mathbf{\Gamma}\mathbf{F}\lambda_{i}+\mathbf{\Gamma}\varepsilon_{i}$
We are in a situation very similar to the one we considered in Section 4, with the only difference that the covariance matrix of the idiosyncratic errors differs. In the case considered in Section 4:
$\mathbf{D}=\mathbb{E}[\varepsilon_{i}\varepsilon_{i}^{\prime}]=\begin{bmatrix}d_{1}&0&\dots&0\\ 0&d_{2}&\dots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&\dots&0&d_{T}\end{bmatrix}$
whereas for the differenced data the covariance matrix is no longer diagonal and it equals:
$\mathbf{\dot{D}}=\mathbb{E}[\varepsilon_{i}\varepsilon_{i}^{\prime}]=\begin{bmatrix}\sigma_{1}^{2}&\sigma_{c}&0&\dots&0\\ \sigma_{c}&2\sigma^{2}&-\sigma^{2}&\dots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&0&-\sigma^{2}&2\sigma^{2}&-\sigma^{2}\\ 0&0&0&-\sigma^{2}&2\sigma^{2}\end{bmatrix}$
a tridiagonal matrix. The parameters $\sigma_{1}^{2}$ and $\sigma_{c}$ capture the fact that the initial idiosyncratic error $\varepsilon_{i1}$ may follow a different distribution from the subsequent errors. Specifically,
$\sigma_{1}^{2}=\mathbb{E}[\varepsilon_{i1}^{2}],\qquad\sigma_{c}=\mathbb{E}[\varepsilon_{i1}\varepsilon_{i2}]=\mathbb{E}[\varepsilon_{i1}\Delta u_{i2}]=-\,\mathbb{E}[\varepsilon_{i1}u_{i1}].$
For any $t\geq 3$, we have $\mathbb{E}[\varepsilon_{i1}\varepsilon_{it}]=\mathbb{E}[\varepsilon_{i1}\Delta u_{it}]=0$, implying that the projection of the initial observation affects only the leading $2\times 2$ block of $\mathbf{\dot{D}}$.
Then, the expected value of the sample variance of the differenced data is equal to:
$\mathbb{E}[\mathbf{S}_{N}]=\mathbf{\dot{\Sigma}}(\theta^{0})=\mathbf{\Gamma}(\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}+\mathbf{\dot{D}})\mathbf{\Gamma}^{\prime}$ (24)
which is equal to the one derived in Section 2 changing $\mathbf{D}$ for $\mathbf{\dot{D}}$. Thus, showing global identification in the differenced model is equivalent to showing:
$\mathbf{\Gamma}(\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}+\mathbf{\dot{D}})\mathbf{\Gamma}^{\prime}=\mathbf{\widetilde{\Gamma}}(\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}+\mathbf{\widetilde{\dot{D}}})\mathbf{\widetilde{\Gamma}}^{\prime}$ (25)
if and only if $\mathbf{\widetilde{\dot{D}}}=\mathbf{\dot{D}}$, $\widetilde{\alpha}=\alpha$, and $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}=\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}$.
While the proof derived in previous sections does not directly apply to this case, it can be adapted very easily. For instance, throughout the paper, we used the fact that $\mathbf{D}$ was diagonal to isolate the identification problem of $\alpha$ from the one of the rest of the parameters. In the case of the differenced data, the counterpart of the matrix $\mathbf{D}$, $\mathbf{\dot{D}}$, is not diagonal, so we cannot use the same argument. However, it is easy to note that the matrix:
$\mathbf{\dot{O}}\equiv(\mathbf{\widetilde{\Gamma}}^{-1}\mathbf{\Gamma})(\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}+\mathbf{\dot{D}})(\mathbf{\widetilde{\Gamma}}^{-1}\mathbf{\Gamma})^{\prime}-\mathbf{\widetilde{\dot{D}}}$ (26)
is such that the elements outside the main diagonal and the adjacent diagonals do not depend on $\mathbf{\widetilde{\dot{D}}}$. In particular, the tridiagonal exclusion minors of $\mathbf{\dot{O}}$ only depend on the true DGP parameters $\theta^{0}$ and $\widetilde{\alpha}$.
Definition 2.
We refer to tridiagonal exclusion minors as minors that do not include any entries from the tridiagonal band (the main diagonal together with the first super- and subdiagonal) of the original matrix. Formally, let $\mathbf{A}$ be a $T\times T$ matrix, and define $\mathbf{\dot{M}}^{\mathbf{A},k}_{R,C}$ as a tridiagonal exclusion minor of $\mathbf{A}$ of dimension $k$, consisting of rows $R=\{r_{1},r_{2},\dots,r_{k}\}$ and columns $C=\{c_{1},c_{2},\dots,c_{k}\}$.
These restrictions ensure that no entry of the tridiagonal band (main diagonal, first super-, or subdiagonal) is included. The properties of tridiagonal exclusion minors of $\mathbf{\dot{O}}$ are the same as those derived for diagonal exclusion minors of $\mathbf{O}$ in Section 4. The change of $\mathbf{D}$ to $\mathbf{\dot{D}}$ does not affect the essence of the result, it simply requires re-defining some of the objects. Thus, our proof follows smoothly for tridiagonal exclusion minors if we impose similar assumptions on the parameters as we did before. Let $T$ denote the dimension of the data in levels, so that we have T1T-1 periods when working in differences. The assumptions required for almost sure identification in the differenced data are:
Assumption C1.
The number of distinct factors is known and equal to $\bar{r}$
Assumption C2.
$\mathbf{\Psi}$ is an unrestricted symmetric and PD matrix, and $\mathbf{F}=(\mathbf{I}_{\bar{r}},\mathbf{F}_{2}^{\prime})^{\prime}$
Assumption C3.
$T\geq 2(\bar{r}+1)+3$
Assumption C4.
$u_{it}$ are i.i.d. over $i$. $\mathbb{E}[u_{it}]=0$, $\mathbb{V}(u_{it})=\sigma^{2}>0$, and $u_{it}$ are independent over $t$. $u_{it}$ is independent from the factor loadings $\lambda_{i}$ and the factors $\mathbf{F}$.
Assumption C5.
$\lambda_{i}$ are either random or fixed constants with $\mathbf{\Psi}_{N}\to\mathbf{\Psi}$ as $N\to\infty$.
Assumption C6.
$\alpha,\sigma^{2},\sigma_{1}^{2},\sigma_{c},\mathbf{\Psi}$ are fixed population parameters.
Assumption C7.
Let $\mathbf{F}_{2}=\begin{bmatrix}f_{\bar{r}+1,1}&f_{\bar{r}+1,2}&\dots&f_{\bar{r}+1,\bar{r}}\\ \vdots&\vdots&\vdots&\vdots\\ f_{T,1}&f_{T,2}&\dots&f_{T,\bar{r}}\end{bmatrix}$ be the random elements of $\mathbf{F}$. Then, $\mathbf{F}_{2}$ is a continuous random matrix with support in $\Pi\in\mathbb{R}^{(T-\bar{r})\times\bar{r}}$.
The assumptions are basically the same we imposed in Section 3 but applied to the differenced processes. The only major difference appears in the time periods required for identification (Assumption C3). We need three more periods to achieve identification in this case than in the baseline case of the paper. There are two reasons for this. First, differencing the data mechanically reduces the number of observations by 1, so in order to have $T$ observations for the differenced model there must be $T+1$ observations in levels.
The second reason is that in order to have $(\bar{r}+1)\times(\bar{r}+1)$ tridiagonal exclusion minors the original matrix should at least be of dimension $2(\bar{r}+1)+1$. However, when $T=2(\bar{r}+1)+1$ we can only have one tridiagonal exclusion minor because $\mathbf{\dot{O}}$ is symmetric (see Appendix C for a proof), and our identification strategy requires having at least two different tridiagonal exclusion minors. Thus, we need to have at least one additional observation to be able to build many tridiagonal exclusion minors. For example, if there is a single factor $\bar{r}=1$ and we want to build $2\times 2$ tridiagonal exclusion minors of $\mathbf{\dot{O}}$, we should have that $dim(\mathbf{\dot{O}})\geq 6$. In our original proof we only needed $dim(\mathbf{O})\geq 4$. In general, this adds two additional observations. Once this is taken into account, we can use the techniques employed in the previous sections to show that the autoregressive coefficient $\alpha$ is almost surely identified when we consider the differenced data:
Theorem 5.
Let $y_{it}$ be generated from (23). Then, under Assumptions C1-C7, $\alpha$ is almost surely globally identified.
Proof.
The entries outside the tridiagonal band of $\mathbf{\dot{O}}$ are defined in the same way as the entries outside the tridiagonal band of $\mathbf{O}$ substituting $\mathbf{D}$ for $\mathbf{\dot{D}}$. The structure of $\mathbf{D}$ was only instrumental for Claims 2 and 3, and Corollary 1 and 3. When we substitute $\mathbf{D}$ for $\mathbf{\dot{D}}$ all these results hold analogously if we consider tridiagonal exclusion minors instead of diagonal exclusion minors. This implies that if $T\geq 2(\bar{r}+1)+3$ we can write the determinant of every tridiagonal exclusion minor as:
$\det\left(\mathbf{\dot{M}}^{\mathbf{\dot{O}},\bar{r}+1}_{R,C}\right)=(\alpha-\widetilde{\alpha})\widetilde{\dot{J}}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)=0$
Provided that $\dim(\mathbf{\dot{O}})\geq 2(\bar{r}+1)+2$ we can construct (more than) two tridiagonal exclusion minors of dimension $(\bar{r}+1)\times(\bar{r}+1)$. Then, the rest of the proof follows as in Theorem 1. $\square$
This result contradicts the claim stated in Hayakawa et al., (2023) regarding the impossibility of attaining global identification in dynamic panel models additive time and fixed effects, as well as interactive effects. Moreover, Williams, (2020) extends Anderson and Rubin, (1956) to more general structures of dependence between the errors, not requiring their covariance matrix to be diagonal. We can use their results to show that the rest of the parameters are globally identified as well.
Theorem 6.
Let $y_{it}$ be generated from (23). Then, under Assumptions C1-C7, all the parameters in the model are almost surely globally identified.
Proof.
Once we have shown that $\widetilde{\alpha}=\alpha$ is a necessary condition for (25) to hold, we are left to show that:
$\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}+\mathbf{\dot{D}}=\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}+\mathbf{\widetilde{\dot{D}}}$
if and only if $\mathbf{\widetilde{\dot{D}}}=\mathbf{\dot{D}}$, and $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}=\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}$.
This is shown in Theorem 2.1 of Williams, (2020) $\square$
5.3 Dynamic panel with Individual Fixed Effects
Consider a dynamic panel with individual fixed effects as in the canonical work of Arellano and Bond, (1991):
$y_{it}\;=\;\alpha\,y_{it-1}\;+\;\gamma_{i}\;+\;\varepsilon_{it},\qquad t\geq 2,$ (27)
where $\varepsilon_{it}$ are i.i.d. across $i$, independent of $\gamma_{i}$, satisfy $\mathbb{E}[\varepsilon_{it}]=0$, and have $\mathbb{V}(\varepsilon_{it})=d_{t}$.
In Arellano and Bond, (1991), the restriction $|\alpha|<1$ is imposed. Subsequent work argues that $\alpha$ is not identified in the unit-root case based on their estimation method (see Blundell and Bond, (1998) and Sentana, (2024)). In this subsection we show that, in fact, $\alpha$ is almost surely identified for any value, including $\alpha=1$.
To proceed, we augment (27) with time effects:
$y_{it}\;=\;\alpha\,y_{it-1}\;+\;\gamma_{i}\;+\;\delta_{t}\;+\;\varepsilon_{it}.$ (28)
As in previous sections, we treat the first observation separately and project $y_{i1}$ onto the span of $1,\gamma_{i}]$:
$y_{i1}\;=\;\delta_{1}^{*}\;+\;f_{\gamma}\,\gamma_{i}\;+\;\varepsilon^{*}_{i1}$
where $(\delta_{1}^{*},f_{\gamma})$ are projection coefficients and $\varepsilon^{*}_{i1}$ is orthogonal to $\gamma_{i}$ by construction. Since we place no restrictions on the distribution of $\varepsilon_{i1}$, we simply relabel $(\delta_{1}^{*},\varepsilon_{i1}^{*})$ as $(\delta_{1},\varepsilon_{i1})$ without loss of generality.
We include time effects $\delta_{t}$ for generality, even if our method does not rely on time effects to identify the model. If (27) is the true DGP, one may impose $\delta_{t}=0$ for $t\geq 2$ (and retain a free $\delta_{1}$ arising from the projection at $t=1$) to improve efficiency. In any case, $\delta_{t}$ can be concentrated out as in Section 2.2, and our identification arguments are based on second-moment conditions that are invariant to this step.
As we have outlined in subsection 5.1, we can write a dynamic panel model with individual fixed effects as an interactive fixed-effects model with a unique factor given by $\mathbf{F}=[f_{\gamma},1,\dots,1]^{\prime}$ and factor loadings given by $\gamma_{i}$ (this is a particular case of the model discussed in subsection 5.1 when $\bar{r}=1$ and the only factor corresponds to the individual fixed-effects). Then, we can stack observations over $t$ as before and write:
$y_{i}=\mathbf{\Gamma}\delta+\mathbf{\Gamma}\mathbf{F}\gamma_{i}+\mathbf{\Gamma}\varepsilon_{i}$
where everything is defined as before except that now $\mathbf{F}=[f_{\gamma},1,\dots,1]^{\prime}$ and $\gamma_{i}$ is a scalar containing the individual fixed effect of observation $i$. Also note that the matrix $\mathbf{\Psi}$ is now a scalar $\Psi$ corresponding to the population variance of the individual fixed effects $\mathbb{V}(\gamma_{i})$. The assumptions required for identification are:
Assumption D1.
$T\geq 4$
Assumption D2.
$\varepsilon_{i}$ are i.i.d. over $i$. $\mathbb{E}[\varepsilon_{it}]=0$, $\mathbb{V}(\varepsilon_{it})=d_{t}>0$, and $\varepsilon_{it}$ are independent over $t$. $\varepsilon_{it}$ is independent from the individual effects $\gamma_{i}$ and the factors $\mathbf{F}$.
Assumption D3.
$\gamma_{i}$ are either random or fixed constants with $\Psi_{N}\to\Psi>0$ as $N\to\infty$.
Assumption D4.
$\alpha,d_{1},d_{2},\dots,d_{T},\Psi$ are fixed population parameters.
Assumption D5.
$f_{\gamma}$ is drawn from a continuous probability distribution
Remark: Assumption D5 is just a technical device to simplify exposition. Another way of seeing this, is that the parameters of the model are globally identified except for a set of projection coefficients of the initial observation that have Lebesgue measure zero in the support of $f_{\gamma}$. This means that unless very particular and strict conditions are imposed on the initial observation, any value of $\alpha$ can be identified in a dynamic panel with individual effects, including $\alpha=1$. Moreover, we also conjecture that this assumption is not strictly needed to attain identification, but it substantially simplifies the proof exposition without having major consequences in terms of generalization.
Under these assumptions $\alpha$ can be almost surely globally identified:
Theorem 7.
Let $y_{it}$ be generated from (27). Then, under Assumptions D1-D5 $\alpha$ is almost surely globally identified.
Proof.
In a dynamic panel model with individual effects the expected value of the sample covariance matrix of $y_{i}$ is then given by
$\mathbb{E}[\mathbf{S}_{N}]=\mathbf{\Sigma}(\theta^{0})=\mathbf{\Gamma}(\mathbf{F}\Psi\mathbf{F}^{\prime}+\mathbf{D})\mathbf{\Gamma}^{\prime}$
where $\theta^{0}=(\alpha,\mathbf{F},\Psi,\mathbf{D})$ is the true vector of structural parameters. The matrix $\mathbf{\Sigma}$ is a $T\times T$ symmetric PD matrix, $\Psi$ is a positive scalar corresponding to the cross-sectional variance of the individual effects, $\mathbf{F}\Psi\mathbf{F}^{\prime}$ is a $T\times T$ symmetric positive semi-definite (PSD) matrix of rank $1$, and $\mathbf{D}$ is a $T\times T$ diagonal PD matrix.
[…]
Note that most of the results derived in Section 4 can be applied with no modifications for the case in which we have a single factor given by $\mathbf{F}=[f_{\gamma},1,\dots,1]$, as they did not require any distributional assumption on $\mathbf{F}$. In particular, we can define $\mathbf{O}$ in the same way as before and Claim 1 still implies that the determinant of every diagonal exclusion minor of $\mathbf{O}$ of dimension $2\times 2$ must be equal to 0 (we only have $\bar{r}=1$ factor given by the individual effects). Moreover, it will still be the case that we can write the determinant of every diagonal exclusion minor of $\mathbf{O}$ of dimension $2\times 2$ as:
$\det\left(\mathbf{M}_{R,C}^{\mathbf{O},2}\right)=(\alpha-\widetilde{\alpha})\widetilde{J}(\widetilde{\alpha},\theta^{0})=0$
Then, the proof of this result follows a similar logic than the proof outlined in Section 4 for the case of a single factor. However, we need to distinguish between the case when $\alpha=1$ from the case when $\alpha\neq 1$. Paradoxically, $\alpha$ can be very easily identified when $\alpha=1$, contrary to the canonical result of Blundell and Bond, (1998). However, the difficulty relies in showing that alternative models when $\alpha\neq 1$ do not allow an alternative representation with $\widetilde{\alpha}=1$. This will become clearer later on.
Case 1: $\alpha=1$
Consider the diagonal exclusion minor of $\mathbf{O}$ associated with rows $R=(1,2)$ and columns $C=(3,4)$. We can show that:
$\det\left(\mathbf{M}_{(1,2),(3,4)}^{\mathbf{O},2}\right)=\Psi(\widetilde{\alpha}-1)^{2}(d_{1}-d_{2}f_{\gamma})=0$ (29)
Thus, in order for (29) to be satisfied it must that either:
Case 1”.
$\Psi=0$
Case 2”.
$f_{\gamma}=\frac{d_{1}}{d_{2}}$
Case 3”.
$\widetilde{\alpha}=\alpha=1$
Case 1” is ruled out by assumption, as we have assumed that $\mathbb{V}(\gamma_{i})>0$. The case in which we allow $\mathbb{V}(\gamma_{i})=0$ (i.e. $\gamma_{i}=\gamma\ \forall\ i$) can also be identified, but we omit the discussion here for simplicity.
Case 2” can only occur in a set of probability 0 because we have assumed that $f_{\gamma}$ is drawn from a continuous probability distribution, so by Lemma 2 the case in which $f_{\gamma}=\frac{d_{1}}{d_{2}}$ is an event of probability 0.
Case 3” is the only alternative making the determinant equal to zero with positive probability. Then, it must be the case that $\widetilde{\alpha}=1$ almost surely. This shows that $\alpha$ can be almost surely globally identified if $\alpha=1$$\square$
Case 2: $\alpha\neq 1$
When $\alpha\neq 1$ the determinant of the minor associated with rows $R=(1,2)$ and columns $C=(3,4)$ can be written as:
$\det\left(\mathbf{M}_{(1,2),(3,4)}^{\mathbf{O},2}\right)=\Psi(\widetilde{\alpha}-\alpha)(\widetilde{\alpha}-1)(\alpha d_{1}-d_{2}f_{\gamma})=0$ (30)
The cases where $\Psi=0$ or $f_{\gamma}=\frac{\alpha d_{1}}{d_{2}}$ can be ruled-out in an analogous way as in the previous case. The difficulty here is ruling out that there exists an alternative parametrization such that $\widetilde{\alpha}=1$. Interestingly, diagonal exclusion minors built using other rows and columns do not provide useful information in this regard, so we need to resort to other methods.
[…]
$\begin{cases}(d_{1}+\Psi f_{\gamma}^{2})-(\widetilde{d_{1}}+\widetilde{\Psi}\widetilde{f}_{\gamma}^{2})=0\\ (\alpha(\Psi f_{\gamma}^{2}+d_{1})+\Psi f_{\gamma})-((\widetilde{\Psi}\widetilde{f}_{\gamma}^{2}+\widetilde{d}_{1})+\widetilde{\Psi}\widetilde{f}_{\gamma})=0\\ (\alpha^{2}(\Psi f_{\gamma}^{2}+d_{1})+\alpha\Psi f_{\gamma}+\Psi f_{\gamma})-((\widetilde{\Psi}\widetilde{f}_{\gamma}^{2}+\widetilde{d}_{1})+2\widetilde{\Psi}\widetilde{f}_{\gamma})=0\end{cases}$ (31)
This is a nonlinear system of three equations in three parameters $(\widetilde{d}_{1},\widetilde{\Psi},\widetilde{f}_{\gamma})$. However, this system admits a solution only for a set of realizations of $f_{\gamma}$ of probability 0. To see this, note that we can re-write the first equation as:
$\widetilde{d}_{1}=d_{1}+\Psi f_{\gamma}^{2}-\widetilde{\Psi}\widetilde{f}_{\gamma}^{2}$
Plugging into the second equation yields:
$\alpha(\Psi f_{\gamma}^{2}+d_{1})+\Psi f_{\gamma}-\widetilde{\Psi}\widetilde{f}_{\gamma}^{2}-(d_{1}+\Psi f_{\gamma}^{2}-\widetilde{\Psi}\widetilde{f}_{\gamma}^{2})-\widetilde{\Psi}\widetilde{f}_{\gamma}=0$
Rearranging in terms of $\widetilde{\Psi}\widetilde{f}_{\gamma}$:
$\widetilde{\Psi}\widetilde{f}_{\gamma}=\Psi f_{\gamma}-(1-\alpha)(d_{1}+\Psi f_{\gamma}^{2})$
Finally, substituting $\widetilde{d}_{1}$ and $\widetilde{\Psi}\widetilde{f}_{\gamma}$ into the third equation and rearranging yields:
$(\alpha-1)\left((\alpha-1)(d_{1}+\Psi f_{\gamma}^{2})+\Psi f_{\gamma}\right)=0$
Note that the final condition only depend on the true structural parameters $(\alpha,d_{1},\Psi,f_{\gamma})$. It defines a manifold in which these parameters must lie in order for an alternative solution with $\widetilde{\alpha}=1$ can exist when $\alpha\neq 1$. However, this is a low dimensional manifold with respect to the dimension of these parameters. Concretely, given that we are in the case in which $\alpha\neq 1$ by the fundamental theorem of algebra there are at most two values of $f_{\gamma}$ in the real line under which $\widetilde{\alpha}=1$ can potentially be a solution. Under Assumption D5 and $\alpha\neq 1$ this is an event of probability 0, so this case can be ruled-out.
Then, we have shown that when $\alpha\neq 1$ it must be the case that $\widetilde{\alpha}=\alpha$ almost surely in order for (30) to hold. This shows that $\alpha$ is almost surely globally identified.
$\square$
Once $\alpha$ is identified the rest of the parameters are identified as well:
Theorem 8.
Let $y_{it}$ be generated from (27). Then, under Assumptions D1-D5, all the parameters in the model are almost surely globally identified.
Proof.
Identical to the Proof of Theorem 2 in Section $4.$ $\square$
5.4 Possible extensions
The results derived in this paper can be used to establish identification in a wide range of dynamic models featuring time, individual, or interactive effects. The techniques of the proof we have introduced can be used to accommodate models with more sophisticated structures in the error term than a heteroskedastic white noise and provide a novel framework for addressing identification in these settings.
For example, going back to the baseline model of the paper (1), an immediate implication of the results derived in the previous subsections is that if the error $\varepsilon_{it}$ is an MA(q) process then almost sure global identification of $\alpha$ can be derived using minors that exclude all entries within the band formed by the main diagonal and the first $q$ super- and subdiagonals of $\mathbf{O}$ provided that the time dimension is big enough.
More generally, we believe that some modifications of our original proof can be used to show identification in a wide range of dynamic models where the covariance between the error terms feature some level of sparsity (i.e. the matrix $\mathbf{D}$ has ‘enough’ zeros).
6 Concluding remarks
In this paper, we demonstrate that the parameters of a dynamic panel data model with interactive effects can be globally identified. We extend the seminal result of Anderson and Rubin, (1956) to dynamic panels with interactive effects, a problem that was previously thought to be non-globally identifiable.
The significance of our result lies in the fact that, despite the high degree of nonlinearity in the problem, when the data-generating process corresponds to the specified model, there exists a one-to-one correspondence between the population moments and the model parameters.
Our proof shows that global identification is achievable when $T\geq 2(\bar{r}+1)$. An interesting open question is whether global identification can be attained under weaker conditions, such as when $T$ satisfies the Ledermann bound plus one. The requirement for an additional period, beyond the Ledermann bound, arises from the dynamic nature of the model, which includes an extra parameter compared to the static case.
Appendix A Explicit expression for $O_{i,j}$ in Section 4 and $\bar{r}=1$
As $\mathbf{O}$ is symmetric, then $O_{i,j}=O_{j,i}$.The positions of $\mathbf{O}$ used in section 4.3 are:
$O_{1,2}$ $=\left(d_{1}+\Psi\right)\,\left(\alpha-\widetilde{\alpha}\right)+f_{2}\Psi$
[…]
$\quad-f_{2}\,\Psi\,\left(\alpha\,\widetilde{\alpha}-\alpha^{2}\right)$
$O_{2,3}$ $=\left(\alpha-\widetilde{\alpha}\right)\,\left(\Psi\,{f_{2}}^{2}+\Psi\,\left(\alpha-\widetilde{\alpha}\right)\,f_{2}+d_{2}\right)$
$\quad-\left(\alpha\,\widetilde{\alpha}-\alpha^{2}\right)\,\left(\left(d_{1}+\Psi\right)\,\left(\alpha-\widetilde{\alpha}\right)+f_{2}\,\Psi\right)$
[…]
$\quad-\left(\alpha\,\widetilde{\alpha}-\alpha^{2}\right)\,\left(\Psi\,{f_{2}}^{2}+\Psi\,\left(\alpha-\widetilde{\alpha}\right)\,f_{2}+d_{2}\right)$
[…]
$\quad-\left(\alpha\,\widetilde{\alpha}-\alpha^{2}\right)\,\left(\left(\Psi\,{f_{2}}^{2}+d_{2}\right)\,\left(\alpha-\widetilde{\alpha}\right)+f_{2}\,f_{3}\,\Psi-f_{2}\,\Psi\,\left(\alpha\,\widetilde{\alpha}-\alpha^{2}\right)\right)$
[…]
Appendix B Rotational indeterminacy in the presence of individual fixed-effects and $\bar{r}=2$
In general, factor models exhibit rotational indeterminacy, meaning that identification is possible up to a $\bar{r}\times\bar{r}$ rotation matrix $\mathbf{A}$. The presence of individual fixed effects imposes a restriction on the possible values of the factor matrix $\mathbf{F}$. We explain why, in the case of individual fixed effects and $\bar{r}=2$, only two restrictions are needed to eliminate rotational indeterminacy, as opposed to $\bar{r}^{2}=4$ restrictions in the unrestricted case. Let $\mathbf{A}$ be an invertible $2\times 2$ matrix, defined as:
$\mathbf{A}=\begin{bmatrix}a&b\\ c&d\\ \end{bmatrix}$
Next, recall the factor structure with individual fixed effects:
$\mathbf{F}=\begin{bmatrix}f_{\gamma}&f_{1}\\ 1&f_{2}\\ \vdots&\vdots\\ 1&f_{T}\\ \end{bmatrix},\quad\eta_{i}=\begin{bmatrix}\gamma_{i}\\ \lambda_{i}\end{bmatrix}$
Rotational indeterminacy implies that:
$\mathbf{F}\eta_{i}=\mathbf{F}\mathbf{A}\mathbf{A}^{-1}\eta_{i}$
However, for $\mathbf{F}\mathbf{A}$ to maintain the structure of $\mathbf{F}$ (which includes a column of nearly all elements equal to one), we immediately conclude that $a=1$ and $c=0$, leaving only two free parameters, $b$ and $d$. Hence, only two restrictions are required on $\mathbf{F}$ to resolve rotational indeterminacy.
Under the tail normalization discussed in the main text, we implicitly assume that the last $2\times 2$ block of $\mathbf{F}$ is an invertible matrix, which is equivalent to the condition $f_{T-1}\neq f_{T}$. The case where $f_{T-1}=f_{T}$ is considered a zero-probability event by assumption.
Appendix C Number of unique tridiagonal exclusion minors
Proposition 3.
If $T=2(\bar{r}+1)+1$, then there exists exactly one unique tridiagonal exclusion minor of dimension $\bar{r}+1$ of $\mathbf{\dot{O}}$.
Proof.
A $(\bar{r}+1)\times(\bar{r}+1)$ tridiagonal exclusion minor is defined by two index sets
$R=\{r_{1},\dots,r_{\bar{r}+1}\},\qquad C=\{c_{1},\dots,c_{\bar{r}+1}\},$
with $|R|=|C|=\bar{r}+1$, such that every selected entry lies outside the tridiagonal band, i.e.
$|r-c|\geq 2\quad\text{for all }r\in R,\ c\in C.$
For a fixed $R$, define the tridiagonal neighborhood
$N(R):=\{\,r-1,\,r,\,r+1:r\in R\,\}\cap[T].$
Then, any valid column set must be contained in the complement $T]\setminus N(R)$, so the number of admissible columns equals
$|[T]\setminus N(R)|=T-|N(R)|.$
In order to choose a $(\bar{r}+1)\times(\bar{r}+1)$ tridiagonal exclusion minor we require
$T-|N(R)|\;\geq\;\bar{r}+1.$
Let $|R|=\bar{r}+1$. Each $r\in R$ contributes up to three elements $\{r-1,r,r+1\}$ to $N(R)$. Overlaps may occur. We can check that
$|N(R)|\;\geq\;\bar{r}+2,$
with equality if and only if $R$ is a block of $\bar{r}+1$ consecutive integers that touches a boundary (i.e. $R=\{1,\dots,\bar{r}+1\}$ or $R=\{T-\bar{r},\dots,T\}$). Any other placement yields $|N(R)|>\bar{r}+2$.
Now, let $T=2(\bar{r}+1)+1$. In this case
$T-|N(R)|\;\leq\;2(\bar{r}+1)+1-(\bar{r}+2)=\bar{r}+1.$
Hence the inequality $T-|N(R)|\geq\bar{r}+1$ forces equality, so we must have $|N(R)|=\bar{r}+2$. By Step 1 this occurs only when
$R=\{1,\dots,\bar{r}+1\}\quad\text{or}\quad R=\{T-\bar{r},\dots,T\}.$
We can use this result to pin down the possible composition of $C$. If $R=\{1,\dots,\bar{r}+1\}$ then
$N(R)=\{1,2,\dots,\bar{r}+2\},\qquad[T]\setminus N(R)=\{\bar{r}+3,\dots,2(\bar{r}+1)+1\}.$
Thus $C$ is forced to be $\{\bar{r}+3,\dots,2(\bar{r}+1)+1\}$. If instead $R=\{T-\bar{r},\dots,T\}$ then symmetrically $C=\{1,\dots,\bar{r}+1\}$.
The two ordered pairs $(R,C)$ obtained above are transposes of each other. In a symmetric matrix they yield the same minor value. Therefore, up to equality, there is exactly one tridiagonal exclusion minor of dimension $\bar{r}+1$. $\square$